Texinfo

The GNU Documentation Format

Edition 2.0 December 1990

by Robert J. Chassell and Richard M. Stallman

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988, 1990 Free Software Foundation, Inc.

This is edition 2.0 of the Texinfo documentation, and is consistent with version 2 of ‘texinfo.tex’.

Published by the Free Software Foundation 675 Massachusetts Avenue, Cambridge, MA 02139 USA Printed copies are available for $10 each. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

Texinfo Copying Conditions

The programs currently being distributed that relate to Texinfo include portions of GNU Emacs, plus other separate programs (makeinfo.c, info.c, texindex, and texinfo.tex). These programs are free; this means that everyone is free to use them and free to redistribute them on a free basis. The Texinfo-related programs are not in the public domain; they are copyrighted and there are restrictions on their distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of these programs that they might get from you.

Specifically, we want to make sure that you have the right to give away copies of the programs that relate to Texinfo, that you receive source code or else can get it if you want it, that you can change these programs or use pieces of them in new free programs, and that you know you can do these things.

To make sure that everyone has such rights, we have to forbid you to deprive anyone else of these rights. For example, if you distribute copies of the Texinfo related programs, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must tell them their rights.

Also, for our own protection, we must make certain that everyone finds out that there is no warranty for the programs that relate to Texinfo. If these programs are modified by someone else and passed on, we want their recipients to know that what they have is not what we distributed, so that any problems introduced by others will not reflect on our reputation.

The precise conditions of the licenses for the programs currently being distributed that relate to Texinfo are found in the General Public Licenses that accompany them. The programs that are part of GNU Emacs are covered by the GNU Emacs copying terms (see section ‘License’ in The GNU Emacs Manual), and other programs are covered by licenses that are contained in their source files.

2 Texinfo 2.0

1 Overview of Texinfo

Texinfo is a documentation system that uses a single source file to produce both on-line help (and other information) and a printed manual. This means that instead of writing two different documents, one providing on-line information and the other for a printed manual, you need write only one document. When the system is revised, you need revise only one file. (You can read the on-line help, known as an Info file, with the Info documentationreading program.)

Using Texinfo, you can create a printed document with the normal features of a book, including chapters, sections, cross references, and indices. From the same Texinfo source file, you can create a menu-driven, on-line Info file with nodes, menus, cross references, and indices. You can, if you wish, make the chapters and sections of the printed document correspond to the nodes of the on-line information; and you use the same cross references and indices for both the Info file and the printed document. The GNU Emacs Manual is a good example of a Texinfo file.

To make a printed manual, process a Texinfo source file with the TEX typesetting program. This creates a dvi file that you can typeset and print as a book. (If you do not have TEX, you can use the texi2roff program instead.) To create an Info file, process a Texinfo source file with the makeinfo or texinfo-format-buffer command; this creates an Info file that you can install on-line.

Info works with almost every type of computer terminal; similarly, TEX works with many types of printer. This power makes Texinfo a general purpose system, but brings with it a constraint, which is that a Texinfo file may contain only the customary “typewriter” characters (letters, numbers, spaces, and punctuation marks) but no special graphics.

A Texinfo file is a plain ASCII file containing text and @-commands (words preceded by an ‘@’) that tell the typesetting and formatting programs what to do. You may edit a Texinfo file with any text editor; but it is especially convenient to use GNU Emacs since that editor has a special mode, called Texinfo mode, that provides various Texinfo-related features. (See Chapter 2 [Texinfo Mode], page 15.)

Before writing a Texinfo source file, you should become familiar with the Info documentation reading program and learn about nodes, menus, cross references and the rest. (See Info file ‘info’, node ‘Info’, for more information.)

Texinfo creates both on-line help and a printed manual; moreover, it is freely redistributable. For these reasons, Texinfo is the standard format in which documentation for GNU utilities and libraries should be written.

1.1 Info files

A Texinfo file can be transformed into a printed manual and an Info file. An Info file is a file formatted so that the Info documentation reading

program can operate on it. Info files are divided into pieces called nodes, each of which contains the discussion of one topic. Each node has a name, and contains both text for the user to read and pointers to other nodes, which are identified by their names. The Info program displays one node at a time, and provides commands with which the user can move to the other related nodes.

Each node of an Info file may have any number of child nodes that describe subtopics of the node’s topic. The names of these child nodes, if any, are listed in a menu within the parent node; this allows you to use certain Info commands to move to one of the child nodes. Generally, a Texinfo file is organized like a book. If a node is at the logical level of a ‘chapter’, its child nodes are at the level of ‘sections’; likewise, the child nodes of sections are at the level of subsections.

All the children of any one parent are linked together in a bidirectional chain of ‘Next’ and ‘Previous’ pointers. This means that all the nodes that are logically parallel to sections within a chapter are all linked together. Normally the order in this chain is the same as the order of the children in the parent’s menu. Each child node records the parent node name, as its ‘Up’ pointer. The last child has no ‘Next’ pointer, and the first child normally has the parent both as its ‘Previous’ and as its ‘Up’ pointer.

The book-like structuring of an Info file into nodes that correspond to chapters, sections, and the like is a matter of convention, not a requirement. The ‘Up’, ‘Previous’, and ‘Next’ pointers of a node can point to any other nodes, and a menu can contain any other nodes. Thus, the node structure can be any directed graph. But it is usually more comprehensible to follow a structure that corresponds to the structure of chapters and sections in a printed manual.

In addition to ‘Next’, ‘Previous’, and ‘Up’ pointers and menus, Info provides another kind of pointer between nodes, called a reference, that can be sprinkled throughout a text. This is usually the best way to represent links that do not fit a hierarchical structure.

Usually, you will design a document so that its nodes match the structure of chapters and sections in the printed manual. But there are times when this is not right for the material being discussed. Therefore, Texinfo uses separate commands to specify the node structure of the Info file and the section structure of the printed manual.

Generally, you enter an Info file through a node that by convention is called ‘Top’. This node normally contains just a brief summary of the file’s purpose, and a large menu through which the rest of the file is reached. From this node, you can either traverse the file systematically by going from node to node, or you can go to a specific node listed in the main menu, or you can search the index menus and then go directly to the node that has the information.

If you want to read through an Info file in sequence, as if it were a printed manual, you can get the whole file with the advanced Info command g* RET. (See Info file ‘info’, node ‘Expert’.)

The ‘dir’ file in the ‘…/emacs/info’ directory serves as the departure point for the whole Info system. From it, you can reach the top nodes of each of the documents in a complete Info system.

1.2 Printed Manuals

A Texinfo file can be formatted and typeset as a printed manual. To do this, you need to use TEX, a powerful, sophisticated typesetting program written by Donald Knuth.^{1}

A Texinfo-based printed manual will be similar to any other book; it will have a title page, copyright page, table of contents, and preface, as well as chapters, numbered or unnumbered sections and subsections, page headers, cross references, footnotes, and indices.

You can use Texinfo to write a book without ever having the intention of converting it into on-line information. You can use Texinfo for writing a printed novel, and even to write a printed memo, although this latter application is not recommended since electronic mail is so much easier.

TEX is a general purpose typesetting program. Texinfo provides a file called ‘texinfo.tex’ that contains information (definitions or macros) that TEX uses when it typesets a Texinfo file. (The macros tell TEX how to convert the Texinfo @-commands to TEX commands, which TEX can then process to create the typeset document.) ‘texinfo.tex’ contains the specifications for printing a document, either with 7 inch by 9.25 inch pages or with 8.5 inch by 11 inch pages. (This is 178 mm by 235 mm or else 216 mm by 280 mm.) By changing the parameters in ‘texinfo.tex’, you can change the

1 You can also use the texi2roff program if you do not have TEX; since Texinfo is designed for use with TEX, texi2roff is not described here. texi2roff is part of the standard Emacs distribution. size of the printed document. In addition, you can change the style in which the printed document is formatted; for example, you can change the sizes and fonts used, the amount of indentation for each paragraph, the degree to which words are hyphenated, and the like. By changing the specifications, you can make a book look dignified, old and serious, or light-hearted, young and cheery.

TEX is freely distributable. It is written in a dialect of Pascal called WEB and can be compiled either in Pascal or (by using a conversion program that comes with the TEX distribution) in C. (See section ‘TeX Mode’ in The GNU Emacs Manual, for information about TEX.)

TEX is very powerful and has a great many features. Because a Texinfo file must be able to present information both on a character-only terminal in Info form and in a typeset book, the formatting commands that Texinfo supports are necessarily limited.

1.2.1 Obtaining TEX

TEX is freely distributable. You can obtain TEX for Unix systems from the University of Washington for a distribution fee.

To order a full distribution, send $140.00 for a 1/2 inch 9-track tape, $165.00 for two 4-track 1/4 inch cartridge tapes (foreign sites $150.00, for 1/2 inch, $175.00 for 1/4 inch, to cover the extra postage) payable to the University of Washington to:

The Director

Northwest Computer Support Group, DW-10 University of Washington Seattle, Washington 98195

Purchase orders are acceptable, but there is an extra charge of $10.00, to pay for processing charges. (Total of $150 for domestic sites, $175 for foreign sites).

The normal distribution is a tar tape, blocked 20, 1600 bpi, on an industry standard 2400 foot half-inch reel. The physical format for the 1/4 inch streamer cartridges uses QIC-11, 8000 bpi, 4-track serpentine recording for the SUN. Also, SystemV tapes can be written in cpio format, blocked 5120 bytes, ASCII headers.

1.3 @-commands

In a Texinfo file, the commands that tell TEX how to typeset the printed

manual and tell makeinfo and texinfo-format-buffer how to create an

Info file are preceded by ‘@’; they are called @-commands. For example, @node is the command to indicate a node and @chapter is the command to indicate the start of a chapter.

The Texinfo @-commands are a strictly limited set of constructs. The strict limits make it possible for Texinfo files to be understood both by TEX and by the code that converts them into Info files. You can display Info files on any terminal that displays alphabetic and numeric characters. Similarly, you can print the output generated by TEX on a wide variety of printers.

Depending on what they do or what arguments ^{2} they take, you need to write @-commands on lines of their own or as part of sentences:

• Write a command such as @page at the beginning of a line as the only text on the line.

• Write a command such as @chapter at the beginning of a line followed by the command’s arguments, in this case the chapter title, on the rest of the line.

• Write a command such as @dots{} where ever you wish, within a sentence or on a line of its own.

• Write a command such as @code{} where ever you wish (but usually

within a sentence) with its argument between the braces. As you gain experience with Texinfo, you will rapidly learn how to write the different commands: the different ways to write commands make sense and are designed to make Texinfo files easier to write and easier to read than if all commands followed exactly the same syntax.

All the @-commands, with the exception of the @TeX{} command, must be written entirely in lower case.

1.4 General Syntactic Conventions

All ASCII printing characters except ‘@’, ‘{’ and ‘}’ can appear in a Tex-

0 Texinfo 2.0

2 The word argument comes from the way it is used in mathematics and does not refer to a disputation between two people; it refers to the information presented to the command. According to the Oxford English Dictionary, the word derives from the Latin for to make clear, prove; thus it came to mean ‘the evidence offered as proof’, which is to say, ‘the information offered’, which led to its mathematical meaning. In its other thread of derivation, the word came to mean ‘to assert in a manner against which others may make counter assertions’, which led to the meaning of ‘argument’ as a disputation. info file and stand for themselves. ‘@’ is the escape character which introduces commands. ‘{’ and ‘}’ should be used only to surround arguments to certain commands. To put one of these special characters into the document, put an ‘@’ character in front of it, like this: ‘@@’, ‘@{’, and ‘@}’.

It is customary in TEX to use doubled single-quote characters to begin and end quotations: ‘``’ and ‘''’. This convention should be followed in Texinfo files. TEX converts doubled single-quote characters to left- and right-hand doubled quotation marks, “like this,” and Info converts doubled single-quote characters to regular double-quotes: ‘``’ and ‘''’ to ‘"’.

Use three hyphens in a row, ‘---’, for a dash—like this. In TEX, a single or even a double hyphen produces a printed dash that is shorter than you want. Info reduces three hyphens to two for display on the screen.

To prevent a paragraph from being indented in the printed manual, put the command @noindent on a line by itself before the paragraph.

If you mark off a region of the Texinfo file with the @iftex and @end iftex commands, that region will appear only in the printed copy; and you can use TEX commands in that region that cannot be used in Info. Likewise, if you mark off a region with the @ifinfo and @end ifinfo commands, that region will appear only in the Info file, and you can use Info commands in that region that cannot be used in TEX. (See Chapter 17 [Conditionals], page 129.)

Caution: Don’t use tabs in a Texinfo file! TEX has trouble with tabs: it treats them like single spaces, and that is not what they look like. This is a problem with TEX. (To avoid puting tabs into your file, you can set the indent-tabs-mode variable in Emacs to nil so that Emacs inserts multiple spaces when you press the TAB key. To get convert tabs in a file to multiple spaces, you can run untabify to convert tabs in a region to multiple spaces.)

1.5 Comments

You can write comments in a Texinfo file that will not appear in either the Info file or the printed manual by using the @comment command (which may be abbreviated to @c). Such comments are for the person who reads the Texinfo file. All the text on a line that follows either @comment or @c is a comment; the rest of the line does not appear in either the Info file or the printed manual.

You can write long stretches of text that will not appear in either the Info file or the printed manual by using the @ignore and @end ignore commands. Write each of these commands on a line of its own, starting each command at the beginning of the line. Text between these two commands does not appear in the processed output. You can use @ignore and @end ignore for writing comments or for holding text you may wish to use in another version of your document. Often, @ignore and @end ignore is used to enclose a part of the licensing permissions that applies to the Texinfo source file of a document, but not to the Info or printed version of the document.

1.6 What a Texinfo File Must Have

By convention, the names of Texinfo files end with either of the three extensions ‘.texinfo’, ‘.texi’, or ‘.tex’. The longer extension is preferred since it describes more clearly to a human reader the nature of the file. The shorter extensions are for operating systems that cannot handle long file names.

In order to be made into a printed manual and an Info file, a Texinfo file must begin with lines like this:

    \input texinfo

@setfilename info-file-name @settitle Name of Manual

The contents of the file follow this beginning, and then you must end a Texinfo file with a line like this:

    @bye

The ‘\input texinfo’ line tells TEX to use the ‘texinfo.tex’ file, which tells TEX how to translate the Texinfo @-commands into TEX typesetting commands. The ‘@setfilename’ line provides a name for the Info file and the ‘@settitle’ line specifies a title for the page headers (or footers) of the printed manual.

The @bye line at the end of the file on a line of its own tells TEX that the file is ended and to stop typesetting.

Usually, you won’t use quite such a spare format, but will include mode setting and start-of-header and end-of-header lines at the beginning of a Texinfo file, like this:

    \input texinfo   @c -*-texinfo-*-
    @c %**start of header

@setfilename info-file-name @settitle Name of Manual

    @c %**end of header

In the first line, ‘-*-texinfo-*-’ causes Emacs to switch into Texinfo mode when you edit the file.

The second and fifth lines, which surround the @setfilename and @settitle lines, are optional, but you need them if you run TEX or Info on just part of a file. (See Section 3.2.2 [Start of Header], page 29, for more information.)

Furthermore, you will usually provide a Texinfo file with a title page, master menu, and the like. But the minimum, which can be useful for short documents, is just the three lines at the beginning and the one line at the end.

1.7 Six Parts of a Texinfo File

Generally, a Texinfo file contains contains more than the minimal beginning and end—it usually contains six parts: Header The Header names the file, tells TEX which definitions’ file to

use, and performs other “housekeeping” tasks.

Summary Description and Copyright

The Summary Description and Copyright segment describes the document and contains the copyright notice and copying permissions for the Info file. The segment must be enclosed between @ifinfo and @end ifinfo commands so that it appears only in the Info file. It

Title and Copyright

The Title and Copyright segment contains the title and copyright pages and copying permissions for the printed manual. The segment must be enclosed between @titlepage and @end titlepage commands and appears only in the printed manual.

Top Node and Master Menu

The Master Menu contains a complete menu of all the nodes in the whole Info file. It appears only in the Info file, in the Top node.

Body The Body of the document may be structured like a traditional

book or encyclopedia or it may be free form.

End The End contains commands for printing indices and generating the table of contents, and the @bye command on a line of its own.

1.8 A Short Sample Texinfo File

Here is a complete but very short Texinfo file. The first three parts of the file, from ‘\input texinfo’ through to ‘@end titlepage’, look more intimidating than they are. Most of the material is standard boilerplate; when you write a manual, you just write the names for your own manual in this segment. (See Chapter 3 [Beginning a File], page 27.)

• Part 1. Header:

\input texinfo   @c -*-texinfo-*-
@c %**start of header
@setfilename sample.info
@settitle Sample Document
@c %**end of header
@setchapternewpage odd

• Part 2. Summary Description and Copyright

The summary description and copyright segment does not appear in the printed document.

@ifinfo
This is a short example of a complete Texinfo file.
Copyright @copyright{} 1990 Free Software Foundation, Inc.
@end ifinfo

• Part 3. Titlepage and Copyright

The titlepage segment does not appear in the Info file.

@titlepage
@sp 10
@comment The title is printed in a large font.
@center @titlefont{Sample Title}
@c The following two commands start the copyright page.
@page
@vskip 0pt plus 1filll
Copyright @copyright{} 1990 Free Software Foundation, Inc.
@end titlepage

• Part 4. Top Node and Master Menu

The Top node contains the master menu for the Info file. Since a printed manual uses a table of contents rather than a menu, the master menu appears only in the Info file.

@node    Top,       First Chapter, (dir),    (dir)
@comment node-name, next,          previous, up
@menu
* First Chapter::    The first chapter is the
                     only chapter in this sample.
@end menu

• Part 5: The Body of the Document

@node     First Chapter,    , Top,      Top
@comment  node-name,    next, previous, up
@chapter First Chapter
@cindex Sample index entry
This is the contents of the first chapter.
@cindex Another sample index entry
Here is a numbered list.
@enumerate
@item
This is the first item.
@item
This is the second item.
@end enumerate
The @code{makeinfo} and @code{texinfo-format-buffer}
commands transform a Texinfo file such as this into
an Info file; and @TeX{} typesets it for a printed
manual.

• Part 6: The End of the Document

@node    Concept Index,    ,  First Chapter, Top
@comment node-name,    next,  previous,      up
@unnumbered Concept Index
@printindex cp
@contents
@bye

Here is what the contents of the first chapter of the sample look like:

This is the contents of the first chapter. Here is a numbered list.

1. This is the first item.

2. This is the second item.

The makeinfo and texinfo-format-buffer commands transform

a Texinfo file such as this into an Info file; and TEX typesets it for a printed manual.

14 Texinfo 2.0

2 Using Texinfo Mode

In GNU Emacs, Texinfo mode provides commands and features especially designed for working with Texinfo files:

• Insert commonly used strings of text.

• Show the structure of a Texinfo source file.

• Automatically insert or update the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node.

• Automatically insert or update menus.

• Automatically create a master menu.

• Format a part or all of a file for Info.

• Typeset and print part or all of a file.

The special Texinfo commands are in addition to the usual editing commands, which are generally the same as the commands of Text mode. However, in Texinfo mode the paragraph separation variable and syntax table are redefined so that Texinfo commands that should be on lines of their own are not inadvertently included in paragraphs. Thus, the M-q (fill-paragraph) command will refill a paragraph but not mix an indexing command on a line adjacent to it into the paragraph.

You may name a Texinfo file however you wish, but the convention is to end a Texinfo file name with one of the three extensions ‘.texinfo’, ‘.texi’, or ‘.tex’. A longer extension is preferred, since it is explicit, but a shorter extension may be necessary for operating systems that limit the length of file names. GNU Emacs automatically enters Texinfo mode when you visit a file with any of these extensions. Also, Emacs switches to Texinfo mode for a file that has ‘-*-texinfo-*-’ in its first line. If ever you are in another mode and wish to switch to Texinfo mode, type M-x texinfo-mode.

Like all other Emacs features, you can customize or enhance Texinfo mode as you wish. In particular, the keybindings are very easy to change. The keybindings described here are the default or standard ones.

2.1 Inserting Frequently Used Commands

Texinfo mode provides commands to insert various frequently used @- commands into the buffer. You can use these commands to save keystrokes.

The insert commands are invoked by typing C-c twice and then the first letter of the @-command.

C-c C-c c
texinfo-insert-@code

Insert @code{} and put the cursor between the braces.

C-c C-c d
texinfo-insert-@dfn

Insert @dfn{} and put the cursor between the braces.

C-c C-c e
texinfo-insert-@end

Insert @end{}.

C-c C-c i
texinfo-insert-@item

Insert @item{} and put the cursor at the beginning of the next line.

C-c C-c n
texinfo-insert-@node

Insert @node and a comment line listing the sequence for the ‘Next’, ‘Previous’, and ‘Up’ nodes. Leave point after the @node.

C-c C-c s
texinfo-insert-@samp

Insert @samp{} and put the cursor between the braces.

C-c C-c v
texinfo-insert-@var

Insert @var{} and put the cursor between the braces.

C-c C-c {
texinfo-insert-braces

Insert {} and put the cursor between the braces.

C-c C-c }
up-list

Move forward past one set of closing braces. This set of insert commands was created after analyzing the frequency

with which different @-commands are used in the GNU Emacs Manual and the GDB Manual. If you wish to add your own insert commands, you can bind a keyboard macro to a key, use abbreviations, or extend the code in

‘texinfo.el’.

2.2 Showing the Section-Structure of a File

You can show the section-structure of a Texinfo file by using the Cc C-s command (texinfo-show-structure). This command shows the section-structure of a Texinfo file by listing the lines that begin with the @- commands for @chapter, @section, and the like. The command constructs what amounts to a table of contents. These lines are displayed in another window called the ‘*Occur*’ window. In that window, you can position the cursor over one of the lines and use the C-c C-c command (occur-mode-goto-occurrence), to jump to the corresponding spot in the Texinfo file.

C-c C-s
texinfo-show-structure

Show the @chapter, @section, and such lines of a Texinfo file.

C-c C-c
occur-mode-goto-occurrence

Go to the line in the Texinfo file corresponding to the line under the cursor in the ‘*Occur*’ buffer.

If you call texinfo-show-structure with a prefix argument by typing Cu C-c C-s, it will list not only those lines with the @-commands for @chapter, @section, and the like, but also the @node lines. (This is how the texinfoshow-structure command worked without an argument in the first version of Texinfo. It was changed because @node lines clutter up the the ‘*Occur*’ window and are usually not needed.) You can use texinfo-show-structure with a prefix argument to check whether the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node line are correct.

Often, when you are working on a manual, you will be interested only in the structure of the current chapter. In this case, you can mark off the region of the buffer that you are interested in with the C-x n (narrow-to-region) command and texinfo-show-structure will work on only that region. To see the whole buffer again, use C-x w (widen). (See section ‘Narrowing’ in The GNU Emacs Manual, for more information about the narrowing commands.)

2.3 Updating Nodes and Menus

Texinfo mode provides commands for automatically creating or updating menus and node pointers. The commands are called ‘update’ commands because their most frequent use is for updating a Texinfo file after you have worked on it; but you can use them to insert the ‘Next’, ‘Previous’, and ‘Up’ pointers into a node line that has none and to create menus in a file that has none.

You can use the updating commands • to insert or update the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node,

• to insert or update the menu for a section, and

• to create a master menu for a Texinfo source file. You can also use the commands to update all the nodes and menus in a

region or in a whole Texinfo file.

Texinfo mode has five updating commands that are used most often: two are for updating the node pointers or menu of a single node (or a region), two are for updating every node pointer and menu in a file, and one is for creating a master menu for a complete file. (When using a command interactively, type C-u before typing the command to give it a prefix argument.) The major updating commands are:

C-c C-u C-n
texinfo-update-node

Insert the ‘Next’, ‘Previous’, and ‘Up’ pointers for the node point is within (i.e., for the @node line preceding point). It does not matter whether the @node line has pre-existing ‘Next’, ‘Previous’, or ‘Up’ pointers in it. The old are removed and new inserted. With an argument (prefix argument, if interactive), this command updates the region (which is the text between point and mark).

C-c C-u C-m
texinfo-make-menu

Create or update the menu in the node point is within. With an argument (prefix argument, if interactive), the command makes or updates menus for the nodes within or part of the region. Whenever texinfo-make-menu updates an existing menu, the descriptions from that menu are incorporated into the new menu. This is done by copying descriptions from the existing menu to the entries in the new menu that have the same node names. If the node names are different, the descriptions are not copied to the new menu. Menu entries that refer to other Info files are removed since they do not refer to nodes within the current buffer. This is a deficiency.

C-c C-u C-e
texinfo-every-node-update

Insert or update the ‘Next’, ‘Previous’, and ‘Up’ pointers for every node in the buffer .

C-c C-u C-a
texinfo-all-menus-update

Create or update all the menus in the buffer, except for the master menu. With an argument (prefix argument, if interactive), first insert or update all the node pointers before working on the menus.

If a master menu exists, texinfo-all-menus-update removes the detailed part of the master menu, replacing the master menu with a regular top-level menu. (A master menu lists all the nodes in an Info file; a regular top-level menu lists just the chapters. Since the descriptions in the detailed part of a master menu come from the lower level menus, removing the detailed part does not destroy any information.)

M-x texinfo-master-menu

Create or update a master menu that includes all the other menus (incorporating the descriptions from pre-existing menus, if any). With an argument (prefix argument, if interactive), first create or update all the nodes and all the regular menus in the buffer before constructing the master menu. (See Section 3.5 [The Top Node and Master Menu], page 37, for more about a master menu.)

For texinfo-master-menu to work, the Texinfo file must have a ‘Top’ node.

After extensively editing a Texinfo file, it is common to type C-u M-x texinfo-master-menu to update all the nodes and menus completely and all at once.

The texinfo-column-for-description variable specifies the column to

which menu descriptions are indented. By default, the value is 32 although it is often useful to reduce it to as low as 24. You can set the variable with the M-x edit-options command (see section ‘Editing Variable Values’ in The GNU Emacs Manual) or with the M-x set-variable command (see section ‘Examining and Setting Variables’ in The GNU Emacs Manual).

Also, the texinfo-indent-menu-description command may be used

to indent existing menus to a specified column. See Section 2.3.2 [Other Updating Commands], page 20.

Finally, if you wish, you can use the texinfo-insert-node-lines command to insert missing @node lines into a file. (See Section 2.3.2 [Other Updating Commands], page 20, for more information.)

2.3.1 Updating Requirements

To use the updating commands, you must organize the Texinfo file hierarchically with chapters, sections, subsections, and the like. Each @node line, with the exception of the line for the ‘Top’ node, must be followed by a line with a structuring command such as @chapter, @section, or

@unnumberedsubsec.

Each @node line/structuring-command line combination must look either like this:

    @node     Comments,  Minimum, Conventions, Overview
    @comment  node-name, next,    previous,    up
    @section Comments

or like this (without the @comment line):

    @node Comments, Minimum, Conventions, Overview
    @section Comments

If the file has a ‘Top’ node, it must be called ‘top’ or ‘Top’ and be the first node in the file.

Incidently, the makeinfo command will create an Info file for hierarchically organized Texinfo file that lacks ‘Next’, ‘Previous’ and ‘Up’ pointers. Thus, if you can be sure that your Texinfo file will be formatted with makeinfo, you have no need for the ‘update node’ commands. (See Chapter 19 [Creating an Info File], page 139, for more information about

makeinfo.) However, both makeinfo and the texinfo-format-. . . com-

mands require menus.

2.3.2 Other Updating Commands

In addition to the five major updating commands, Texinfo possesses several less frequently used updating commands.

M-x texinfo-indent-menu-description

Indent every description in the menu following point to the specified column. You can use this command to give yourself more space for descriptions. With an argument (prefix argument, if in-

          teractive), the texinfo-indent-menu-description command

indents every description in every menu in the region. However, this command does not indent the second and subsequent lines of a multi-line description.

M-x texinfo-insert-node-lines

Insert @node before the @chapter, @section, and other sectioning commands in a region in a Texinfo file where the @node lines are missing. With an argument (prefix argument, if inter-

active), the texinfo-insert-node-lines command not only

inserts @node lines but also inserts the chapter or section titles as the names of the corresponding nodes; and it inserts their titles for node names in pre-existing @node lines that lack names. Since node names should be more concise than section or chapter titles, node names so inserted should be edited manually.

M-x texinfo-sequentially-update-node

Insert the names of the nodes immediately following and preceding the current node as the ‘Next’ or ‘Previous’ pointers regardless of those nodes’ hierarchical level. This means that the ‘Next’ node of a subsection may well be the next chapter. Sequentially ordered nodes are useful for novels and other documents that you read through sequentially. (However, in Info, the g* RET command lets you look through the file sequentially, so sequentially ordered nodes are not strictly necessary.) With an argument (prefix argument, if interactive), the texinfo-sequentially-update-node command sequentially updates all the nodes in the region.

M-x texinfo-multiple-files-update

Update nodes and menus in a document built from several separate files. With a prefix argument if called interactively (a non-nil ‘make-master-menu’ argument, if called from noninteractively), create and insert a master menu in the outer file. With a numeric prefix argument if called interactively (a nonnil ‘update-everything’ argument if called non-interactively), first update all the menus and all the ‘Next’, ‘Previous’, and ‘Up’ pointers of all the included files before creating and inserting a master menu in the outer file. The texinfo-multiple-files-update command is described in the appendix on @include files. See Appendix B[Include Files], page 165.

2.4 Formatting for Info

Texinfo mode provides several commands for formatting part or all of a Texinfo file for Info. Often, when you are writing a document, you want to format only part of a file, a region.

You can use either the texinfo-format-region or the makeinfo-region

command to format a region.

C-c C-i C-r
texinfo-format-region
C-c C-m C-r
makeinfo-region

Format the current region for Info.

   You can use either the texinfo-format-buffer or the makeinfo-buffer

command to format a whole buffer:

C-c C-i C-b
texinfo-format-buffer
C-c C-m C-b
makeinfo-buffer

Format the current buffer for Info. For any of the Info formatting commands to work, the file must include

a line that has @setfilename in its header.

Not all systems support the makeinfo-based formatting commands. See Chapter 19 [Creating an Info File], page 139, for details about Info

formatting.

2.5 Formatting and Printing

Typesetting and printing a Texinfo file is a multi-step process in which you first create a file for printing (called a DVI file), and then you print the file. Optionally, also, you may create indices.

Often, when you are writing a document, you want to typeset and print only part of a file, to see what it will look like. You can use the texinfotex-region and related commands for this purpose. Use the texinfo-tex-buffer command to format all of a buffer.

C-c C-t C-r
texinfo-tex-region

Run TEX on the region.

C-c C-t C-b
texinfo-tex-buffer

Run TEX on the buffer.

C-c C-t C-p
texinfo-tex-print

Print the file (or the part of the file) previously formatted with

          texinfo-tex-buffer or texinfo-tex-region.

For texinfo-tex-region or texinfo-tex-buffer to work, the file must

start with a ‘\input texinfo’ line and must include an @settitle line between start of header and end of header lines. The file must end with @bye on a line by itself.

See Chapter 18 [Printing Hardcopy], page 131, for a description of the other TEX related commands, such as texinfo-texindex and tex-show-

print-queue.

2.6 Texinfo Mode Summary

In Texinfo mode, each set of commands has default keybindings that begin with the same keys. All the commands that are custom-created for Texinfo mode begin with C-c. The keys that follow are arranged mnemonically.

Insert Commands

The insert commands are invoked by typing C-c twice and then the first letter of the @-command to be inserted.

    C-c C-c c  Insert `@code'.
    C-c C-c d  Insert `@dfn'.
    C-c C-c e  Insert `@end'.
    C-c C-c i  Insert `@item'.
    C-c C-c n  Insert `@node'.
    C-c C-c s  Insert `@samp'.
    C-c C-c v  Insert `@var'.

C-c C-c { Insert braces.

C-c C-c } Move out of enclosing braces.

Show Structure

The texinfo-show-structure command is most often used within a narrowed region.

C-c C-s List all the headings.

Update Pointers

The update pointer commands are invoked by typing C-c C-u and then

either C-n for texinfo-update-node or C-e for texinfo-every-node-
update.

C-c C-u C-n Update a node.

C-c C-u C-e Update every node in the buffer.

Update Menus

The update menu commands are invoked by typing C-c C-u and then

either C-m for texinfo-make-menu or C-a for texinfo-all-menus-update.

The texinfo-master-menu command lacks a keybinding since it is used infrequently.

C-c C-u C-m Make or update a menu. C-c C-u C-a Make or update all the menus

in a buffer; with an argument, first update all the nodes.

    M-x texinfo-master-menu

Create or update a master menu. With an argument, first create or update all nodes and regular menus.

Format for Info

The Info formatting commands that are written in Emacs Lisp are invoked by typing C-c C-i and then either C-r for a region or C-b for the whole buffer.

The Info formatting commands that are written in C and based on the makeinfo.c program are invoked by typing C-c C-m and then either C-r for a region or C-b for the whole buffer. Not all systems support the makeinfo program used by these commands. Use the texinfo-format… commands:

C-c C-i C-r Format the region. C-c C-i C-b Format the buffer.

Use makeinfo:

C-c C-m C-r Format the region. C-c C-m C-b Format the buffer.

Typeset and Print

The TEX typesetting and printing commands are invoked by typing C-c C-t and then another control command: C-r for texinfo-tex-region, C-b

for texinfo-tex-buffer, and so on.

C-c C-t C-r Run TEX on the region. C-c C-t C-b Run TEX on the buffer.

C-c C-t C-i  Run texindex.

C-c C-t C-p Print the dvi file. C-c C-t C-q Show the print queue. C-c C-t C-k Kill the current job. C-c C-t C-l Recenter the output buffer.

26 Texinfo 2.0

3 Beginning a Texinfo File

Various pieces of information must be provided at the beginning of a Texinfo file, such as the name of the file and the title of the document.

Generally, the beginning of a Texinfo file has four parts:

1. The header, with its start of header and end of header lines, that includes the commands for naming the Texinfo file and telling TEX what definitions’ file to use when processing the file.

2. A short statement of what the file is about, with a copyright notice and copying permissions. This is enclosed in @ifinfo and @end ifinfo commands so that it appears only in the Info file.

3. A title page and copyright page, with a copyright notice and copying permissions. This is enclosed in @titlepage and @end titlepage commands and appears only in the printed manual.

4. The ‘Top’ node that contains a master extensive menu for the whole Info file. The contents of this node appear only in the Info file. If the Texinfo file has a section containing licensing information and a

warranty disclaimer, that section usually follows the ‘Top’ node. The licensing section will be followed by an introduction or else by the first chapter of the manual.

Since the copyright notice and the copying permissions are in sections that appear only in the Info file or only in the printed manual, this information must be given twice.

3.1 Sample Texinfo File Beginning

The following sample shows what is needed.

\input texinfo   @c -*-texinfo-*-
@c %**start of header

@setfilename name-of-info-file @settitle Name of Manual

@setchapternewpage odd
@c %**end of header
@ifinfo
This file documents . . .
Copyright @copyright{} year copyright-owner

Permission is granted to . . .
@end ifinfo

@c  This title page illustrates only one of the
@c  two methods of forming a title page.
@titlepage

@title Name of Manual When Printed @subtitle Subtitle, If Any @subtitle Second Subtitle @author by Author

@c  The following two commands
@c  start the copyright page.
@page
@vskip 0pt plus 1filll
Copyright @copyright{} year copyright-owner
Published by . . .
Permission is granted to . . .
@end titlepage

@node Top, Overview, (dir), (dir)
@ifinfo
Summary . . .
@end ifinfo
@menu
* License::          Your rights and freedoms.
* First Chapter::    Getting started . . .
* Second Chapter:: . . .
  <many more menu items here>
@end menu
@node    First Chapter, Second Chapter, top,      top
@comment node-name,     next,           previous, up
@chapter First Chapter
@cindex Reference to First Chapter

3.2 The Texinfo File Header

Texinfo files start with at least three lines that provide Info and TEX with necessary information. These are the \input texinfo line, the @settitle line, and the @setfilename line. If you want to run TEX on just a part of the Texinfo File, you also have to surround the ‘@settitle’ and @setfilename lines with start-of-header and end-of-header lines.

3.2.1 The First Line of a Texinfo File

Every Texinfo file that is to be the top-level input to TEX must begin with a line that looks like this:

    \input texinfo   @c -*-texinfo-*-

The line serves two functions:

1. When the file is processed by TEX, the ‘\input texinfo’ command tells TEX to load the macros needed for processing a Texinfo file. These are in a file called ‘texinfo.tex’, which is usually located in

    the ‘/usr/lib/tex/macros’ directory.

2. When the file is edited in GNU Emacs, the ‘-*-texinfo-*-’ mode specification tells Emacs to switch to Texinfo mode.

3.2.2 Start of Header

Your should write a start-of-header line on the second line of a Texinfo file. Follow the start-of-header line with @setfilename and @settitle lines (and, optionally, an @setchapternewpage command line), and then by an end-of-header line.

With these lines, you can format part of a Texinfo file for Info or typeset part for printing.

A start-of-header line looks like this:

    @c %**start of header

The odd string of characters (‘%**’) is so the region formatting command does not accidentally find something it shouldn’t when it looks for the header.

3.2.3 @setfilename

In order to be made into an Info file, a Texinfo file must contain a line that looks like this:

@setfilename info-file-name

Write the @setfilename command at the beginning of a line followed by the Info file name.

The @setfilename line specifies the name of the Info file to be generated. The Info formatting commands ignore everything written before this line, which is why the very first line of the the file (the \input line) does not need to be commented out. The @setfilename line is ignored when you typeset a printed manual.

3.2.4 @settitle

In order to be made into a printed manual, a Texinfo file must contain a line that looks like this:

    @settitle title

Write the @settitle command at the beginning of a line followed by the title name. This tells TEX the title to use in a header or footer.

Conventionally, TEX formats a Texinfo file for double-sided output so as to print the title in the left-hand (even-numbered) page headings and the current chapter titles in the right-hand (odd-numbered) page headings. (TEX learns the name of each chapter from each @chapter command.) Page footers are not printed.

Even if you are printing in a single-sided style, TEX looks for an @settitle command line, in case you include the manual title in the heading.

The @settitle command should precede everything that generates actual output in TEX.

The title in the @settitle command is usually the same as the real title on the title page, but not always. It can be a shortened or expanded version of the title.

TEX prints page headings only for that text that comes after the @end titlepage command in the Texinfo file, or that comes after an @headings command that turns on headings. (See Section 3.4.6 [The @headings Command], page 36, for more information.)

You may, if you wish, create your own, customized headings and footing. See Appendix C[Page Headings], page 169, for a detailed discussion of this process.

3.2.5 @setchapternewpage

In a book or a manual, usually text is printed on both sides of the sheets of paper, chapters start on right-hand pages, and right-hand pages have odd numbers. But in short reports, text often is printed only on one side of the sheets of paper. Also, in short reports, chapters sometimes do not start on new pages, but are printed on the same page as the end of the preceding chapter, after a small amount of vertical whitespace.

You can use the @setchapternewpage command with various arguments to specify how TEX should start chapters and whether it should typeset pages for printing on one side of sheets of paper (single-sided printing) or on both sides (double-sided printing).

Write the @setchapternewpage command at the beginning of a line followed by its argument.

For example, you would write the following to cause each chapter to start on a fresh odd-numbered page:

    @setchapternewpage odd

There are three alternatives for the @setchapternewpage command:

@setchapternewpage off

This tells TEX to typeset for double-sided printing, but not to start new chapters on new pages.

@setchapternewpage on

This tells TEX to typeset for single-sided printing and to start new chapters on new pages. This is the form most often used for short reports. This alternative is the default.

@setchapternewpage odd

This tells TEX to typeset for double-sided printing and to start new chapters on new, odd-numbered pages (right-handed pages). This is the form most often used for books and manuals.

Texinfo does not have an @setchapternewpage even command. (You can countermand or modify an @setchapternewpage command with an @headings command. See Section 3.4.6 [The @headings Command], page 36.)

At the beginning of a manual or book, pages are not numbered—for example, the title and copyright pages of a book are not numbered. By convention, table of contents pages are numbered in roman numerals as their own series.

Since an Info file does not have pages, the @setchapternewpage command has no effect on it.

Usually, you do not use an @setchapternewpage command when you are printing single-sidedly, and use the @setchapternewpage odd command when you are printing double-sidedly.

3.2.6 End of Header

Follow the line that contains the @settitle or @setchapternewpage command with the end-of-header line.

An end-of-header line looks like this:

    @c %**end of header

If you include the @setchapternewpage command between the start-of-header and end-of-header lines, TEX will typeset a region as that command specifies.

See Section 3.2.2 [Start of Header], page 29.

3.3 Summary and Copying Permissions for Info

Since the title page and the copyright page appear only in the printed copy of the manual, the same information must be inserted in a section that appears only in the Info file. This section usually contains a brief description of the contents of the Info file, a copyright notice, and copying permissions.

The copyright notice should read:

Copyright year copyright-owner and be put on a line by itself.

Standard text for the copyright permissions is contained in the appendix. See Section F.1 [ifinfo Permissions], page 188, for the complete text.

3.4 The Title and Copyright Pages

A manual’s name and author are usually printed on a title page. Sometimes copyright information is printed on the title page as well; more often, copyright information is printed on the back of the title page.

The title and copyright pages appear in the printed manual, but not in the Info file. Because of this, it is possible to use several slightly obscure TEX typesetting commands that could not be used in an Info file. In addition, this part of the beginning of a Texinfo file contains the text of the copying permissions that will appear in the printed manual.

3.4.1 @titlepage

Start the material for the title page and following copyright page with @titlepage on a line by itself and end it with @end titlepage on a line by itself.

The @end titlepage command starts a new page and turns on page numbering. (See Appendix C[Page Headings], page 169, for details about the generation of headings.) All the material that you want to appear on unnumbered pages should be put between the @titlepage and @end titlepage commands. By using the @page command you can force a page break within the region delineated by the @titlepage and @end titlepage commands and create more than one unnumbered page. This is how the copyright page is produced. (The @titlepage command might perhaps have been better named the @titleandadditionalpages command, but that would have been rather long!)

Texinfo provides you with two methods for creating a title page. One method uses the @titlefont, @sp, and @center commands to generate a title page in which the words on the page are centered.

The second method uses the @title, @subtitle, and @author to create a title page with black rules under the title and author lines and the subtitle text set flush to the right hand side of the page. You may use either method.

3.4.2 @titlefont, @center, and @sp

You can use the @titlefont, @sp, and optionally the @center commands to create a title page for a printed document.

Use the @titlefont command to select a large font suitable for the title itself. For example:

    @titlefont{Texinfo Manual}

Use the @center command at the beginning of a line to center the remaining text on that line. Thus,

    @center @titlefont{Texinfo Manual}

centers the title, which in this example is “Texinfo Manual” printed in the title font.

Use the @sp command to adjust vertical spacing. For example:

    @sp 2

This inserts two blank lines on the printed page. (See Section 16.3 [@sp], page 126, for more information about the @sp command.)

A typical title page looks like the following:

    @titlepage
    @sp 10

@center @titlefont{Name of Manual When Printed}

    @sp 2

@center Subtitle, If Any

    @sp 2
    @center Author

The spacing of the example fits an 8 1/2 by 11 inch manual.

3.4.3 @title, @subtitle, and @author

You can use the @title, @subtitle, and @author commands to create a title page in which the vertical and horizontal spacing is done for you. This contrasts with the method described in the previous section, in which the @sp command is needed to adjust vertical spacing.

Write the @title, @subtitle, or @author commands at the beginning of a line followed by the title, subtitle, or author.

The @title command produces a line in which the title is set flush to the left-hand side of the page in a larger then normal font. The title is underlined with a black rule.

The @subtitle command sets subtitles in a normal-sized font flush to the right-hand side of the page.

The @author command sets the names of the author in a middle-sized font flush to the left-hand side of the page on a line near the bottom of the title page. The author’s name is underlined with a black rule that is thinner than the rule that underlines the title. (The black rule only occurs if the @author command line is followed by an @page command line.)

There are two ways to use the @author command: you can write the name or names on the remaining part of the line that starts with an @author command:

    @author by Jane Smith and John Doe

or you can write the names one above each other by using two (or more) @author commands:

    @author Jane Smith
    @author John Doe

(Only the bottom name is underlined with a black rule.)

Thus, a typical title page looks like the following:

@title Name of Manual When Printed @subtitle Subtitle, If Any @subtitle Second Subtitle

    @author Author

3.4.4 Copyright Page and Permissions

By international treaty, the copyright notice for a book should be either on the title page or on the back of the title page. The copyright notice should include the year followed by the name of the person or organization who owns the copyright.

When the copyright notice is on the back of the title page, that page is customarily not numbered. Therefore, in Texinfo, the information on the copyright page should be within @titlepage and @end titlepage commands.

Use the @page command to cause a page break. To push the copyright notice and the other text on the copyright page towards the bottom of the page, you can write a somewhat mysterious line after the @page command that reads like this:

    @vskip 0pt plus 1filll

This is a line that uses TEX commands that are not supported by the Info formatting commands. The @vskip command means to skip lines and put in white space. The ‘0pt plus 1filll’ means to put in zero points of mandatory white space, and as much optional white space as needed to push the following text towards the bottom of the page. Note the use of three ‘l’s in the word ‘filll’; this is the correct usage in TEX.

In a printed manual, the @copyright{} command generates a ‘c’ inside a circle. (In Info, it generates ‘(C)’.) The copyright notice itself has the following legally defined sequence:

Copyright \begin{matrix}c \\ ◯ \end{matrix} year copyright-owner It is customary to put information on how to get a manual after the

copyright notice, before the permissions.

Note that permissions must be given here as well as in the summary segment within @ifinfo and @end ifinfo that immediately follows the header since this section appears only in the printed manual and the ‘ifinfo’ segment appears only in the Info file.

See Appendix F[Sample Permissions], page 187, for recommended permission text.

3.4.5 Heading Generation

An @end titlepage command (on a line by itself) not only marks the end of the title and copyright pages but it causes TEX to start generating page headings and page numbers.

To repeat what is said elsewhere (see Section 3.2.5 [@setchapternewpage], page 31), Texinfo has two standard heading formats, one for documents printed on one side of each sheet of paper (single-sided printing), and the other for documents printed on both sides of each sheet (double-sided printing). You can specify these formats in different ways:

• The conventional way is to write an @setchapternewpage command before the title page commands, and then have the @end titlepage command start generating headings in the manner desired. (See Section 3.2.5 [@setchapternewpage], page 31.)

• Alternatively, you can use the @headings command to prevent headings from being generated or to start them for either single or double-sided printing. See Section 3.4.6 [The @headings Command], page 36.

• Or, you may specify your own heading and footing format. See Appendix C[Page Headings], page 169, for detailed information about headings and footings. Most documents are formatted with the standard single-sided or double-

sided format, using @setchapternewpage odd for double-sided printing and no @setchapternewpage command for single-sided printing.

3.4.6 The @headings Command

The @headings command is rarely used. It specifies what kind of page headings and footings to print on each page. Usually, this is controlled by the @setnewchapter command. You need the @headings command only if the @setnewchapter command does not do what you want.

There are four ways to use the @headings command:

@headings off

Turns off printing of headings.

@headings single

Turns on headings appropriate for single-sided printing.

@headings double
@headings on

Turns on headings appropriate for double-sided printing. The

          two commands, @headings on and @headings double, are syn-

onymous. To prevent TEX from generating any page headings, write @headings

off on a a line of its own immediately after the line containing the @end titlepage command. The @headings off command overrides the @end titlepage command, which would otherwise cause TEX to print headings.

You can also specify your own style of heading and footing. See Appendix C[Page Headings], page 169, for more information.

3.5 The Top Node and Master Menu

A ‘Top’ node should contain a brief description of the file and an extensive, master menu for the whole Info file. The contents of this node should appear only in the Info file; none of it should appear in printed output.

TEX does not print either an @node line or a menu; they appear only in Info. But other text will be printed unless you enclose it between @ifinfo and @end ifinfo commands.

Sometimes, you will want to place an @chapter line containing the title of the document after the @node line, followed by a short summary. This helps the reader understand what the Info file is about. You must enclose these lines between @ifinfo and @end ifinfo commands; otherwise, they will appear in the printed manual and its table of contents.

For example, the beginning of the Top node of this manual looks like this:

…

    @end titlepage
    @ifinfo
    @node Top, Copying, (dir), (dir)
    @chapter Texinfo
    Texinfo is a documentation system. . .

…

    @end ifinfo
    @menu
    * Copying::                 Texinfo is freely
                                  redistributable.

* Overview::                What is Texinfo?

…

@end menu

(As you can see from this example, you can place the @ifinfo command before the @node line as well as after it. See Chapter 17 [Conditionals], page 129, for more information about the use of @ifinfo.)

In a ‘Top’ node, the ‘Previous’, and ‘Up’ nodes usually refer to the top level directory of the whole Info system, which is called ‘(dir)’.

3.5.1 Parts of a Master Menu

The master menu is enclosed in @menu and @end menu commands. The menu does not appear in TEX output.

Generally, a master menu is divided into parts.

• The first part contains the major nodes in the Texinfo file: the nodes for the chapters, chapter-like sections, and the major appendices.

• The second part contains nodes for the indices.

• The third and subsequent parts contain a listing of the other, lower level nodes, often ordered by chapter. This way, rather than go through an intermediary menu, an inquirer can go directly to a particular node when searching for specific information. (These menu items are not required; use them if you think they are a convenience.) Each section in the menu can be introduced by a descriptive line. So long

as the line does not begin with an asterisk, it will not be treated as a menu item. (See Section 6.3 [Making Menus], page 55, for more information.)

For example, the master menu for this manual looks like the following (but has many more entries):

@menu
* Copying::             Texinfo is freely
                          redistributable.
* Overview::            What is Texinfo?
* Texinfo Mode::        Special features in GNU Emacs.

… …

* Command Index::       An item for each @-command.
* Concept Index::       An item for each concept.
 --- The Detailed Node Listing ---

    Overview of Texinfo
    * Info Files::          What is an Info file?
    * Printed Manuals::     Characteristics of
                              a printed manual.

… …

    Using Texinfo Mode
    * Info on a Region::    Formatting part of a file
                              for Info.

… …

    @end menu

3.6 Licensing and Distribution Information

If the Texinfo file has a section containing the “General Public License” and the distribution information and a warranty disclaimer, this section usually follows the ‘Top’ node. The General Public License is very important to Project GNU software. It ensures that you and others will continue to have a right to use and share the software.

The copying and distribution information and the disclaimer are followed by an introduction or else by the first chapter of the manual.

Although an introduction is not a required part of a Texinfo file, it is very helpful. Ideally, it should state clearly and concisely what the file is about and who would be interested in reading it. In general, an introduction would follow the licensing and distribution information, although sometimes people put it earlier in the document. Usually, an introduction is put in an

@unnumbered section. (See Section 5.4 [The @unnumbered and @appendix

Commands], page 47.)

40 Texinfo 2.0

4 Ending a Texinfo File

The end of a Texinfo file should include the commands that create indices and generate detailed and summary tables of contents. And it must include the @bye command that marks the last line that TEX processes.

For example,

    @node    Concept Index,     , Variables Index, Top
    @comment node-name,     next, previous,      up
    @unnumbered Concept Index
    @printindex cp
    @contents
    @bye

4.1 Index Menus and Printing an Index

To print an index means to include it as part of a manual or Info file. This does not happen automatically just because you use @cindex or other index-entry generating commands in the Texinfo file; those just cause the raw data for the index to be accumulated. To generate an index, you must include the @printindex command at the place in the document where you want the index to appear. Also, as part of the process of creating a printed manual, you must run a program called texindex (see Chapter 18 [Printing Hardcopy], page 131) to sort the raw data to produce a sorted index file. The sorted index file is what will actually be used to print the index.

Texinfo offers six different types of predefined index: the concept index, the function index, etc. (See Section 15.2 [Predefined Indices], page 119.) Each index type has a two-letter name. You may merge indices, or put them into separate sections (See Section 15.4 [Combining Indices], page 121.).

The @printindex command takes a two-letter index name, reads the corresponding sorted index file and formats it appropriately into an index.

The @printindex command does not generate a chapter heading for the index. Consequently, you should precede the command with a suitable section or chapter command (usually @unnumbered) to supply the chapter heading and put the index into the table of contents. Precede the @unnumbered command with an @node line. For example,

     @node Variables Index, Concept Index, Function Index, Top
     @comment    node-name,          next,       previous,  up
     @unnumbered Variable Index

@printindex vr
@node     Concept Index,     , Variables Index, Top
@comment      node-name, next,        previous,  up
@unnumbered Concept Index
@printindex cp
@summarycontents
@contents
@bye

(Readers often prefer that the concept index come last in a book, since that makes it easiest to find.)

4.2 Generating a Table of Contents

The @chapter, @section, and other structuring commands supply the information to make up a table of contents, but they do not cause an actual table to appear in the manual. To do this, you must use the @contents and

@summarycontents commands.
@contents

Generate a table of contents in a printed manual, including all chapters, sections, subsections, etc., as well as appendices and unnumbered chapters. (Headings generated by the @heading series of commands do not appear in the table of contents.) The @contents command should be written on a line by itself.

@shortcontents
@summarycontents
(@summarycontents is a synonym for @shortcontents; the two

commands are exactly the same.) Generate a short or summary table of contents that lists only the chapters (and appendices and unnumbered chapters); sections, subsections and subsubsections are omitted. Write the @shortcontents command on a line by itself immediately before the @contents command. Only large manuals need a short table of contents.

The table of contents commands automatically generate a chapter-like heading at the top of the first table of contents page. Write tables of contents commands at the very end of a Texinfo file, just before the @bye command; the commands should follow the index sections—nothing in the Texinfo file that is written after the table of contents commands will be listed in the table of contents. When you print a manual with a table of contents, the table of contents will be printed last. You need to place those pages in their proper place, after the title page, yourself. (This is the only collating you need to do for a printed manual. The table of contents is printed last because it is generated after the rest of the manual is typeset.)

Here is an example of where to write table of contents commands:

indices…

    @shortcontents
    @contents
    @bye

Since an Info file uses menus instead of tables of contents, the Info formatting commands ignore the @contents and @shortcontents commands.

4.3 @bye File Ending

An @bye command terminates TEX processing and forces out unfinished pages. TEX does not see any of the file following @bye. An @bye command should be on a line by itself.

Optionally, you may follow an @bye line with a local variables list. See Section 18.4 [Using Local Variables and the Compile Command], page 135, for more information.

44 Texinfo 2.0

5 Chapter Structuring

The chapter structuring commands divide a document into a hierarchy of chapters, sections, subsections, and subsubsections. These commands generate large headings; they also provide information for the table of contents of a printed manual (see Section 4.2 [Generating a Table of Contents], page 42).

The chapter structuring commands do not create an Info node structure, so normally you should put an @node command immediately before each chapter structuring command (see Chapter 6 [Nodes and Menus], page 51). The only time you are likely to use the chapter structuring commands without using the node structuring commands is if you are writing a document that contains no cross references and will never be transformed into Info format.

It is unlikely that you will ever write a Texinfo file that is intended only as an Info file and not as a printable document. If you do, you might still use chapter structuring commands to create a heading at the top of each node—but you don’t have to.

5.1 Tree Structure of Sections

A Texinfo file is usually structured like a book with chapters, sections, subsections, and the like. This structure can be visualized as a tree (or rather as an upside-down tree) with the root at the top and the levels corresponding to chapters, sections, subsection, and subsubsections. In Info format, ‘Next’ and ‘Previous’ pointers of a node usually lead to other nodes at the same level; an ‘Up’ pointer usually leads to a node at the level above; and a ‘Menu’ leads to nodes at a level below. (Cross references can point to nodes at any level. See Chapter 7 [Cross References], page 59.)

Here is a diagram that shows a Texinfo file with three chapters, each of which has two sections.

                              top
                               |
             -------------------------------------
            |                  |                  |
         Chapter 1          Chapter 2          Chapter 3
            |                  |                  |
         --------           --------           --------
        |        |         |        |         |        |
     Section  Section   Section  Section   Section  Section
       1.1      1.2       2.1      2.2       3.1      3.2

In a Texinfo file that has this structure, the beginning of Chapter 2 looks like this:

@node    Chapter 2,  Chapter 3, Chapter 1, top
@chapter Chapter 2

The chapter structuring commands are described in the sections that follow; the @node and @menu commands are described in the following chapter (see Chapter 6 [Nodes & Menus], page 51).

5.2 Types of Structuring Command

The chapter structuring commands fall into four groups or series, each of which contains structuring commands corresponding to the hierarchical levels of chapters, sections, subsections, and subsubsections.

The four groups are the @chapter series, the @unnumbered series, the @appendix series, and the @heading series.

Each command produces titles that have a different appearance on the printed page or Info file; and only some of the commands produce titles that are listed in the table of contents of a printed book or manual.

• The @chapter and @appendix series of commands produce numbered or lettered entries both in the body of a printed work and in its table of contents.

• The @unnumbered series of commands produce unnumbered entries both in the body of a printed work and in its table of contents.

• The @heading series of commands produce unnumbered headings that do not appear in a table of contents. The heading commands never start a new page.

• The @majorheading command produces results similar to using the @chapheading command but generates a larger vertical whitespace before the heading.

• When an @setchapternewpage command says to do so, the @chapter, @unnumbered, and @appendix commands start new pages in the printed manual; the @heading commands do not.

Here are the four groups of chapter structuring commands:

No new pages

Numbered  Unnumbered  Lettered and numbered Unnumbered
In contents  In contents  In contents  Not in contents

                                                            @majorheading
     @chapter       @unnumbered          @appendix          @chapheading
     @section       @unnumberedsec       @appendixsec       @heading
     @subsection    @unnumberedsubsec    @appendixsubsec    @subheading
     @subsubsection @unnumberedsubsubsec @appendixsubsubsec @subsubheading

5.3 @chapter

@chapter identifies a chapter in the document. Write the command at the beginning of a line and follow it on the same line by the title of the chapter.

For example, this chapter in this manual is entitled “Chapter Structuring”; the @chapter line looks like this:

    @chapter Chapter Structuring

In TEX, the @chapter command creates a chapter in the document, specifying the chapter title. The chapter is numbered automatically.

In Info, the @chapter command causes the title to appear on a line by itself, with a line of asterisks inserted underneath. Thus, in Info, the above example produces the following output:

    Chapter Structuring
    *******************

5.4 @unnumbered, @appendix

Use the @unnumbered command to create a chapter that appears in a printed manual without chapter numbers of any kind. Use the @appendix command to create an appendix in a printed manual that is labelled by letter instead of by number.

For Info file output, the @unnumbered and @appendix commands are equivalent to @chapter: the title is printed on a line by itself with a line of asterisks underneath. (See Section 5.3 [@chapter], page 47.)

To create an appendix or an unnumbered chapter, write an @appendix or @unnumbered command at the beginning of a line and follow it on the same line by the title, as you would if you were creating a chapter.

5.5 @majorheading, @chapheading

The @majorheading and @chapheading commands put chapter-like headings in the body of a document.

However, neither command causes TEX to produce a numbered heading or an entry in the table of contents; and neither command causes TEX to start a new page in a printed manual.

In TEX, a @majorheading command generates a larger vertical whitespace before the heading than a @chapheading command but is otherwise the same.

For Info file output, the @majorheading and @chapheading commands commands are equivalent to @chapter: the title is printed on a line by itself with a line of asterisks underneath. (See Section 5.3 [@chapter], page 47.)

5.6 @section

In a printed manual, an @section command identifies a numbered section within a chapter. The section title appears in the table of contents. In Info, an @section command provides a title for a segment of text, underlined with ‘=’.

To create a section, write the @section command at the beginning of a line and follow it on the same line by the section title.

Thus,

    @section This is a section

produces

    This is a section
    =================

in Info.

5.7 @unnumberedsec, @appendixsec, @heading
   Use the @unnumberedsec, @appendixsec, and @heading commands for
sections within chapters made by @unnumbered, @appendix, @majorheading

or @chapheading respectively. (See Section 5.6 [@section], page 48.)

@appendixsec
@appendixsection
          @appendixsection is a longer spelling of the @appendixsec

command; it is a synonym for the @appendixsec command.

          Conventionally, the @appendixsec or @appendixsection com-

mand is used only within appendices.

@unnumberedsec

The @unnumberedsec command may be used within an unnumbered chapter or within a regular chapter or appendix to provide an unnumbered section.

@heading You may use the @heading command anywhere you wish for a section-style heading that will not appear in the table of contents.

5.8 The @subsection Command

Subsections are to sections as sections are to chapters. (See Section 5.6 [@section], page 48.) In Info, subsection titles are underlined with ‘-’. For example,

    @subsection This is a subsection

produces

    This is a subsection
    --------------------

In a printed manual, subsections are listed in the table of contents and are numbered three levels deep.

5.9 The @subsection-like Commands

   The @unnumberedsubsec, @appendixsubsec, and @subheading com-

mands are, respectively, the unnumbered, appendix-like, and heading-like equivalents of the @subsection command. (See Section 5.8 [@subsection], page 49.)

In Info, the @subsection-like commands generate a title underlined with hyphens. In a printed manual, an @subheading command produces a heading like that of a subsection except that it is not numbered and does not appear in the table of contents. Similarly, an @unnumberedsubsec command produces an unnumbered heading like that of a subsection and an @appendixsubsec command produces a subsection-liike heading labelled with a letter and numbers

5.10 The ‘subsub’ Commands

The fourth and lowest level sectioning commands in Texinfo are the ‘subsub’ commands. They are:

@subsubsection

Subsubsections are to subsections as subsections are to sections. (See Section 5.8 [@subsection], page 49.) In a printed manual, subsubsection titles appear in the table of contents and are numbered four levels deep.

@unnumberedsubsubsec

Unnumbered subsubsection titles appear in the table of contents of a printed manual, but lack numbers. Otherwise, unnumbered subsubsections are the same as subsubsections. In Info, unnumbered subsubsections look exactly like subsubsections.

@appendixsubsubsec

Conventionally, appendix commands are used only for appendices and are lettered and numbered appropriately in a printed manual. In Info, appendix subsubsections look exactly like subsubsections.

@subsubheading

The @subsubheading command may be used anywhere that you need a small heading that will not appear in the table of contents. In Info, subsubheadings look exactly like subsubsections.

In Info, ‘subsub’ titles are underlined with periods. For example,

@subsubsection This is a subsubsection

produces

This is a subsubsection
.......................

6 Nodes and Menus

Most Texinfo files are organized hierarchically like a book, with chapters, sections, subsections, and subsubsections. Such a hierarchy is tree-like; the chapters are the major limbs from which the sections branch out. In a conventional diagram, however, such a hierarchy is drawn with the “root” at the top and the “leaves” at the bottom—as an upside-down tree. The root node is called the ‘Top’ node, and ‘Up’ pointers carry you closer to the root.

6.1 Node and Menu Illustration

Here is a copy of the diagram shown earlier that illustrates a Texinfo file with three chapters, each of which contains two sections.

                              top
                               |
             -------------------------------------
            |                  |                  |
         Chapter 1          Chapter 2          Chapter 3
            |                  |                  |
         --------           --------           --------
        |        |         |        |         |        |
     Section  Section   Section  Section   Section  Section
       1.1      1.2       2.1      2.2       3.1      3.2

In a Texinfo file that has this organization, you would write the beginning of the node for Chapter 2 like this:

    @node     Chapter 2,  Chapter 3, Chapter 1, top
    @comment  node-name,  next,      previous,  up

To go to Sections 2.1 and 2.2 using Info, you need a menu inside of Chapter 2 that says:

        @menu
        * Sect. 2.1::    Description of this section.
        * Sect. 2.2::
        @end menu

You would locate this menu inside Chapter 2, after the beginning of the chapter and before Section 2.1.

The node for Sect. 2.1 will look like this:

        @node     Sect. 2.1, Sect. 2.2,         ,  Chapter 2
        @comment  node-name, next,      previous,  up

(This node does not have a ‘Previous’ pointer.)

Usually, an @node command and a chapter structuring command are used in sequence, along with indexing commands. (The updating commands require this sequence. See Section 2.3.1 [Updating Requirements], page 20.) Also, you may want to follow the @node line with a comment line that reminds you which pointer is which. For example, the beginning of the node for the chapter on ending a file looks like this:

     @node    Ending a File, Structuring, Beginning a File, Top
     @comment node-name,     next,        previous,         up
     @chapter Ending a Texinfo File
     @cindex Ending a Texinfo file
     @cindex Texinfo file ending
     @cindex File ending

The following two sections describe the @node and @menu commands in detail.

6.2 @node

@node defines the beginning of a new node in the Info output file (See Info file ‘info’, node ‘Top’.). Write the command at the beginning of a line, followed by four arguments, separated by commas, that make up the rest of the line. These arguments are the name of the node, and the names of the ‘Next’, ‘Previous’, and ‘Up’ pointers, in that order. You may insert spaces before each pointer if you wish. The spaces are ignored.

The name of the node identifies the node. The pointers, which enable you to reach other nodes, consist of the names of those nodes.

All the node names for a single Info file must be unique. Duplications confuse the Info movement commands. This means, for example, that if you end each chapter which a summary, you must name every summary node differently. You may, however, duplicate section titles (although this practice may confuse a reader).

Since it is often hard to remember the order for listing the arguments to @node, it is useful to insert a comment line listing the arguments. For example,

    @node    Texinfo Mode, Beginning a File,  Overview, Top
    @comment node-name,    next,              previous, up

defines a node named ‘Texinfo Mode’, whose ‘Next’ pointer is to node ‘Beginning a File’, whose ‘Previous’ pointer is to node ‘Overview’, and whose ‘Up’ pointer is to node ‘Top’. (We have added white space in both the @node line and the @comment line to make the two lines more readable. The spaces are ignored.)

Texinfo mode provides the texinfo-insert-@node command (usually bound to C-c C-c n), which inserts @node and the @comment line and leaves point after @node. (Of course, if you use the Texinfo updating node commands to insert pointers for you, an @comment line will be superfluous. See Chapter 2 [Using Texinfo Mode], page 15.)

The pointer names must be the names of nodes defined elsewhere. In this example, the nodes named ‘Beginning a File’, ‘Overview’ and ‘Top’ must be defined elsewhere in the file with other @node commands. It does not matter whether they are before or after the node that refers to them.

Normally, a node’s ‘Up’ pointer should contain the name of the node whose menu mentions that node. The node’s ‘Next’ pointer should contain the name of the node that follows that node and its ‘Previous’ pointer should contain the name of the node that precedes it in that menu. When a node’s ‘Up’ node is the same as its ‘Previous’ node, both node pointers should name the same node.

In TEX, @node is nearly ignored. It generates nothing visible. Its only function is to identify the name to use for cross references to the chapter or section which follows the @node command and which makes up the body of the node. (Cross references, such as the one following this sentence, are made with @xref and its related commands. See Chapter 7 [Cross References], page 59.)

In general, an @node line is followed immediately by a chapter-structuring

command such as @chapter, @section, @subsection, or @subsubsection.

(See Section 5.2 [Types of Structuring Command], page 46.)

6.2.1 Writing a Node Line

The easiest way to write a node line is to use the Texinfo mode keyboard command C-c C-c n to insert ‘@node’ and a comment line listing the names of the pointers in their proper order. This way you won’t lose track of which arguments are for which pointers. The template is especially useful if you are not familiar with Texinfo. Or else you can just write the node name and use the Texinfo update node commands provided by Texinfo mode. (See Chapter 2 [Using Texinfo Mode], page 15.)

It is important to pick a suitable node name. Node names should be informative, but short. The pointer names must all fit into the rest of the line following the @node command. Long names may not fit into the line. However, so long as an @node line continues on the same line and is not accidentally split onto two lines, it does not matter how long the line is. Sometimes it is better to use long, ugly but informative names rather than short names. By convention, node names are capitalized just as they would be for a section or chapter title.

Caution: Do not use any of the Texinfo @-commands in a node name; these commands confuse Info. Do not use commas within a node name; a comma terminates the node name. Some people insert the names of the ‘Next’, ‘Previous’, and ‘Up’ pointers

at the same time they insert the node’s own name. This is because it is easier to keep track of the node structure as you create a document than it is to sort it out after you have dozens of nodes.

Others wait to insert the ‘Next’, ‘Previous’, and ‘Up’ pointers until they have a nearly final version of the document. This is because they expect to change the organization of the document while they write it and insert or delete sections and move them around. The command texinfo-show-structure can be used to find the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node. (See See Section D.3 [Using texinfo-show-structure], page 178.)

Yet other people use the Texinfo updating commands in Texinfo mode, which insert pointers automatically. (See Chapter 2 [Using Texinfo Mode], page 15.)

However you do it, it is best to name the node itself whenever you write the section so you can easily make cross references to the section. A large number of cross references are an especially important feature of a good Info file.

After you have inserted the node-line, you should immediately write an @-command for the chapter or section and insert its name. Next (and this is important!), put in several index entries. Usually, you will find at least two and often as many as four or five ways of referring to the node in the index. Use them all. This will make it much easier for people to find the node.

The top node of the file (which must be named ‘top’ or ‘Top’) should have as its ‘Up’ and ‘Previous’ nodes the name of a node in another file, where there is a menu that leads to this file. Specify the file name in parentheses. If the file is to be installed directly in the Info directory file, use ‘(dir)’ as the parent of the ‘Top’ node; this is short for ‘(dir)top’, and specifies the ‘Top’ node in the ‘dir’ file, which contains the main menu for Info. For example,

    @node    Top,       Overview, (dir),    (dir)
    @comment node-name, next,     previous, up

See Chapter 20 [Installing an Info File], page 147, for more information about installing an Info file in the ‘info’ directory.

6.3 @menu

Info file nodes can contain menus, which point to other nodes. In Info, the @menu line changes into ‘* Menu:’, which indicates the beginning of a menu in an Info file. Otherwise, the contents are unchanged by Texinfo, except for the processing of any other @-commands within. The entire menu construct has no effect in the printed manual and will not appear there.

By convention, a menu is put at the end of a node. This way, it is easy for someone using Info to find the menu, using the M-> (end-of-buffer) command.

A node that has a menu should not contain much text. If you have a lot of text and a menu, move the text into a new subnode.

6.3.1 Writing a Menu Item

In a menu, every line that begins with a ‘* ’ is a menu item. (Note the space after the asterisk.) A line that does not start with a ‘* ’ will also appear in the menu and can be used as a comment.

A menu item has three parts, only the second of which is required:

1. The menu item name.

2. The name of the node.

3. A description of the item. A menu item looks like this:

    * Item name: Node name.      Description.

Follow the menu item name with a single colon and follow the node name with tab, comma, period, or newline.

In Info, a user can select a node with the m (Info-menu) command. The menu item name is what the user types after the m command.

If the menu item name and the node name are the same, you can write the name immediate after the asterisk and space at the beginning of the line and follow the name with two colons.

For example, write

    * Name::

instead of

    * Name: Name.

You should use the node name for the menu item name whenever possible, since it reduces visual clutter in the menu.

The third part of a menu entry is a short descriptive phrase or sentence. Menu item names and node names are often short; the description explains to the reader what the node is about. The description, which is optional, can spread over two or more lines. A useful description complements the node name rather than repeating it.

6.3.2 A Menu Example

A menu looks like this:

@menu
* Menu item name: Node name.    A short description.
* Node name::                   This form is preferred.
@end menu

This produces:

* menu:
* Menu item name: Node name.    A short description.
* Node name::                   This form is preferred.

Here is an example as you might see it in a Texinfo file:

@menu
Larger Units of Text
* Files::                 All about handling files.
* Multiples: Buffers.     Multiple buffers; editing
                            several files at once.
@end menu

This produces:

* menu:
Larger Units of Text
* Files::                 All about handling files.
* Multiples: Buffers.     Multiple buffers; editing
                            several files at once.

In this example, the menu has two entries. ‘Files’ is both a menu item name and the name of the node referred to by that item. In the other entry, ‘Multiples’ is the item name, and it refers to the node named ‘Buffers’.

Since no file name is specified with either ‘Files’ or ‘Buffers’, they must be the names of nodes in the same Info file. (See Section 6.4 [Referring to Other Info Files], page 57.)

The line ‘Larger Units of Text’ is a comment.

6.4 Referring to Other Info Files

You can refer to nodes in other Info files by writing the file name in parentheses just before the node name. In this case, you should use the three-part menu item format, which saves the reader from having to type the file name.

If you do not list the node name, but only name the file, then Info presumes that you are referring to the ‘Top’ node.

The format looks like this:

    @menu
    * Menu item name: (filename)Nodename.   Three part
                                              format.
    * Menu item name: (filename).           This takes you
                                              to the `Top'
                                              node.
    @end menu

The ‘dir’ top level directory for the Info system has menu entries that take you directly to the ‘Top’ nodes of each Info document. (See Chapter 20 [Installing an Info File], page 147.)

For example,

…

    * Info: (info).         Documentation browsing system.
    * Emacs: (emacs).       The extensible, self-documenting
                            text editor.

…

To refer directly to the ‘Outlining’ and ‘Rebinding’ nodes in the Emacs Manual, you would write a menu similar to the following:

    @menu
    * Outlining: (emacs)Outline Mode. The major mode for
                                      editing outlines.
    * Rebinding: (emacs)Rebinding.    How to redefine the
                                      meaning of a key.
    @end menu

6.5 Describing Command Invocation

Documents that describe programs such as Emacs, GCC, and GAWK that are invoked from a shell each contain a section that describes the command line arguments. Unfortunately, the node names and titles for these sections are all different! This makes it hard to search for the section.

We suggest adopting the convention of naming such sections with a phrase beginning with the word ‘Invoking …’, as in ‘Invoking Emacs’; this way users can find the section easily.

7 Making Cross References

Cross references are used to refer the reader to other parts of the same or different Texinfo files. In Texinfo, nodes are the points to which cross references can refer.

Often, but not always, a printed document should be designed so that it can be read sequentially. People tire of flipping back and forth to find information that should be presented to them as they need it.

However, in any document, some information will be too detailed for the current context, or incidental to it; use cross references to provide access to such information. Also, an on-line help system or a reference manual is not like a novel; few read such documents in sequence from beginning to end. Instead, people look up what they need. For this reason, such creations should contain many cross references to help readers find other information that they may not have read.

In a printed manual, a cross reference creates a page reference, unless it is to another manual altogether, in which case it names that manual.

In Info, a cross reference creates an entry that you can follow using the Info ‘f’ command. (See Info file ‘info’, node ‘Some advanced Info commands’.)

The various cross reference commands use nodes to define cross reference locations. This is evident in Info, in which a cross reference takes you to the specified node. TEX also uses nodes to define cross reference locations but the action is less obvious. When TEX generates a dvi file, it records nodes’ page numbers and uses the page numbers in making references. Thus, if you are writing a manual that will only be printed, and will not be used on-line, you must nonetheless write @node lines to name the places to which you make cross references.

7.1 Different Cross Reference Commands

There are several different cross reference commands: @xref Used to start a sentence in the printed manual saying ‘See …’

or an entry in the Info file saying ‘*Note …’.

@ref Used within or, more often, at the end of a sentence; produces just the reference in the printed manual without a preceding ‘See’.

@pxref Used within parentheses to make a reference that starts with a lower case ‘see’ within the printed manual. (‘p’ is for ‘parenthesis’.)

@inforef Used to make a reference to an Info file for which there is no

printed manual.

(The @cite command is used to make references to books and manuals for which there is no corresponding Info file and, therefore, no node to which to point. See Section 8.1.8 [@cite{reference}], page 74.)

7.2 Parts of a Cross Reference

A cross reference command requires only one argument, which is the name of the node to which it refers. But a cross reference command may contain up to four additional arguments. By using these arguments, you can provide a menu item name for Info, a descriptive phrase for the printed output, the name of a different Info file, and the name of a different printed manual.

Most cross references look like this:

    @xref{Node name}.

which produces

    *Note Node name::.

and See section nnn [Node name], page ppp. Here, however, is an example of a full five-part cross reference:

    @xref{Node name, Item name, Topic, info-file-name,
    A Printed Manual}, for details.

which produces

    *Note Item name: (info-file-name)Node name, for details.

and See section Topic of A Printed Manual, for details. The five arguments for a cross reference are:

1. The node name. (This is required.) This is the node to which the cross reference takes you. In a printed document, the location of the node provides the page reference (for references within the same document).

2. The item name for the Info menu, if it is to be different from the node name. It is usually omitted.

3. A topic description or section name. Often, this is the title of the section. This is used as the name of the reference in the printed manual. If omitted, the node name is used.

4. The name of the Info file in which the reference is located, if it is different from the current file.

5. The name of another printed manual.

Cross references with one, two, three, four, and five arguments are described separately following the description of @xref.

You can write cross reference commands within a paragraph, but note how Info and TEX format the output of each of the various commands: write @xref at the beginning of a sentence; write @pxref only within parentheses, and so on.

7.3 @xref

The @xref command generates a cross reference for the beginning of a sentence. The Info formatting commands convert it into an Info cross reference, which the Info ‘f’ command can use to bring you directly to another node. The TEX typesetting commands convert it into a page reference, or a reference to another book or manual.

Most often, an Info cross reference looks like this:

    *Note Node name::

or like this

    *Note Item name: Node name

In TEX, a cross reference looks like this:

See section section [Node name], page page or like this

    See section section [Topic], page page

The @xref command does not generate a period or comma to end the cross reference in either the Info file or the printed output. You must write that period or comma yourself; otherwise, Info will not recognize the end of the reference. (The @pxref command works differently. See Section 7.5 [@pxref], page 66.)

@xref must refer to an Info node by name. Use @node to define the node (see Section 6.2.1 [Writing a Node], page 53).

@xref is followed by an argument inside braces; but actually the text inside the braces is treated as several arguments, separated by commas. Whitespace after these commas is ignored.

A period or comma must follow the closing brace of an @xref. It is required to terminate the cross reference. This period or comma will appear in the output, both in the Info file and in the printed manual.

A cross reference requires only the name of a node; but it may contain up to four additional arguments. Each of these variations produces a cross reference that looks somewhat different.

7.3.1 @xref with One Argument

The simplest form of @xref takes one argument, the name of another node in the same Info file. For example,

    @xref{Node name}.

produces

    *Note Node name::.

and

See section nnn [Node name], page ppp. (Note that in the preceding example the closing brace is followed by a period.)

You can write a clause after the cross reference, like this:

    @xref{Node name}, for more info.

which produces:

    *Note Node name::, for more info.

See section nnn [Node name], page ppp, for more info. (Note that in the preceding example the closing brace is followed by a comma, and then by the clause.)

7.3.2 @xref with Two Arguments

With two arguments, the second one is used as the name of the Info cross reference, while the first argument is still the node that the cross reference points to:

    @xref{Node name, Item name}.

produces

    *Note Item name: Node name.

and

See section nnn [Node name], page ppp. (Note that in the preceding example the closing brace is followed by a period.)

You can write a clause after the cross reference, like this:

    @xref{Node name, Item name}, for more info.

produces

    *Note Item name: Node name, for more info.

and

See section nnn [Node name], page ppp, for more info. (Note that in the preceding example the closing brace is followed by a comma, and then by the clause.)

7.3.3 @xref with Three Arguments

A third argument replaces the node name in the TEX output. It should state the topic discussed by the section being referenced, or be the name of the section. Often, you will want to use initial upper case letters so it will be easier to read when the reference is printed. Use a third argument when the node name is unsuitable because of syntax or meaning.

Remember that a comma or period must follow the closing brace of an @xref command to terminate the cross reference. In the following examples, a clause follows a terminating comma. For example,

    @xref{Node name, Item name, Topic}, for details.

produces

    *Note Item name: Node name, for details.

and See section nnn [Topic], page ppp, for details. If a third argument is given and the second one is empty, then the third

argument serves both. (Note how two commas, side by side, mark the empty second argument.)

    @xref{Node name, , Topic}, for details.

produces

    *Note Topic: Node name, for details.

and See section nnn [Topic], page ppp, for details.

7.3.4 @xref with Four Arguments

A fourth argument specifies the name of the Info file in which the referenced node is located, assuming it is not the one in which the reference appears. @xref with only four arguments is used when the reference is not within one Info file, but is within a single printed manual—when multiple Texinfo files are incorporated into the same TEX run but make separate Info files. (This is seldom the case and usually you will use five arguments if you are making a reference that is outside the current Info file.)

Remember that a comma or period must follow the closing brace of an @xref command to terminate the cross reference. In the following examples, a clause follows a terminating comma. For example,

    @xref{Node name, Item name, Topic, info-file-name},
    for details.

produces

    *Note Item name: (info-file-name)Node name, for details.

and See section nnn [Topic], page ppp, for details.

7.3.5 @xref with Five Arguments

A fifth argument is used when you are making a reference to another Info file which is also part of another printed manual. Write the title of the manual in this slot. Since TEX does not know the page numbers in the other manual, the printed reference will not have a page number.

Remember that a comma or period must follow the closing brace of an @xref command to terminate the cross reference. In the following examples, a clause follows a terminating comma. For example,

    @xref{Node name, Item name, Topic, info-file-name,
    A Printed Manual}, for details.

produces

    *Note Item name: (info-file-name)Node name, for details.

and

See section Topic of A Printed Manual, for details. The name of the printed manual will be typeset in italics.

Often, you will leave out the second argument when you use the long version of @xref. In this case, the third argument, the topic description, will be used as the item name in Info. For example,

    @xref{Node name, , Topic, info-file-name,
    A Printed Manual}, for details.

produces

    *Note Topic: (info-file-name)Node name, for details.

and See section Topic of A Printed Manual, for details. In a cross reference, you must always name a node. This means that in

order to refer to a whole manual, you must identify the ‘Top’ node by writing it as the first argument to the @xref command. (This is different from the way you write a menu entry. See Section 6.4 [Referring to Other Info Files], page 57.) At the same time, to provide a meaningful section topic or title in the printed cross reference (instead of the word ‘Top’), you must write an appropriate entry for the third argument to the @xref command.

Thus, to make a cross reference to The GNU Make Manual, write:

    @xref{Top, , Overview, make, The GNU Make Manual}.

which produces

    *Note Overview: (make)Top.

and

See section Overview of The GNU Make Manual. In this example, ‘Top’ is the name of the node, and ‘Overview’ is the name of the first section of the manual.

7.4 @ref

@ref is nearly the same as @xref except that it does not generate a ‘See’ in the printed output, just the reference itself. This makes it useful as the last part of a sentence. For example,

    For more information, see @ref{Node name, ,
    Topic Description}.

produces

    For more information, see *Note Topic
    Description: Node name.

and For more information, see section nnn [Topic Description], page ppp.

The @ref command sometimes leads writers to express themselves in a manner that is suitable for a printed manual but looks awkward in the Info format. Bear in mind that your audience will be using both the printed and the Info format. For example,

    Sea surges are described in @ref{Hurricanes}.

produces

Sea surges are described in section nnn [Hurricanes]. in a printed document, but

    Sea surges are described in *Note Hurricanes::.

in Info.

Caution: You must write a period or comma immediately after an @ref command with two or more arguments. Otherwise, Info will not find the end of the cross reference entry and attempts to follow the cross reference will fail. As a general rule, you should write a period or comma after every @ref command. This looks best in both the printed and the Info output.

7.5 @pxref

The parenthetical reference command, @pxref, is nearly the same as @xref, but you use it only inside parentheses and you do not type a comma or period after the command’s closing brace. The command differs from @xref in two ways:

1. TEX typesets the reference for the printed manual with a lower case ‘see’ rather than an upper case ‘See’.

2. The Info formatting commands automatically end the reference with a closing colon or period. Because one type of formatting automatically inserts closing punctuation

and the other does not, you should use @pxref only inside parentheses as part of another sentence. Also, you yourself should not insert punctuation after the reference, as you do with @xref.

The parenthetical reference command is designed so that the output looks right and works right both in printed output and in an Info file. In a printed manual, a closing colon or period should not follow a cross reference within parentheses; such punctuation is wrong. But in an Info file, a closing colon or period must follow the cross reference so Info can recognize its end. @pxref spares you the need to use complicated methods to put a terminator into one form of the output and not the other.

Don’t try to use @pxref as a clause in a sentence. It will look bad in either the Info file, the printed output, or both. Use it only as a parenthetical reference.

With one argument, a parenthetical cross reference looks like this:

    . . . large storms (@pxref{Hurricanes}) may cause flooding

…

which produces

    . . . large storms (*Note Hurricanes::) may cause flooding . . .

and … large storms (see section nnn [Hurricanes], page ppp) may cause flooding … With two arguments, a parenthetical cross reference looks like this:

    . . . (@pxref{Node name, Item name}) . . .

which produces

    . . . (*Note Item name: Node name.) . . .

and … (see section nnn [Node name], page ppp) … @pxref can be used with up to five arguments just like @xref (see Sec-

tion 7.3 [@xref], page 61).

7.6 @inforef

@inforef is used for cross references to Info files for which there are no printed manuals. Even in a printed manual, @inforef generates a reference directing the user to look in an Info file.

The command takes either two or three arguments, in the following order:

1. The node name.

2. The item name (optional).

3. The Info file name. Separate the arguments with commas, as with @xref. Also, you must terminate the reference with a comma or period after the ‘}’, as you do with

@xref.

Thus,

    @inforef{Node name, Item name, info-file-name},
    for more information.

produces

*Note Item name: (info-file-name)Node name,
for more information.

and See Info file ‘info-file-name’, node ‘Node name’, for more information.

Similarly,

@inforef{Node name, , info-file-name},
for more information.

produces

*Note (info-file-name)Node name::,
for more information.

and See Info file ‘info-file-name’, node ‘Node name’, for more information. The converse of @inforef is @cite, which is used to refer to printed

works for which no Info form exists. See Section 8.1.8 [@cite], page 74.

8 Marking Words and Phrases

In Texinfo, you can mark words and phrases in a variety of ways. These ways specify, for example, whether a word or phrase is a defining occurrence, a metasyntactic variable, or a symbol used in a program. Also, you can emphasize text.

8.1 Indicating Definitions, Commands, etc.

Texinfo has commands for indicating just what kind of object a piece of text refers to. Metasyntactic variables, for example, are marked by @var and code by @code. Texinfo uses this information to determine how to highlight the text. Since the pieces of text are labelled by commands that tell what kind of object they are, it is easy to change the way Texinfo formats such text. (Texinfo is an intentional formatting language rather than a typesetting formatting language.)

For example, code is usually illustrated in a typewriter font, but it would be easy to change the way Texinfo highlights code to use another font. This change would not effect how keystroke examples are highlighted. If straight typesetting commands were used in the body of the file and you wanted to make a change, you would have to check every single occurrence to make sure that you were changing code and not something else that should not be changed.

The highlighting commands can be used to generate useful information from the file, such as lists of functions or file names. It is possible, for example, to write a program in Emacs Lisp (or a keyboard macro) to insert an index entry after every paragraph that contains words or phrases marked by a specified command. You could do this to construct an index of functions if you had not already made the entries.

The commands serve a variety of purposes: @code{sample-code}

Indicate text that is a literal example of a piece of a program.

@kbd{keyboard-characters}

Indicate keyboard input.

@key{key-name}

Use for the conventional name for a key on a keyboard.

@samp{text}

Indicate text that is a literal example of a sequence of characters.

@var{metasyntactic-variable} Indicate a metasyntactic variable.

@file{file-name}

Indicate the name of a file.

@dfn{term}

Use for the introductory or defining use of a term.

@cite{reference}

Indicate the name of a book.

@ctrl{ctrl-char} Use for an ASCII control character.

8.1.1 @code{sample-code}

Use the @code command to indicate text that is a piece of a program which consists of entire syntactic tokens. Enclose the text in braces.

Thus, you should use @code for an expression in a program, for the name of a variable or function used in a program, or for a keyword. Also, you should use @code for the name of a program, such as diff, that is a name used in the machine. (You should write the name of a program in the ordinary text font if you regard it as a new English word, such as ‘Emacs’ or ‘Bison’.)

Do not use the @code command for a string of characters shorter than a syntactic token. In particular, you should not use the @code command when writing about the characters used in a token; do not, for example, use @code when you are explaining what letters or printable symbols can be used in the names of functions. (Use @samp.) Also, you should not use @code to mark text that is considered input to programs unless the input is written in a language that is like a programming language. For example, you should not use @code for the single character commands of GNU Emacs (use @kbd instead) although you may use @code for the names of the Emacs Lisp functions that the keyboard commands invoke.

Do use @code for command names in command languages that resemble programming languages, such as Texinfo or the shell. Note, however, that you should not use @code for options such as ‘-c’ when such options stand alone. (Use @samp.) Also, an entire shell command often looks better if written using ‘@samp’ rather than ‘@code’.

It is incorrect to alter the case of a word inside an @code command when it appears at the beginning of a sentence. This is a particularly insidious error if the language being documented is case sensitive. If the command is printf, then Printf is a misspelling. If you do not like having such a command with lower case at the beginning of a sentence, rearrange the sentence.

In the printed manual, @code causes TEX to typeset the argument in a typewriter face. In the Info file, it causes the Info formatting commands to use ‘…’ quotation. For example:

    Use @code{diff} to compare two files.

produces this in the printed manual:

Use diff to compare two files. and this in Info file:

    Use `diff' to compare two files.

8.1.2 @kbd{keyboard-characters}

Use the @kbd command for characters of input to be typed by users. For example, to refer to the characters M-a, write

    @kbd{M-a}

and to refer to the characters M-x shell, write

    @kbd{M-x shell}

The @kbd command has the same effect as @code in Info, but may produce a different font in a printed manual.

You can embed another @-command inside the braces of an @kbd command. Here, for example, is the way to describe a command that would be described more verbosely as “press an ‘r’ and then press the RET key”:

    @kbd{r @key{RET}}

This produces: r RET

You also use the @kbd command if you are spelling out the letters you type; for example:

    To give the @code{logout} command,
    type the characters @kbd{1 o g o u t @key{RET}}.

This produces To give the logout command, type the characters l o g o u t RET. (Also, this example shows that you can add spaces for clarity. If you

really want to mention a space character as one of the characters of input, write @key{SPC} for it.)

8.1.3 @key{key-name}

Use the @key command for the conventional name for a key on a keyboard, as in

    @key{RET}

You can use the @key command within the argument of an @kbd command when the sequence of characters to be typed includes one or more keys that are described by name.

For example, to produce C-x ESC you would type:

    @kbd{C-x @key{ESC}}

The recommended names to use for keys are in upper case and are SPC Space RET Return LFD Linefeed

TAB  Tab

BS Backspace ESC Escape DEL Delete SFT Shift CTL Control

META  Meta

There are subtleties to handling words like ‘meta’ or ‘ctrl’ that are names of shift keys. When mentioning a character in which the shift key is used, such as Meta-a, use the @kbd command alone without the @key command, but when you are referring to the shift key in isolation, use the @key command. For example, write ‘@kbd{Meta-a}’ to produce Meta-a and ‘@key{META}’ to produce META. This is because Meta-a refers to keys that you press on a keyboard, but META refers to a key without implying that you press it.

8.1.4 @samp{text}

Use the @samp command to indicate text that is a literal example of a sequence of characters in a file, string, pattern, etc. Enclose the text in braces. The argument appears within ‘…’ quotation in both the Info file and the printed manual; in addition, it is printed in a fixed-width font.

    To match @samp{foo} at the end of the line,
    use the regexp @samp{foo$}.

produces To match ‘foo’ at the end of the line, use the regexp ‘foo$’.

Any time you are referring to single characters, you should use @samp unless @kbd is more appropriate. Basically, @samp is a catchall for whatever is not covered by @code, @kbd, or @key.

Never include the punctuation marks that are part of the English text that surrounds the strings you are specifying within the braces. In the following sentence, for example, the commas and period are outside of the braces:

    In English, the vowels are @samp{a}, @samp{e},
    @samp{i}, @samp{o}, @samp{u}, and sometimes
    @samp{y}.

This produces:

In English, the vowels are ‘a’, ‘e’, ‘i’, ‘o’, ‘u’, and sometimes ‘y’.

8.1.5 @var{metasyntactic-variable}

Use the @var command to indicate metasyntactic variables. A metasyntactic variable is something that stands for another piece of text. For example, you should use a metasyntactic variable in the documentation of a function to describe the arguments that are passed to that function.

Do not use @var for the names of particular variables in programming languages. These are specific names from a program, so @code is correct for them. For example, the Lisp variable texinfo-tex-command is not a metasyntactic variable; it is properly formatted using @code.

The effect of @var in the Info file is to upcase the argument; in the printed manual, to italicize it. For example:

    To delete file @var{filename},
    type @code{rm @var{filename}}.

produces To delete file filename, type rm filename. In some documentation styles, metasyntactic variables are shown with

angle brackets, for example:

    . . ., type rm <filename>

However, that is not the style we use in Texinfo. (You could, of course, modify the sources to TEX and the Info formatting commands to output the <…> format if you wished.)

8.1.6 @file{file-name}

Use the @file command to indicate text that is the name of a file, buffer, or directory. You can also use the command for filename suffixes. Currently, @file is equivalent to @samp in its effects on the output. For example,

    The @file{.el} files are in
    the @file{/usr/local/emacs/lisp} directory.

produces The ‘.el’ files are in the ‘/usr/local/emacs/lisp’ directory.

8.1.7 @dfn{term}

Use the @dfn command to identify the introductory or defining use of a technical term. Use the command only in passages whose purpose is to introduce a term which will be used again or which the reader ought to know. Mere passing mention of a term for the first time doesn’t deserve @dfn. The command generates italics in the printed manual, and double quotation marks in the Info file. For example:

    Getting rid of a file is called @dfn{deleting} it.

produces Getting rid of a file is called deleting it. As a general rule, a sentence containing the defining occurrence of a

term should be a definition of the term. The sentence does not have to say explicitly that it is a definition, but it should contain the information of a definition.

8.1.8 @cite{reference}

Use the @cite command for the name of a book that lacks a companion Info file. The command produces italics in the printed manual, and quotation marks in the Info file.

(If a book is written in Texinfo, it is better to use a cross reference command since you can easily follow such a reference in Info. See Section 7.3 [@xref], page 61.)

8.1.9 @ctrl{ctrl-char}

The @ctrl command is seldom used. It describes an ASCII control character by inserting the actual character into the Info file. Usually, in Texinfo, you talk about a control character by describing it with @kbd: thus, ‘@kbd{C-

a}’ for C-a.

Use @kbd in this way when talking about a control character that is typed on the keyboard by the user. When talking about a control character appearing in a file or a string, don’t use @kbd since the control character is not typed. Also, don’t use ‘C-’ but spell out control-, as in ‘control-a’, to make it easier for a reader to understand.

@ctrl is an idea from the beginnings of Texinfo which may not really fit in to the scheme of things. But there may be times when you may want to use the command. The pattern is @ctrl{ch}, where ch is an ASCII character whose control-equivalent is wanted. For example, to specify ‘control-f’, you would enter

    @ctrl{f}

which produces

    ^f

In the Info file, this generates the specified control character, output literally into the file. This is done so a user can copy the specified control character (along with whatever else he or she wants) into another Emacs buffer and use it. Since the ‘control-h’,‘control-i’, and ‘control-j’ characters are formatting characters, they should not be indicated with @ctrl.

In a printed manual, @ctrl generates text to describe or identify that control character: an uparrow followed by the character ch.

8.2 Emphasizing Text

Usually, Texinfo changes the font to mark words in the text according to what category the words belong to. The @code command, for example, does this. Most often, this is the best way to mark words. However, sometimes you will want to emphasize text without indicating a category. Texinfo has two ways to do this: commands that tell Texinfo to emphasize the text but leave the method to the program, and commands that specify the method to use. The first method is generally the best because it makes it possible to change the style of a document without have to re-edit it line by line.

8.2.1 @emph{text} and @strong{text}

The @emph and @strong commands are for emphasis; @strong is stronger. In printed output, @emph produces italics and @strong produces bold.

For example,

    @quotation
    @strong{Caution:} @code{rm * .??*} removes @emph{all}

    files in the directory.
    @end quotation

produces the following in printed output:

Caution: rm * .??* removes all files in the directory. and the following in Info:

         *Caution*: `rm * .??*' removes *all*
         files in the directory.

The @strong command is seldom used except to mark what is, in effect, a typographical element, such as the word ‘Caution’ in the preceding example.

In the Info file, both @emph and @strong put asterisks around the text.

8.2.2 @sc{text}: The Small Caps Font

Use the ‘@sc’ command to set text in the printed output in a small caps font and set text in the Info file in upper case letters.

Write the text between braces in lower case, like this:

    The @sc{acm} and @sc{ieee} are technical societies.

This produces:

The acm and ieee are technical societies.

TEX typesets the small caps font in a manner that prevents the letters from ‘jumping out at you on the page’. This makes small caps text easier to read than text in all upper case.

TEX typesets any upper case letters in the small caps fonts in FULL-SIZE CAPITALS. Use them sparingly.

The Info formatting commands set all small caps text in upper case. You may also use the small caps font for a jargon word such as ato (a

nasa word meaning ‘abort to orbit’).

There are subtleties to using the small caps font with a jargon word such as cdr, a word used in Lisp programming. In this case, you should use the small caps font when the word refers to the second and subsequent elements of a list (the cdr of the list), but you should use ‘@code’ when the word refers to the Lisp function of the same spelling.

8.2.3 Fonts for Printing, Not Info

Texinfo provides four font commands that specify font changes in the printed manual but have no effect in the Info file. @i requests italic font (in some versions of TEX, a slanted font is used), @b requests bold face, @t requests the fixed-width font used by @code, and @r requests a roman font, which is the usual font in which text is printed. All four commands apply to an argument that follows, surrounded by braces.

Only the @r command has much use: in example programs, you can use the @r command to convert code comments from the fixed-width font to a roman font. This looks better in printed output.

For example,

    @lisp
    (+ 2 2)    ; @r{Add two plus two.}
    @end lisp

produces (+ 2 2) ; Add two plus two. If possible, you should avoid using the other three font commands. If you

need to use one, it probably indicates a gap in the Texinfo language.

78 Texinfo 2.0

9 Special Insertions

Texinfo provides several commands for for inserting single characters that have special meaning in Texinfo, such as braces, and for inserting special graphic symbols that do not correspond to characters, such as dots and bullets.

These are:

• Braces, ‘@’ and periods.

• Dots and bullets.

• The TEX logo and the copyright symbol.

• A minus sign.

9.1 Inserting ‘@’, Braces, and Periods

‘@’ and curly braces are special characters in Texinfo. To insert these characters into text, you must put an ‘@’ in front of these characters to prefent Texinfo from misinterpreting them.

Periods are also special. Depending on whether the period is inside of or at the end of a sentence, less or more space is inserted after a period in a typeset manual. Since it is not always possible for Texinfo to determine when a period ends a sentence and when it is used in an abbreviation, special commands are needed in some circumstances. (Usually, Texinfo figures out how to handle periods, so you don’t have to use the special commands; you just enter a period as you would if you were using a typewriter, which means you put two spaces after the period, question mark, or exclamation mark that ends a sentence.)

Do not put braces after any of these commands; they are not necessary.

9.1.1 Inserting ‘@’—@@

@@ stands for a single ‘@’ in either printed or Info output.

Do not put braces after an @@ command.

9.1.2 Inserting ‘{’ and ‘}’—@{ and @}

@{ stands for a single ‘{’ in either printed or Info output.

@} stands for a single ‘}’ in either printed or Info output.

Do not put braces after either an @{ or an @} command.

9.1.3 Spacing After Colons and Periods

Use the @: command after a period, question mark, exclamation mark, or colon that should not have extra space inserted after it by TEX. For example, use @: after periods that end abbreviations which are not at the ends of sentences. @: has no effect on the Info file output.

For example:

    The U.S.A.@: is a continental nation.

produces The U.S.A. is a continental nation. Use @. instead of a period at the end of a sentence that ends with a single

captial letter. Otherwise, TEX will think the letter is an abbreviation and will not insert the correct end-of-sentence spacing. Here is an example:

    Give it to X. and to Y@.  Give it to Z@.
    Give it to X. and to Y.  Give it to Z@.

If you look carefully at this printed output, you will see a little more whitespace after the Y in the first line than the Y in the second line.

Give it to X. and to Y. Give it to Z. Give it to X. and to Y. Give it to Z. In the Info file output, @. is equivalent to a simple ‘.’. The meanings of @: and @. in Texinfo are designed to work well with the

Emacs sentence motion commands. This means they are different from their meanings in some other formatting systems that use @-commands.

Do not put braces after either an @: or an @. command.

9.2 Inserting Ellipsis, Dots, and Bullets

An ellipsis (a line of dots) is typeset unlike a string of periods, so a special command is used in Texinfo. The @bullet command is special, too. Each of these commands is followed by a pair of braces, ‘{}’, without any whitespace between the name of the command and the braces.

9.2.1 @dots{}

Use the @dots{} command to generate an ellipsis, which is three dots in a row, appropriately spaced, like this: ‘…’. Do not simply write three periods in the input file; that would work for the Info file output, but would produce the wrong amount of space between the periods in the printed manual.

Here is an ellipsis: …

Here are three periods in a row: ...

In printed output, the three periods in a row are closer together than the dots in the ellipsis.

9.2.2 @bullet{}

Use the @bullet{} command to generate a large round dot, or the closest possible thing to one.

Here is a bullet: •

When you use @bullet to specify the mark beginning each entry in an itemized list, you do not need to type the braces.

9.3 Inserting TEX and c_{◯}

The logo TEX is typeset in a special fashion and it needs an @-command, as does the command for inserting the copyright symbol. Each of these commands is followed by a pair of braces, ‘{}’, without any whitespace between the name of the command and the braces.

9.3.1 @TeX{}

Use the @TeX{} command to generate ‘TEX’. In a printed manual, this is a special logo that is different from three ordinary letters. The @TeX{} command is unique among Texinfo commands in that the T and the X are in upper case.

9.3.2 @copyright{}

Use the @copyright\{\} command to generate ‘\begin{matrix}c \\ ◯ \end{matrix}’. In a printed manual, this is a ‘c’ inside a circle, and in Info, this is ‘(C)’.

9.4 @minus{}: Inserting a Minus Sign

Use the @minus{} command to generate a minus sign. In a fixed-width font, this is a single hyphen, but in a proportional font, the symbol is the customary length for a minus sign—a little longer than a hyphen.

You can compare the two forms:

‘−’ is a minus sign generated with ‘@minus{}’,

‘-’ is a hyphen generated with a single hyphen. In the fixed-width font used by Info, @minus{} is the same as a hyphen.

You should not use @minus{} inside of @code or @example environments because the width distinction is not made in the fixed-width font they use.

When you use @minus to specify the mark beginning each entry in an itemized list, you do not need to type the braces.

10 Quotations and Examples

Quotations and examples are blocks of text consisting of one or more whole paragraphs that are set off from the bulk of the text and treated differently. They are usually indented.

In Texinfo, you always begin a quotation or example by writing an @- command at the beginning of a line by itself, and end it by writing an @end command that is also at the beginning of a line by itself. For instance, you begin an example by writing @example by itself at the beginning of a line and end the example by writing @end example on a line by at itself, at the beginning of that line. The text between the beginning and end command is enclosed in the beginning command’s environment.

10.1 The Various Environments

There are a variety of commands for quotations and examples:

@quotation

Used to indicate text that is quoted. The text is filled, indented, and printed in a roman font by default.

@example Used to illustrate code, commands, and the like. The text is

printed in a fixed-width font without filling.

@lisp Used to illustrate Lisp code. The text is printed in a fixed-width

font without filling.

@smallexample

Used to illustrate code, commands, and the like. The text is printed in a fixed-width font without filling; for use with the

          @smallbook command.
@smalllisp

Used to illustrate Lisp code. The text is printed in a fixed-width font without filling; for use with the @smallbook command.

@display Used for illustrative text. The text is indented but not filled,

and no font is specified (so, by default, the font is roman).

@format Used for illustrative text. The text is not indented and not filled

and no font is specified (so, by default, the font is roman). The @exdent command is used within the above constructs to undo the

indentation of a line. The @flushleft and @flushright commands are used to line up the left or right margins of unfilled text.

The @noindent command may be used after one of the above constructs to prevent the following text from being indented as a new paragraph.

10.2 @quotation

The text of a quotation is processed normally except that

• The margins are narrower, so the whole of the quotation is indented.

• The first lines of paragraphs are indented no more than the other lines.

• In the printed output, interline spacing and interparagraph spacing are reduced. This is an example of text written between an @quotation command and an @end quotation command. A @quotation command is most often used to indicate text that is excerpted from another (real or hypothetical) printed work. Write an @quotation command at the beginning of and as the only text

on a line by itself. This line will disappear from the output. Mark the end of the quotation with a line beginning with and containing only @end quotation. The @end quotation line will likewise disappear from the output. Thus, the input

@quotation
This is
a foo.
@end quotation

produces This is a foo.

10.3 @example

The @example command is used to indicate an example that is not part of the running text, such as computer input or output.

This is an example of text written between an
@example command and an @end example
command.  The text is indented but not filled.
In the printed manual, the text is typeset in a
fixed-width font, and extra spaces and blank lines are
significant.  In the Info file, an analogous result is
obtained by indenting each line with five extra spaces.

Write an @example command at the beginning of and as the only text on a line by itself. This line will disappear from the output. Mark the end of the example with a line beginning with and containing only @end example. The @end example will likewise disappear from the output. For example:

    @example
    mv foo bar
    @end example

produces

    mv foo bar

Since the lines containing @example and @end example will disappear, you should put a blank line before the @example and another blank line after the @end example. (Remember that blank lines between the beginning @example and the ending @end example will appear in the output.)

Note: Don’t use tabs in lines of an example (or anywhere else in Texinfo, for that matter)! TEX has trouble with tabs: it treats them like single spaces, and that is not what they look like. This is a problem with TEX. (If necessary, in Emacs, you can use M-x untabify to convert tabs in a region to multiple spaces.)

When you use an @example environment to describe a C function’s calling conventions, use the ansii C syntax, like this:

    void dld_init (char *@var{path});

And in the subsequent discussion, refer to the argument values by writing the same argument names within @var.

Avoid the obsolete style that looks like this:

    #include <dld.h>
    dld_init (path)
    char *path;

Also, it is best to avoid writing #include above the declaration just to indicate that the function is declared in a header file. The practice may give the misimpression that the #include belongs near the declaration of the function. Either state explicitly which header file holds the declaration or, better yet, name the header file used for a group of functions at the beginning of the section that describes the functions.

Examples are often, as it were, “in the middle” of a paragraph, and the text continues after an example should not be indented. The @noindent command prevents a piece of text from being indented as if it were a new paragraph.

(The @code command is used for examples of code that is embedded within sentences, not set off from preceding and following text. See Section 8.1.1 [@code{sample-code}], page 70.)

10.4 @noindent

If you have text following an @example or other similar inclusion that reads as a continuation of the text before the @example, it is good to prevent this text from being indented as a new paragraph. To accomplish this, write @noindent at the beginning of a line by itself preceding the continuation text. For example,

@example
This is an example
@end example
@noindent
This line will not be indented.  As you can see, the
beginning of the line is fully flush left with the line
that follows after it.  (This whole example is within an
@display environment.)

produces

    This is an example

This line will not be indented. As you can see, the beginning of the line is fully flush left with the line that follows after it. (This whole example is within an @display environment.) To adjust the number of blank lines properly in the Info file output,

remember that the line containing @noindent does not generate a blank line, and neither does the @end example line.

In the Texinfo source file for this documentation, each of the lines that says ‘produces’ is preceded by a line containing @noindent.

Do not put braces after an @noindent command; they are not necessary, since @noindent is a command used outside of paragraphs (see Appendix G[Command Syntax], page 191).

10.5 @lisp

The @lisp command is used for Lisp code. It is synonymous with the

@example command.
This is an example of text written between an
@lisp command and an @end lisp command.

Use @lisp instead of @example so as to preserve information regarding the nature of the example. This is useful, for example, if you write a function that evaluates only and all the Lisp code in a Texinfo file. Then you can use the Texinfo file as a Lisp library. ^{1}

Mark the end of an @lisp environment with @end lisp on a line by itself.

10.6 @smallexample and @smalllisp

In addition to the regular @example and @lisp commands, Texinfo has two other “example-style” commands. These are the @smallexample and @smalllisp commands. Both these commands are designed for use with the @smallbook command that causes TEX to produce a printed manual in a 7 by 9.25 inch format rather than the regular 8.5 by 11 inch format.

In TEX, the @smallexample and @smalllisp commands are defined the same as the @example and @lisp commands for 8.5 by 11 inch output and produce similar looking output. But when TEX is preparing a manual in a 7 by 9.25 inch format, the text bounded by @smallexample or @smalllisp commands is typeset in a smaller font than that used for the ordinary examples.

In Info, the @smallexample and @smalllisp commands are equivalent to the @example and @lisp commands.

Mark the end of an @smallexample or @smalllisp environment with

@end smallexample or @end smalllisp, respectively.
     This is an example of text written in an @smallexample
     environment.

The @smallexample and @smalllisp commands make it easier to prepare smaller format manuals without forcing you to edit examples by hand to fit them onto narrower pages.

See Section 18.8 [Printing “Small” Books], page 137, for more information about the @smallbook command.

10.7 @display

The @display command is like @example except that, in the printed manual, @display does not select the fixed-width font. In fact, it does not

1 It would be straightforward to extend Texinfo to include an environment for C code, for Fortran, or for some other language, if someone wished. specify the font at all, so that the text appears in the same font it would have appeared in without the @display command.

This is an example of text written between an @display command and an @end display command. The @display command indents the text, but does not fill it.

10.8 @format

The @format command is similar to @example except that, in the printed manual, @format does not select the fixed-width font and does not narrow the margins. This is an example of text written between an @format command and an @end format command. The @format command does not fill the text.

10.9 @exdent: Undoing a Line’s Indentation

The @exdent command removes any indentation a line might have. The command is written at the beginning of a line and applies only to the text that follows the command that is on the same line. Don’t use braces around the text. In the printed manual, the text on the @exdent line is printed in the roman font.

@exdent is usually used within examples. Thus,

    @example
    This line is within an @@example environment.
    @exdent  This line is exdented.
    This line follows the exdented line.
    @end example

produces

    This line is within an @example environment.

This line is exdented.

    This line follows the exdented line.

As a practical matter, the @exdent command is rarely used. Usually, you un-indent text by ending the indented region and returning the page to its normal width.

10.10 @flushleft and @flushright

The @flushleft and @flushright commands line up the left or right ends of lines on the left and right margins of a page, but do not fill the text. The commands are written on lines of their own, without braces. The environments are ended by @end flushleft and @end flushright commands on lines of their own.

For example,

    @flushleft
    This text is
    written flushleft.
    @end flushleft

produces This text is written flushleft.

Flushright produces the type of indentation often used in the return address of letters.

    @flushright
    Here is an example of text written
    flushright.  The @code{@flushright} command
    right justifies every line but leaves the
    left end ragged.
    @end flushright

produces

Here is an example of text written flushright. The @flushright command right justifies every line but leaves the

left end ragged.

90 Texinfo 2.0

11 Special Glyphs for Examples

In Texinfo, code is often illustrated in examples that are delimited by

@example and @end example, or by @lisp and @end lisp. In such examples,

you can indicate the results of evaluation or an expansion using ‘⇒’ or ‘↦’. Likewise, there are special symbols to indicate printed output, an error message, equivalence of expressions, and the location of point.

The special glyph commands do not have to be used within an example, but most often they are. Every special glyph command is followed by a pair of left- and right-hand braces. ⇒ @result{} points to the result of an expression. ↦ @expansion{} shows the results of a macro expansion. ⊣ @print{} indicates printed output. error @error{} indicates that the following text is an error message. ≡ @equiv{} indicates the exact equivalence of two forms. ⋆ @point{} shows the location of point.

11.1 ^{⇒}: Indicating Evaluation

Use the @result{} command to indicate the result of evaluating an expression.

The @result{} command is displayed as ‘=>’ in Info and as ‘⇒’ in the printed output.

Thus, the following,

    (cdr '(1 2 3))
         ⇒ (2 3)

may be read as “(cdr '(1 2 3)) evaluates to (2 3)”.

11.2 ^{↦}: Indicating an Expansion

When an expression is a macro call, it expands into a new expression. You can indicate the result of the expansion with the @expansion{} command.

The @expansion{} command is displayed as ‘==>’ in Info and as ‘↦’ in the printed output.

For example, the following

    @lisp
    (third '(a b c))
         @expansion{} (car (cdr (cdr '(a b c))))
         @result{} c
    @end lisp

produces

    (third '(a b c))
         ↦ (car (cdr (cdr '(a b c))))
         ⇒ c

which may be read as:

    (third '(a b c)) expands to (car (cdr (cdr '(a b c)))); the re-

sult of evaluating the expression is c. (Often, as in this case, an example looks better if the @expansion{} and @result{} commands are indented five spaces.)

11.3 _{⊣}: Indicating Printed Output

Sometimes an expression will print output during its execution. You can indicate the printed output with the @print{} command.

The @print{} command is displayed as ‘-|’ in Info and as ‘ ⊣ ’ in the printed output.

In the following example, the printed text is indicated with ‘ ⊣ ’, and the the value of the expression follows on the last line.

    (progn (print 'foo) (print 'bar))
         ⊣ foo

         ⊣ bar
         ⇒ bar

In a Texinfo source file, this example is written as follows:

    @lisp
    (progn (print 'foo) (print 'bar))
         @print{} foo
         @print{} bar
         @result{} bar
    @end lisp

11.4 error : Indicating an Error Message

A piece of code may cause an error when you evaluate it. You can designate the error message with the @error{} command.

The @error{} command is displayed as ‘error-->’ in Info and as ‘ error ’ in the printed output.

Thus,

    @lisp
    (+ 23 'x)
    @error{} Wrong type argument: integer-or-marker-p, x
    @end lisp

produces

    (+ 23 'x)
     error Wrong type argument: integer-or-marker-p, x

This indicates that the following error message is printed when you evaluate the expression:

    Wrong type argument: integer-or-marker-p, x

Note that ‘ error ’ itself is not part of the error message.

11.5 _{≡}: Indicating Equivalence

Sometimes two expressions produce identical results. You can indicate the exact equivalence of two forms with the @equiv{} command.

The @equiv{} command is displayed as ‘==’ in Info and as ‘ ≡ ’ in the printed output.

Thus,

    @lisp
    (make-sparse-keymap) @equiv{} (list 'keymap)
    @end lisp

produces

    (make-sparse-keymap) ≡ (list 'keymap)

This indicates that evaluating (make-sparse-keymap) produces identical results to evaluating (list 'keymap).

11.6 Indicating Point in a Buffer

Sometimes you need to show an example of text in an Emacs buffer. In such examples, the convention is to include the entire contents of the buffer in question between two lines of dashes containing the buffer name.

You can use the ‘@point{}’ command to show the location of point in the text in the buffer. (The symbol for point, of course, is not part of the text in the buffer; it indicates the place between two characters where point is located.)

The @point{} command is displayed as ‘-!-’ in Info and as ‘⋆’ in the printed output.

The following example shows the contents of buffer ‘foo’ before and after evaluating a Lisp command to insert the word changed.

---------- Buffer: foo ----------
This is the ⋆contents of foo.
---------- Buffer: foo ----------
(insert "changed ")
     ⇒ nil
---------- Buffer: foo ----------
This is the changed ⋆contents of foo.
---------- Buffer: foo ----------

In a Texinfo source file, the example is written like this:

@example
---------- Buffer: foo ----------
This is the @point{}contents of foo.
---------- Buffer: foo ----------
(insert "changed ")
     @result{} nil
---------- Buffer: foo ----------
This is the changed @point{}contents of foo.
---------- Buffer: foo ----------
@end example

12 Making Lists and Tables

Texinfo has several ways of making lists and two-column tables. Lists can be bulleted or numbered, while two-column tables can highlight the items in the first column.

Texinfo automatically indents the text in lists or tables and numbers an enumerated list. This last feature is useful if you modify the list, since you do not have to renumber it yourself.

Numbered lists and tables begin with the appropriate @-command at the beginning of a line, and end with the corresponding @end command on a line by itself. The table and itemized-list commands also require that you write formatting information on the same line as the beginning @-command.

Begin an enumerated list, for example, with an @enumerate command and end the list with an @end enumerate command. Begin an itemized list with an @itemize command, followed on the same line by a formatting command such as @bullet, and end the list with an @end itemize command.

Precede each element of a list with an @item command.

Here is an itemized list of the different kinds of table and lists:

• Itemized lists with and without bullets.

• Numbered lists.

• Two-column tables with highlighting.

Here is an enumerated list with the same items:

1. Itemized lists with and without bullets.

2. Numbered lists.

3. Two-column tables with highlighting. And here is a two-column table with the same items and their @-

commands: @itemize Itemized lists with and without bullets.

@enumerate

Numbered lists.

@table

@ftable Two-column tables with highlighting.

12.1 Making an Itemized List

The @itemize command is used to produce sequences of indented paragraphs, with marks inside the left margin at the beginning of paragraphs for which a mark is desired. A paragraph can be shorter than single line.

Begin an itemized list by writing @itemize at the beginning of a line. Follow the command, on the same line, with a character or a Texinfo command that generates a mark. Usually, you will write @bullet after @itemize, but you can use @minus, or any character or any special symbol that results in a single character in the Info file. (When you write @bullet or @minus after an @itemize command, you may omit the ‘{}’; this saves typing.)

Write the text of the indented paragraphs themselves after the @itemize, up to another line that says @end itemize.

Before each paragraph for which a mark in the margin is desired, place a line that says just @item. Don’t put any other text on this line.

Usually, you should put a blank line before an @item. This puts a blank line in the Info file. (TEX inserts the proper interline whitespace in either case.) Except when the entries are very brief, these blank lines make the list look better.

Here is an example of the use of @itemize, followed by the output it produces. Note that @bullet produces a ‘*’ in Info and a round dot in TEX.

    @itemize @bullet
    @item
    Some text for foo.
    @item
    Some text
    for bar.
    @end itemize

produces • Some text for foo.

• Some text for bar.

Itemized lists may be embedded within other itemized lists. Here is a list marked with dashes embedded in a list marked with bullets:

@itemize @bullet
@item
First item.
@itemize @minus
@item
Inner item.
@item
Second inner item.
@end itemize
@item
Second outer item.
@end itemize

produces

• First item.

− Inner item.

− Second inner item.

• Second outer item.

12.2 Making a Numbered List

@enumerate is like @itemize except that the marks in the left margin contain successive integers starting with 1. (See Section 12.1 [@itemize], page 95.) Do not put any argument on the same line as @enumerate.

@enumerate
@item
Some text for foo.
@item
Some text
for bar.
@end enumerate

produces

1. Some text for foo.

2. Some text for bar.

Normally, you should put a blank line between the entries in the list. This generally makes it easier to read the Info file. For example,

    @enumerate
    @item
    This is the first item.
    @item
    This is the second item.
    @end enumerate

12.3 Making a Two-column Table

@table is similar to @itemize, but the command allows you to specify a name or heading line for each item. (See Section 12.1 [@itemize], page 95.) The @table command is used to produce two-column tables, and is especially useful for glossaries and explanatory exhibits.

Write the @table command at the beginning of a line and follow it on the same line by an argument that is a Texinfo command such as @code, @samp, @var, or @kbd. Also, you may use the @asis command. Although these commands are usually followed by arguments in braces, in this case you use the command name without an argument because @item will supply the argument. This command will be applied to the text that goes into the first column of each item and determines how it will be highlighted. For example, @samp will cause the text in the first column to be highlighted with an @samp command.

@asis is a command that does nothing; if you write this command after @table, TEX and the Info formatting commands output the first column entries without added highlighting, (‘as is’).

(The @table command may work with other commands besides those listed here. You can experiment. However, you can only use commands that normally take arguments in braces.)

Begin each table entry with an @item command at the beginning of a line. Write the first column text on the same line as the @item command. Write the second column text on the line following the @item line and on subsequent lines. (You don’t have to type anything for an empty second column entry.) You may write as many lines of supporting text as you wish, even several paragraphs. But only text on the same line as the @item will be placed in the first column.

The following table, for example, highlights the text in the first column with an @samp command:

    @table @samp
    @item foo
    This is the text for
    @samp{foo}.

    @item bar
    Text for @samp{bar}.
    @end table

produces ‘foo’ This is the text for ‘foo’. ‘bar’ Text for ‘bar’. If you want to list two or more named items with a single block of text,

use the @itemx command. (See Section 12.3.2 [@itemx], page 99.)

12.3.1 @ftable

The @ftable command is the same as the @table command except that it automatically enters each of the items in the first column of the table into the index of functions, which makes it easier to create indices. Only the items on the same line as the @item commands are indexed, and they are indexed in exactly the form that they appear on that line. See Chapter 15 [Creating Indices], page 119, for more information about indices.

Begin an @ftable environment by writing @ftable at the beginning of a line, followed on the same line by an argument that is a Texinfo command such as @code, exactly as you would for an @table command; and end the environment with an @end ftable command on a line by itself.

12.3.2 @itemx

Use the @itemx command inside a table when you have two or more first column entries for the same item, each of which should appear on a line of its own. Use @itemx for all but the first entry. The @itemx command works exactly like @item except that it does not generate extra vertical space above the first column text. For example,

    @table @code
    @item upcase
    @itemx downcase
    These two functions accept a character or a string as
    argument, and return the corresponding upper case (lower
    case) character or string.@refill
    @end table

produces

    upcase

downcase These two functions accept a character or a string as argument, and return the corresponding upper case (lower case) character or string.

(Note also that this example illustrates multi-line supporting text in a twocolumn table.)

13 Definition Commands: @deffn, etc.

The @deffn command and the other definition commands enable you to describe functions, variables, macros, commands, user options, special forms and other such artifacts in a uniform format.

For example, in the Info file, an @deffn command causes the name of the type of the item—‘Function’, ‘Variable’, or whatever—to appear at the beginning of the first line of the definition, followed by the item’s name and arguments. In the printed manual, the command causes TEX to print the item’s name and its arguments on the left margin and print the name of the type next to the right margin. In both output formats, the body of the definition is indented. Also, the name of the item is entered into the appropriate index: @deffn enters the name into the index of functions, @defvr enters it into the index of variables, and so on.

13.1 The Template for a Definition

To write a definition using the @deffn command, write the command at the beginning of a line and follow it on the same line by the name of the type of the item, the name of the item itself, and its arguments (if any). Then write the body of the definition on succeeding lines. (You may embed examples in the body.) Finally, end the definition with an @end deffn command written on a line of its own. (The other definition commands follow the same format.)

The template for a definition looks like this:

@deffn category name arguments…

body-of-definition

    @end deffn

For example,

    @deffn Command forward-word count
    This function moves point forward @var{count} words
    (or backward if @var{count} is negative). . . .
    @end deffn

produces

Command forward-word count

This function moves point forward count words (or backward if count is negative)….

Some of the definition commands are more general than others. The @deffn command, for example, is the general definition command for functions and the like—for entities that may take arguments. When you use this command, you specify the category to which the entity belongs. The @deffn command possesses three predefined, specialized variations, @defun, @defmac, and @defspec, that specify the category for you: “Function Definition”, “Macro”, and “Special Form” respectively. The @defvr command also is accompanied by several predefined, specialized variations for describing particular kinds of variables.

The template for a specialized definition, such as @defun, is similar to the template for a generalized definition, except that you don’t have to specify the category:

@defun name arguments…

body-of-definition

    @end defun

Thus,

    @defun buffer-end flag
    This function returns @code{(point-min)} if @var{flag}
    is less than 1, @code{(point-max)} otherwise.

…

    @end defun

produces

Function buffer-end flag

This function returns (point-min) if flag is less than 1, (pointmax) otherwise. …

See Section 13.5 [A Sample Function Definition], page 113, for a more detailed example of a function definition, including the use of an @example environment inside of the definition.

The other specialized commands are similar to the @deffn and @defun commands. Thus, macro (@defmac) and special form (@defspec) definitions have the same Info and printed format as function definitions, but the word ‘Function’ is replaced by ‘Macro’ or ‘Special Form’.

13.2 Optional and Repeated Parameters

Some entities take optional or repeated parameters, which may be specified by a distinctive special glyph that uses square brackets and ellipses. For example, a special form often breaks its argument list into separate arguments in more complicated ways than a straightforward function.

An argument enclosed within square brackets is optional. Thus, the phrase ‘[optional-arg]’ means that optional-arg is optional. An argument followed by an ellipsis is optional and may be repeated more than once. Thus, ‘repeated-args…’ stands for zero or more arguments. Parentheses are used when several arguments are grouped into additional levels of list structure in Lisp.

Here is the @defspec line of an example of an imaginary special form:

Special Form foobar (var [from to [inc]]) body…

In this example, the arguments from and to are optional, but must both be present or both absent. If they are present, inc may optionally be specified as well. These arguments are grouped with the argument var into a list, to distinguish them from body, which includes all remaining elements of the form.

In a Texinfo source file, this @defspec line is written like this (except it would not be split over two lines, as it is in this example).

    @defspec foobar (@var{var} [@var{from} @var{to}
         [@var{inc}]]) @var{body}@dots{}

Incidentally, the function is listed in the Command Index under ‘foobar’.

13.3 The Definition Commands

Texinfo provides more than a dozen definition commands, all of which are described in this section.

The definition commands automatically enter the name of the item in the appropriate index: for example, @deffn, @defun, and @defmac enter function names in the index of functions; @defvr and @defvar enter variable names in the index of variables.

Although the examples that follow mostly illustrate Lisp, the commands can be used for other programming languages: indeed, each command is itself described using an @deffn command!

13.3.1 Functions and Similar Entities

Here are the commands for functions and similar entities:

Def Command @deffn category name arguments… The @deffn command is the general definition command for functions, interactive commands, and similar entities that may take arguments. The @deffn command is written at the beginning of a line and is followed on the same line by the category of entity being described, the name of this particular entity, and its arguments, if any. For example,

    @deffn Command forward-char nchars

…

    @end deffn

shows the @deffn line you would use to begin a definition of a “command” named forward-char that takes nchars as its only argument. Terminate the definition with @end deffn on a line of its own. The template is:

@deffn category name arguments…

body-of-definition

    @end deffn

Capitalize the category name like a title. If the category name contains spaces, as in the phrase ‘Interactive Command’, write braces around it. For example,

    @deffn {Interactive Command} isearch-forward

…

    @end deffn

Otherwise, the second word will be mistaken for the name of the entity. The first line of this definition is written using an @deffn command and looks like this:

     @deffn {Def Command} @@deffn category name arguments. . .

@deffn creates an entry in the index of functions.

Def Command @defun name arguments… The @defun command is the definition command for functions.

@defun is equivalent to ‘@deffn Function . . .’.

For example,

    @defun symbol-value symbol

…

    @end defun

shows the @defun line you would use to begin a definition of the symbol-value function that takes symbol as its only argument. Terminate the definition with @end defun on a line of its own.

The template is:

@defun function-name arguments…

body-of-definition

    @end defun

@defun creates an entry in the index of functions.

Def Command @defmac name arguments… The @defmac command is the definition command for macros. @defmac is equivalent to ‘@deffn Macro …’.

For example,

    @defmac loop &rest body

…

    @end defmac

shows the @defmac line you would use to begin a definition of the loop macro that takes body as any number of arguments. Terminate the definition with @end defmac on a line of its own.

The template is:

@defmac macro-name arguments…

body-of-definition

    @end defmac

@defmac creates an entry in the index of functions.

Def Command @defspec name arguments… The @defspec command is the definition command for special forms. (In Lisp, a special form is an entity much like a function.)

@defspec is equivalent to ‘@deffn {Special Form} . . .’.

For example, the following line shows how you would begin a definition of the defconst special form for which symbol is a required argument, value is an optional argument, and if value is present, doc-string may be an argument. (See Section 13.2 [Optional and Repeated Parameters], page 102, for an explanation of the conventions for indicating optional and repeated parameters.)

    @defspec defconst symbol [value [doc-string]]

…

    @end defspec

Terminate the definition with @end defspec on a line of its own.

The template is:

@defspec special-form-name arguments… body-of-definition

    @end defspec

@defspec creates an entry in the index of functions.

13.3.2 Variables and Similar Entities

Here are the commands for variables and similar entities:

Def Command @defvr category name The @defvr command is a general definition command for something like a variable—an entity that records a value. You must choose a term to describe the category of entity being defined; for example, “Variable” could be used if the entity is a variable. The @defvr command is written at the beginning of a line and is followed on the same line by the category of the entity and the name of the entity. Capitalize the category name like a title. If the name of the category contains spaces, as in the name ‘User Option’, write braces around it. Otherwise, the second word will be mistaken for the name of the entity. For example,

    @defvr {User Option} fill-column
    This buffer-local variable specifies
    the maximum width of filled lines.

…

    @end defvr

Terminate the definition with @end defvr on a line of its own. The template is:

@defvr category name

body-of-definition

    @end defvr

@defvr creates an entry in the index of variables.

Def Command @defvar name

The @defvar command is the definition command for variables.

@defvar is equivalent to ‘@defvr Variable . . .’.

For example,

    @defvar kill-ring

…

    @end defvar

The template is:

    @defvar name

body-of-definition

    @end defvar

@defvar creates an entry in the index of variables.

Def Command @defopt name

The @defopt command is the definition command for user options. @defopt is equivalent to ‘@defvr {User Option} …’. For example,

    @defopt paragraph-start

…

    @end defopt

The template is:

@defopt option-name body-of-definition

    @end defopt

@defopt creates an entry in the index of variables.

13.3.3 Functions in Typed Languages

The @deftypefn command and its variations are for describing functions in C or any other language with types.

Def Command @deftypefn classification data-type name

arguments… The @deftypefn command is the general definition command for functions and similar entities that may take arguments and that are typed. The @deftypefn command is written at the beginning of a line and is followed on the same line by the classification of entity being described, the type, the name of this particular entity, and its arguments, if any. For example,

    @deftypefn {Library Function} int foobar (int @var-
    foo, float @varbar)

…

    @end deftypefn

produces the following in Info:

     * Library Function: int foobar (int FOO, float BAR)

…

In a printed manual, it produces:

Library Function int foobar (int foo, float bar)

…

This means that foobar is a Library Function that returns an int, and takes foo (an int) and bar (a float) as its arguments. Note that you need to specify which of the arguments to @deftypefn are metasyntactic variables (in this case, foo and bar). The template for @deftypefn is:

@deftypefn classification data-type name arguments …

body-of-description

     @end deftypefn

Note that if the classification or data type is more than one word then it must be enclosed in braces to make it a single argument. If you are describing a procedure in a language that has packages, such as Ada, you might consider using @deftypefn in a manner somewhat contrary to the convention described in the preceding paragraphs. For example:

    @deftypefn stacks private push (@vars:in out stack;@varn:in
    integer)

…

    @end deftypefn

In this instance, the procedure is classified as belonging to the package stacks rather than classified as a ‘procedure’ and its data type is described as private. (The name of the procedure is push, and its arguments are s and n.) @deftypefn creates an entry in the index of functions.

Def Command @deftypefun data-type name arguments… The @deftypefun command is the specialized definition command for functions in typed languages. The command is equiv-

alent to ‘@deftypefn Function . . .’.
    @deftypefun int foobar (int foo, float bar)

…

    @end deftypefun

produces the following in Info:

    * Function: int foobar (int FOO, float BAR)

…

and the following in a printed manual:

Function int foobar-int (int foo, float bar)

…

The template is:

@deftypefun type name arguments… body-of-description

    @end deftypefun

@deftypefun creates an entry in the index of functions.

13.3.4 Variables in Typed Languages

Variables in typed languages are handled in a manner similar to functions in typed languages. See Section 13.3.3 [Typed Functions], page 107. The general definition command @deftypevr corresponds to @deftypefn and the specialized definition command @deftypevar corresponds to @deftypefun.

Def Command @deftypevr classification data-type name The @deftypevr command is the general definition command for something like a variable in a typed language—an entity that records a value. You must choose a term to describe the classification of the entity being defined; for example, “Variable” could be used if the entity is a variable. The @deftypevr command is written at the beginning of a line and is followed on the same line by the classification of the entity being described, the data type, and the name of this particular entity. For example:

    @deftypevr {Global Flag} int enable

…

    @end deftypevr

produces the following in Info:

    Global Flag: int enable

…

and the following in a printed manual:

Global Flag int enable

…

The template is:

@deftypevr classification data-type name body-of-description

    @end deftypevr

@deftypevr creates an entry in the index of variables.

Def Command @deftypefvar data-type name The @deftypevar command is the specialized definition command for variables in typed languages. @deftypevar is equiva-

lent to ‘@deftypevr Variable . . .’.

For example,

    @deftypevar int foobar

…

    @end deftypevar

produces the following in Info:

    * Variable: int foobar

…

and the following in a printed manual:

Variable int foobar

…

The template is:

@deftypevar data-type name

body-of-description

    @end deftypevar

@deftypevar creates an entry in the index of variables.

13.3.5 Object-Oriented Programming

Here are the commands for formatting descriptions about abstract objects, such as are used in object-oriented programming:

Def Command @defcv category class name The @defcv command is the general definition command for variables associated with classes in object-oriented programming. The @defcv command is followed by three arguments: the category of thing being defined, the class to which it belongs, and its name. Thus,

    @defcv {Class Option} Window  border-pattern

…

    @end defcv

illustrates how you would write the first line of a definition of the border-pattern class option of the class ‘Window’. The template is @defcv category class name

…

    @end defcv

@defcv creates an entry in the index of variables.

Def Command @defivar class name The @defivar command is the definition command for instance variables in object-oriented programming. @defivar is equiva-

lent to ‘@defcv {Instance Variable} . . .’

The template is:

@defivar class instance-variable-name body-of-definition

    @end defivar

@defivar creates an entry in the index of variables.

Def Command @defop category class name arguments… The @defop command is the general definition command for entities that may resemble methods in object-oriented programming. These entities take arguments, as functions do, but are associated with particular classes of objects. For example, some systems have constructs called wrappers that are associated with classes as methods are, but that act more like macros than like functions. You could use @defop Wrapper to describe one of these. Sometimes it is useful to distinguish methods and operations. You can think of an operation as the specification for a method. Thus, a window system might specify that all window classes have a method named expose; we would say that this window system defines an expose operation on windows in general. Typically, the operation has a name and also specifies the pattern of arguments; all methods that implement the operation must accept the same arguments, since applications that use the operation do so without knowing which method will implement it. Often it makes more sense to document operations than methods. For example, window application developers need to know about the expose operation, but need not be concerned with whether a given class of windows has its own method to implement this operation. To describe this operation, you would write:

    @defop Operation windows expose

The @defop command is written at the beginning of a line and is followed on the same line by the overall name of the category of operation, the name of the class of the operation, the name of the operation, and its arguments, if any. The template is:

@defop category class name arguments… body-of-definition

    @end defop

If the class name is truly a name defined in the programming system for a class, then you should write an @code around it. Otherwise, it is printed in the usual text font. @defop creates an entry, such as ‘expose on windows’, in the index of functions.

Def Command @defmethod class name arguments… The @defmethod command is the definition command for methods in object-oriented programming. A method is a kind of function that implements an operation for a particular class of objects and its subclasses. In the Lisp Machine, methods actually were functions, but they were usually defined with defmethod. @defmethod is equivalent to ‘@defop Method …’. The command is written at the beginning of a line and is followed by the name of the class of the method, the name of the method, and its arguments, if any. For example,

    @defmethod bar-class bar-method argument

…

    @end defmethod

illustrates the definition for a method called bar-method of the class bar-class. The method takes an argument. The first argument is printed in roman type unless you put it in @code, which you must do if it is the name of a specific class. The template is:

@defmethod class method-name arguments… body-of-definition

    @end defmethod
@defmethod creates an entry, such as ‘bar-method on bar-

class’, in the index of functions.

13.3.6 Data Types

Here is the command for data types:

Def Command @deftp category name attributes… The @deftp command is the generic definition command for data types. The command is written at the beginning of a line and is followed on the same line by the category, by the name of the type (which is a word like ‘int’ or ‘float’), and then by names of attributes of objects of that type. Thus, you could use this command for describing ‘int’ or ‘float’, in which case you could use ‘data type’ as the category. (A data type is a classification of certain objects for purposes of deciding which operations can be performed on them.) In Lisp, for example, a dotted pair names a particular data type, and an object of that type has two attributes called the car and the cdr. Here is how you would write the first line of a definition of ‘dotted-pair’.

    @deftp {Data type} dotted-pair car cdr

…

    @end deftp

The template is:

@deftp category name-of-type attributes… body-of-definition

    @end deftp

@deftp creates an entry in the index of data types.

13.4 Conventions for Writing Definitions

When you write a definition using @deffn, @defun, or one of the other definition commands, please take care that any argument whose name contains the name of a type (e.g., integer, integer1 or buffer) actually is of that type. Please ensure that an argument name in the plural (such as buffers) is a list or collection of objects of that type. By convention, arguments named object may be of any type. Readers find it helpful if you discuss the types of arguments explicitly when the names do not state the types.

13.5 A Sample Function Definition

A function definition uses the @defun and @end defun commands. The name of the function follows immediately after the @defun command and it is followed, on the same line, by the parameter list.

Here is a definition of an imaginary function foo:

Function foo integer1 &optional integer2 &rest integers The function foo subtracts integer1 from integer2, then adds all the rest of the arguments to the result. If integer2 is not supplied, then the number 19 is used by default.

     (foo 1 5 3 9)
          ⇒ 16
     (foo 5)
          ⇒ 14

More generally,

     (foo w x y z)

≡

     (+ (- x w) y z)

In a Texinfo source file, this example looks like this:

@defun foo integer1 &optional integer2 &rest integers
The function @code{foo} subtracts @var{integer1} from
@var{integer2}, then adds all the rest of the arguments
to the result.  If @var{integer2} is not supplied, then
the number 19 is used by default.
@example
(foo 1 5 3 9)
     @result{} 16
(foo 5)
     @result{} 14
@end example
More generally,
@example
(foo @var{w} @var{x} @var{y} @var{z})
@equiv{}
(+ (- @var{x} @var{w}) @var{y} @var{z})
@end example
@end defun

In this manual, this function is listed in the Command Index under ‘foo’.

13.6 A Sample Variable Definition

Ordinary variables and user options are described using a format like that for functions except that variables do not take arguments. (See Section 13.5 [A Sample Function Definition], page 113.)

Here is a definition of the imaginary electric-future-map variable.

Variable electric-future-map

The value of this variable is a full keymap used by electric command future mode. The functions in this map will allow you to edit commands you have not yet thought about executing.

In a Texinfo source file, this variable definition looks like this:

@defvar electric-future-map
The value of this variable is a full keymap used by
electric command future mode.  The functions in this map
will allow you to edit commands you have not yet thought
about executing.
@end defvar

116 Texinfo 2.0

14 Footnotes

A footnote is for a reference that documents or elucidates the primary text. ^{1}

In Texinfo, footnotes are created with the @footnote command. This command is followed immediately by a left brace, then by the text of the footnote, and then by a terminating right brace. The template is:

@footnote{text}.

Footnotes may be of any length, but are usually short.

For example, this clause is followed by a sample footnote ^{2}; in the Texinfo source, it looks like this:

    . . .a sample footnote @footnote{Here is the sample
    footnote.}; in the Texinfo source. . .

In a printed manual or book, the reference mark for a footnote is a small, superscripted number; the text of the footnote is written at the bottom of the page, below a horizontal line.

In Info, the reference mark for a footnote is a pair of parentheses, ‘()’, with nothing between them.

Info has two footnote styles, which determine where the test of the footnote is:

‘End Node’

In the End Node style, all the footnotes for a single node are placed at the end of that node. The footnotes are separated from the rest of the node by a line of dashes with the word ‘Footnotes’ within it. Each footnote begins with a ‘()’ reference mark.

Here is an example of a single footnote in the “End Node” style:

     --------- Footnotes ---------
    ()  Here is a sample footnote.
‘Make Node’

1 A footnote should complement or expand upon the primary text, but a reader should not need to read a footnote to understand the primary text. For a thorough discussion of footnotes, see The Chicago Manual of Style, which is published by the University of Chicago Press.

2 Here is the sample footnote.

In the Make Node style, all the footnotes for a single node are placed in an automatically constructed node of their own. In this style, a “footnote reference” follows each ‘()’ reference mark in the body of the node. Each footnote reference has the same form as a cross reference and you use it to reach the footnote. The node containing the footnotes is given a name that is constructed by appending ‘-Footnotes’ to the name of the node that contains the footnote command. (Consequently, the foot-

          notes’ node for the ‘Footnotes’ node is Footnotes-Footnotes!)

The footnotes’ node has an ‘Up’ node pointer that leads back to its parent node.

Here is an example of a footnote for this ‘Footnotes’ chapter in the “Make Node” style:

    File: Texinfo, Node: Footnotes-Footnotes, Up: Footnotes
    ()  Here is a sample footnote.

A Texinfo file may be formatted into an Info file with either footnote style.

If you use makeinfo to create the Info file, the +footnote-style option determines which style is used, ‘EN’ for the “End Node” style or ‘MN’ for the “make node” style. Thus, to format the Texinfo manual in the “Make Node” style, you would use the following shell command:

    makeinfo +footnote-style MN texinfo.texinfo

To format the Texinfo manual in the “End Node” style, you would type:

    makeinfo +footnote-style EN texinfo.texinfo
   If you use texinfo-format-buffer or texinfo-format-region to create

the Info file, the value of the texinfo-footnote-style variable controls the footnote style. It can be either ‘MN’ for the “Make Node” style or ‘EN’ for the “End Node” style. (You can change the value of this variable with the M-x edit-options command (see section ‘Editing Variable Values’ in The GNU Emacs Manual), or with the M-x set-variable command (see section ‘Examining and Setting Variables’ in The GNU Emacs Manual).

The texinfo-footnote-style variable also controls the style if you use

the M-x makeinfo-region or M-x makeinfo-buffer command in Emacs.

This chapter contains two footnotes.

15 Creating Indices

Using Texinfo, you can generate indices without having to sort and collate entries manually. In an index, the entries are listed in alphabetical order, together with information on how to find the discussion of each entry. In a printed manual, this information consists of page numbers. In an Info file, this information is a menu item leading to the first node referenced.

Texinfo provides several predefined kinds of index: an index for functions, an index for variables, an index for concepts, and so on. You can combine indices or use them for other than their canonical purpose. If you wish, you can define your own indices.

15.1 Making Index Entries

When you are making index entries, it is good practice to think of the different ways people may look for something. Different people do not think of the same words when they look something up. A helpful index will have items indexed under all the different words that people may use. For example, someone might think it obvious that the two-letter names for indices should be listed under “Indices, two-letter names”, since the word “Index” is the general concept. But another reader may remember the specific concept of two-letter names and search for the entry listed as “Two letter names for indices”. A good index will have both entries and will help both kinds of user.

Like typesetting, the construction of an index is a highly skilled, professional art, the subtleties of which are not appreciated until you have to do it yourself.

See Section 4.1 [Printing an Index and Generating Index Menus], page 41, for information about printing an index at the end of a book or creating an index menu in an Info file.

15.2 Predefined Indices

Texinfo provides six predefined indices:

• A concept index listing concepts that are discussed.

• A function index listing functions (such as, entry points of libraries).

• A variables index listing variables (such as, global variables of libraries).

• A keystroke index listing keyboard commands.

• A program index listing names of programs.

• A data type index listing data types (such as, structures defined in

header files). Not every manual needs all of these. This manual has two indices: a concept index and an @-command index (that is actually the function index but is called a command index in the chapter heading). Two or more indices can be combined into one using the @synindex or @syncodeindex commands. See Section 15.4 [Combining Indices], page 121.

15.3 Defining the Entries of an Index

The data to make an index comes from many individual indexing commands scattered throughout the Texinfo source file. Each command says to add one entry to a particular index; after processing, it will give the current page number or node name as the reference.

An index entry consists of an indexing command at the beginning of a line followed, on the rest of the line, by the entry.

For example, this section begins with the following five entries for the concept index:

    @cindex Defining indexing entries
    @cindex Index entries
    @cindex Entries for an index
    @cindex Specifying index entries
    @cindex Creating index entries

Each predefined index has its own indexing command—@cindex for the concept index, @findex for the function index, and so on.

By default, entries for a concept index are printed in a small roman font and entries for the other indices are printed in a small @code font. You may change the way part of an entry is printed with the usual Texinfo commands, such as @file for file names and @emph for emphasis (see Chapter 8 [Marking Text], page 69).

The six indexing commands are:

@cindex concept

Make an entry in the concept index for concept.

@findex function

Make an entry in the function index for function.

@vindex variable Make an entry in the variable index for variable.

@kindex key

Make an entry in the key index for key.

@pindex program

Make an entry in the program index for program.

@tindex data type

Make an entry in the data type index for data type. If the same name is indexed on several pages, all the pages are listed in the

printed manual’s index. However, only the first node referenced will appear in the index of an Info file. This means that it is best to write indices in which each entry will refer to only one place in the Texinfo file. Fortunately, this constraint is a feature rather than a loss since it means that the index will be easy to use. Otherwise, it can be easy to create an index which has many pages listed for an entry and you don’t know which one you need. If you have two entries for one topic, change the topics slightly, or qualify them to indicate the difference.

For example, the following pair of index entries refer to different pages:

    @cindex Installing an Info file
    @cindex Installing Info in another directory

as do the following:

    @cindex Info file installation
    @cindex Info installed in another directory

You are not actually required to use the predefined indices for their canonical purposes. For example, suppose you wish to index some C preprocessor macros. You could put them in the function index along with actual functions, just by writing @findex commands for them; then, when you print the “function index” as an unnumbered chapter, you could give it the title ‘Function and Macro Index’ and all will be consistent for the reader. Or you could put the macros in with the data types by writing @tindex commands for them, and give that index a suitable title so the reader will understand. (See Section 4.1 [Printing Indices & Menus], page 41.)

15.4 Combining Indices

Sometimes you will want to combine two disparate indices such as functions and concepts, perhaps because you have few enough of one of them that a separate index for them would look silly.

You could put functions into the concept index by writing @cindex commands for them instead of @findex commands, and produce a consistent manual by printing the concept index with the title ‘Function and Concept Index’ and not printing the ‘Function Index’ at all; but this is not a robust procedure. It works only if your document is never included in part of or together with another document that is designed to have a separate function index; if your document were to be included with such a document, the functions from your document and those from the other would not end up together. Also, to make your function names appear in the right font in the concept index, you would have to enclose every one of them in an @code environment.

What you should do instead when you want functions and concepts in one index is to index the functions with @findex as they should be, but use the @syncodeindex command to redirect these @findex commands to the concept index.

The @syncodeindex command takes two arguments: the name of an index to redirect, and the name of an index to redirect it to:

    @syncodeindex from to

For this purpose, the indices are given two-letter names: ‘cp’ the concept index ‘vr’ the variable index ‘fn’ the function index ‘ky’ the key index ‘pg’ the program index ‘tp’ the data type index

Thus, to merge a function index with a concept index, place the following at the front of a Texinfo file, within or right after the end-of-header line:

    @syncodeindex fn cp

This will cause all entries designated for the function index to go to the concept index instead.

The @syncodeindex command puts all the entries from the redirected index into the @code font. This way all the function names are in the @code font as you would expect. Do the same when you merge a variable or other index with a concept index.

The @synindex command is nearly the same as the @syncodeindex command, except that it does not put the redirected index into the @code font, but puts it in the roman font.

See Section 4.1 [Printing an Index and Generating Index Menus], page 41, for information about printing an index at the end of a book or creating an index menu in an Info file.

15.5 Defining New Indices

In addition to the predefined indices, you may use the @defindex and @defcodeindex commands to define new indices. These commands create new indexing @-commands with which you mark index entries. The @defindex command is used like this:

    @defindex name

The name of an index should be a two letter word, such as ‘au’.

    @defindex au

In the preceding example, the @defindex command defines a new index, called the ‘au’ index. At the same time, the @defindex command creates a new indexing command, @auindex, that you can use to mark index entries. Use the new indexing command just as you would use a predefined indexing command.

For example, here is a section heading followed by a concept index entry and two ‘au’ index entries.

    @section Cognitive Semantics
    @cindex kinesthetic image schemas
    @auindex Johnson, Mark
    @auindex Lakoff, George

(Evidently, ‘au’ serves here as a abbreviation for “author”.) Texinfo constructs the new indexing command by concatenating the name of the index with ‘index’; thus, defining an ‘au’ index leads to the automatic creation of an @auindex command.

Use the @printindex command to print the index, as you do with the predefined indices. For example,

    @node Author Index, Subject Index, , Top
    @unnumbered Author Index
    @printindex au

The @defcodeindex is like the @defindex command, except that in the printed output, it prints entries in an @code font, like the @findex command, rather than in a roman font, like the @cindex command.

You should define new indices within or right after the end-of-header line of a Texinfo file, before any @synindex or @syncodeindex commands (see Section 3.2 [Header], page 29).

124 Texinfo 2.0

16 Making and Preventing Breaks

Usually, a Texinfo file is processed both by TEX and by one of the Info formatting commands. Sometimes line, paragraph, or page breaks occur in the ‘wrong’ place in one or other form of output. You must ensure that text looks right both in the printed manual and in the Info file.

For example, in a printed manual, page breaks may occur awkwardly in the middle of an example; to prevent this, you can hold text together using a grouping command that keeps the text from being split across two pages. Conversely, you may want to force a page break where none would occur normally.

Fortunately, problems like these do not often arise.

The break commands create line and paragraph breaks: @* Force a line break. @br Force a paragraph break.

The line-break-prevention command holds text together all on one line and the skip command creates blank lines. @w{text} Prevent text from being split across two lines. @sp n Skip n blank lines.

The pagination commands apply only to printed output since Info files do not have pages. @page Start a new page in the printed manual. @group Hold text together that must appear on one printed page.

@need mils

Start a new printed page if not enough space on this one.

16.1 @*: Generate Line Breaks

The @* command forces a line break in both the printed manual and in Info.

For example,

    This line @* is broken @*in two places.

produces

    This line
     is broken
    in two places.

(Note that the space after the first @* command is faithfully carried down to the next line.)

The @* command is often used in a file’s copyright page:

    This is version 2.0 of the Texinfo documentation,@*
    and is for . . .

In this case, the @* command keeps TEX from stretching the line across the whole page in an ugly manner.

Do not write braces after an @* command; they are not needed.

Do not write an @refill command at the end of a paragraph containing an @* command; it will cause the paragraph to be refilled after the line break occurs, negating the effect of the line break.

16.2 @w{text}: Prevent Line Breaks

@w{text} outputs text and prohibits line breaks within text.

You can use the @w command to prevent TEX from automatically hyphenating a long name or phrase that accidentally falls near the end of a line.

    You can copy GNU software from @w{@file{prep.ai.mit.edu}}.

produces You can copy GNU software from ‘prep.ai.mit.edu’. In the Texinfo file, you must write the @w command and its argument (all

the affected text) all on one line.

Do not write an @refill command at the end of a paragraph containing an @w command; it will cause the paragraph to be refilled and may thereby negate the effect of the @w command.

16.3 @sp n: Insert Blank Lines

A line beginning with and containing only @sp n generates n blank lines of space in both the printed manual and the Info file. For example,

    @sp 2

generates two blank lines.

The @sp command is most often used in the title page.

16.4 @br: Generate Paragraph Breaks

The @br command forces a paragraph break. It inserts a blank line. You can use the command within or at the end of a line. If used within a line, the @br{} command must be followed by left and right braces (as shown here) to mark the end of the command.

For example,

    This line @br{}contains and is ended by paragraph breaks@br
    and is followed by another line.

produces

    This line
    contains and is ended by paragraph breaks
    and is followed by another line.

The @br command is seldom used.

16.5 @page: Start a New Page

A line containing only @page starts a new page in a printed manual. The command has no effect on Info files since they are not paginated. An @page command is often used in the title section of a Texinfo file to start the copyright page.

16.6 @group: Prevent Page Breaks

The @group command (on a line by itself) begins an unsplittable vertical group, which will appear entirely on one page in the printed output. The group is terminated by a line containing only @end group. These two lines produce no output of their own, and in the Info file output they have no effect at all.

The @group command is most often used to hold an example together on one page. In this Texinfo manual, about 100 examples are enclosed between

@group and @end group.

If you forget to end a group, you may get strange and unfathomable error messages when you run TEX. This is because TEX keeps trying to put the rest of the Texinfo file onto the one page and does not start to generate error messages until it has processed considerable text. It’s a good rule of thumb to look for a missing @end group if you get incomprehensible error messages in TEX.

16.7 @need mils: Prevent Page Breaks

A line containing only @need n starts a new page in a printed manual if fewer than n mils (thousandths of an inch) remain on the current page. Do not use braces around the argument n. The @need command has no effect on Info files since they are not paginated.

This paragraph is preceded by an @need command that tells TEX to start a new page if fewer than 300 mils (nearly one-third inch) remain on the page. It looks like this:

    @need 300
    This paragraph is preceded by . . .

The @need command is useful for preventing orphans (single lines at the bottoms of printed pages).

17 Conditionally Visible Text

Sometimes it is not good to use the same text for both a printed manual and a corresponding Info file. In this case, you can use the conditional commands to specify which text is for the printed manual and which is for the Info file.

@ifinfo begins text that should be ignored by TEX when it typesets the printed manual. The text appears only in the Info file. The @ifinfo command should appear on a line by itself. End the Info-only text with a line containing @end ifinfo by itself. At the beginning of a Texinfo file, the Info permissions are contained within a region marked by @ifinfo and @end ifinfo. (See Section 3.3 [Info Summary and Permissions], page 32.)

The @iftex and @end iftex commands are used similarly but to delimit text that will appear in the printed manual but not in the Info file.

For example,

    @iftex
    This text will appear only in the printed manual.
    @end iftex
    @ifinfo
    However, this text will appear only in Info.
    @end ifinfo

The preceding example produces the following. Note how you only see one of the two lines, depending on whether you are reading the Info version or the printed version of this manual.

This text will appear only in the printed manual.

The @titlepage command is a special variant of @iftex that is used for making the title and copyright pages of the printed manual. (See Section 3.4.1 [@titlepage], page 33.)

17.1 Using Ordinary TEX Commands

Inside a region delineated by @iftex and @end iftex, you can embed ordinary TEX commands. Info will ignore these commands since they are only in that part of the file that is seen by TEX. The TEX commands are the same as any TEX commands except that an ‘@’ replaces the ‘\’ used by TEX. For example, in the @titlepage section of a Texinfo file, the TEX command @vskip is used to format the copyright page. (The @titlepage command causes Info to ignore the region automatically, as it does with the @iftex command.)

You can enter TEX completely, and use ‘\’ in the TEX commands by delineating a region with the @tex and @end tex commands. (The @tex command causes Info to ignore the region automatically, as it does with the @iftex command.)

For example, here is some mathematics:

    @tex
    $\bigl(x\in A(n)\bigm|x\in B(n)\bigr)$
    @end tex

The output of this example will appear only in the printed manual. If you are reading this in Info, you will not see anything more. In the printed manual, the above mathematics looks like this:

(x ∈ A(n) | x ∈ B(n))

18 Printing Hardcopy

There are three shell commands for printing hardcopy of a Texinfo file. One is for formatting the file, the second is for sorting the index, and the third is for printing the formatted document. When you use the shell commands, you can either work directly in the operating system shell or work within a shell inside of GNU Emacs.

You can also format a file, sort the indices, and print the result using Emacs commands provided by Texinfo mode.

The typesetting program TEX is used for formatting a Texinfo file.^{1} TEX is a very powerful typesetting program and, if used right, does an exceptionally good job.

   The makeinfo, texinfo-format-region, and texinfo-format-buffer

commands read the very same @-commands in the Texinfo file as does TEX but process them differently from TEX to make an Info file. (See Chapter 19 [Creating an Info File], page 139.)

By convention, Texinfo files end with either of the three extensions ‘.texinfo’, ‘.texi’, or ‘.tex’. The longer extension is preferred since it describes more clearly to a human reader the nature of the file.

18.1 How to Print Using Shell Commands

Format the Texinfo file with the shell command tex followed by the name of the Texinfo file. This produces a formatted dvi file as well as several auxiliary files containing indices, cross references, etc. The dvi file (for DeVice Independent file) can be printed on a wide variety of printers.

The tex formatting command itself does not sort the indices; it writes an output file of unsorted index data. This is a misfeature of TEX. Hence, to generate a printed index, you first need a sorted index to work from. The texindex command sorts indices.

TEX outputs unsorted index files under names that obey a standard convention. These names are the name of your main input file to TEX, with everything after the first period thrown away, and the two letter names of indices added at the end. For example, the raw index output files for the input

file ‘foo.texinfo’ would be ‘foo.cp’, ‘foo.vr’, ‘foo.fn’, ‘foo.tp’, ‘foo.pg’

1 If you do not have TEX, you can use the texi2roff program for formatting. and ‘foo.ky’. Those are exactly the arguments to give to texindex. Or else, you can use ‘??’ as “wild-cards” and give the command in this form:

    texindex foo.??

This command will run texindex on all the unsorted index files, including any that you have defined yourself using @defindex or @defcodeindex.

For each file specified, texindex generates a sorted index file whose name is made by appending ‘s’ to the input file name. The @printindex command knows to look for a file of that name. (You may execute texindex foo.?? even if there are similarly named files with two letter extensions that are not index files, such as ‘foo.el’. The texindex command reports but otherwise ignores such files.) texindex does not alter the raw index output file.

After you have sorted the indices, you need to rerun the tex command on the Texinfo file. This regenerates a formatted dvi file with up-to-date index entries.

To summarize, this is a three step process:

1. Run the tex command on the Texinfo file. This generates the formatted dvi file as well as the raw index files with two letter extensions.

2. Run the shell command texindex on the raw index files to sort them. This creates the corresponding sorted index files.

3. Rerun the TEX command on the Texinfo file. This regenerates a formatted dvi file with the index entries in the correct order. This second run also makes all the cross references correct as well. (The tables of contents are always correct.) You need not run texindex after each TEX run. If you don’t, the next

TEX run will use whatever sorted index files happen to exist from the previous use of texindex. This is usually ok while you are debugging.

Finally, you can print the document with the dvi print command. The name of the command depends on the system; lpr -d is common. The dvi print command may require a file name without any extension or with a ‘.dvi’ extension.

The following commands, for example, sort the indices, format, and print the Bison Manual (where ‘%’ is the shell prompt):

    % tex bison.texinfo
    % texindex bison.??
    % tex bison.texinfo
    % lpr -d bison.dvi

(Remember that the shell commands may be different at your site; but these are commonly used versions.)

18.2 Printing from an Emacs Shell

You can give formatting and printing commands from a shell within GNU Emacs. To create a shell within Emacs, type M-x shell. In this shell, you can format and print the document. See Section 18.1 [How to Print Using Shell Commands], page 131, for details.

You can switch to and from the shell buffer while tex is running and do other editing. If you are formatting a long document on a slow machine, this can be very convenient.

Running a shell that you create yourself in Emacs is an alternative to using the predefined key commands in Texinfo mode (see Section 18.3 [Texinfo Mode Printing], page 133) and is sometimes more convenient. In both cases, the formatting and printing commands are run in a shell that you can switch to and from. However, in Texinfo mode, the auxiliary files and the dvi file are placed in the ‘/tmp’ directory; on the other hand, when you create your own shell, these files are placed in whatever directory you specify. (Of course, you can redefine the value of the tex-directory to accomplish the same end when using Texinfo mode.)

18.3 Formatting and Printing in Texinfo Mode

Texinfo mode provides several predefined key commands for TEX formatting and printing. These include commands for sorting indices, looking at the printer queue, killing the formatting job, and recentering display of the buffer in which the operations occur.

Some of the commands are taken directly from TEX mode, but are bound to keys in Texinfo mode.

C-c C-t C-r
texinfo-tex-region

Run TEX on the current region.

C-c C-t C-b
texinfo-tex-buffer

Run TEX on the current buffer.

C-c C-t C-i
texinfo-texindex

Sort the indices of a Texinfo file formatted with texinfo-tex-

          region or texinfo-tex-buffer.
C-c C-t C-p
texinfo-tex-print

Print a dvi file that was made with texinfo-tex-region or

          texinfo-tex-buffer.
C-c C-t C-q
tex-show-print-queue

Show the print queue.

C-c C-t C-k
tex-kill-job
          Kill the shell created by texinfo-tex-region or texinfo-tex-
          buffer.
C-c C-t C-l
tex-recenter-output-buffer

Recenter the display showing the shell buffer in which the TEX printing and formatting commands are run.

Thus, the usual sequence of commands for formatting a buffer is as follows (with comments to the right):

C-c C-t C-b Run TEX on the buffer.

C-c C-t C-i Sort the indices.

C-c C-t C-b Rerun TEX to regenerate indices.

C-c C-t C-p Print the dvi file.

C-c C-t C-q Display the printer queue.

The Texinfo mode TEX formatting commands start a subshell in Emacs

called the ‘*tex-shell*’. The texinfo-tex-command, texinfo-texindex-

command, and tex-dvi-print-command commands are all run in this shell. Temporary files, including the dvi and index files, go into the ‘/tmp’ directory.

You can watch the commands operate in the ‘*tex-shell*’ buffer, and you can switch to and from and use the ‘*tex-shell*’ buffer as you would any other shell buffer.

The formatting and print commands depend on the values of several variables. The default values are:

Variable Default value

    texinfo-tex-command             "tex"
    texinfo-texindex-command        "texindex"
    tex-dvi-print-command           "lpr -d"
    tex-show-queue-command          "lpq"
    tex-directory                   "/tmp"
   The default values of texinfo-tex-command and texinfo-texindex-

command are set by the code for Texinfo mode; the default values of tex-

dvi-print-command, tex-show-queue-command and tex-directory are set

by the code for TEX mode.

You can change the values of these variables with the M-x edit-options command (see section ‘Editing Variable Values’ in The GNU Emacs Manual), with the M-x set-variable command (see section ‘Examining and Setting Variables’ in The GNU Emacs Manual), or with your ‘.emacs’ initialization file (see section ‘Init File’ in The GNU Emacs Manual).

18.4 Using the Local Variables List

Yet another way to apply the TEX formatting command to a Texinfo file is to put that command in a local variables list at the end of the Texinfo file. You can then specify the TEX formatting command as a compilecommand and have Emacs run the TEX formatting command by typing M-x compile. This creates a special shell called the ‘*compilation buffer*’ in which Emacs runs the compile command. For example, at the end of the ‘gdb.texinfo’ file, after the @bye, you would put the following:

    @c Local Variables:
    @c compile-command: "tex gdb.texinfo"
    @c End:

This technique is most often used by programmers who also compile programs this way. (See section ‘Compilation’ in The GNU Emacs Manual.)

18.5 TEX Formatting Requirements Summary

Every Texinfo file that is to be input to TEX must begin with an \input command like this:

    \input texinfo

This instructs TEX to load the macros it needs to process a Texinfo file. You must also include a line that specifies the title of the printed manual: @settitle Name of Manual Finally, every Texinfo file must end with a line saying

    @bye

which terminates TEX processing and forces out unfinished pages.

Strictly speaking these three lines are all a Texinfo file needs for TEX, besides the body.

(The @setfilename line is the only line that a Texinfo file needs for Info formatting.)

Usually, of course, the file’s first line contains an ‘@c -*-texinfo-*-’ comment that causes Emacs to switch to Texinfo mode when you edit the file. In addition, the beginning usually includes an @setfilename for Info formatting, an @setchapternewpage command, a title page, a copyright page, and permissions.

(See Section 3.2.5 [@setchapternewpage], page 31, for more information about the @setchapternewpage command. See Appendix C[Page Headings], page 169, for information about customizing headings and footings. See Section 3.4 [Titlepage & Copyright Page], page 32, for information about creating a title page.)

Besides an @bye, the end of a file usually includes indices and a table of contents.

(See Section 4.1 [Index Menus and Printing an Index], page 41, for information about printing indices. See Section 4.2 [Contents], page 42, for information about creating a table of contents.)

18.6 TEX Input Initialization

TEX needs to know where to find the ‘texinfo.tex’ file that that you have told it to input with the ‘\input texinfo’ command at the beginning of the first line. The ‘texinfo.tex’ file tells TEX how to handle @-commands.

Usually, the ‘texinfo.tex’ file is put in the default directory that contains TEX macros, namely ‘/usr/lib/tex/macros’ directory, when GNU Emacs is installed. In this case, TEX will find the file and you don’t have to do anything special. Alternatively, you can put ‘texinfo.tex’ in the directory in which the Texinfo source file is located, and TEX will find it there.

However, if you may want to specify the location of the \input file yourself. One way to do this is to write the complete path for the file after the \input command. Another way is to set the ‘TEXINPUTS’ environment variable in your ‘.cshrc’ or ‘.profile’ file. The ‘TEXINPUTS’ environment variable will tell TEX where to find the ‘texinfo.tex’ file and any other file that you might want TEX to use.

Whether you use a ‘.cshrc’ or ‘.profile’ file depends on whether you use ‘csh’ or ‘sh’ for your shell command interpreter. When you use ‘csh’, it looks to the ‘.cshrc’ file for initialization information, and when you use ‘sh’, it looks to the ‘.profile’ file.

In a ‘.cshrc’ file, you could use the following csh command sequence:

    setenv TEXINPUTS .:/usr/me/mylib:/usr/lib/tex/macros

In a ‘.profile’ file, you could use the following sh command sequence:

    TEXINPUTS=.:/usr/me/mylib:/usr/lib/tex/macros
    export TEXINPUTS

This would cause TEX to look for ‘\input’ file first in the current directory, indicated by the ‘.’, then in a hypothetical user’s ‘me/mylib’ directory, and finally in the system library.

18.7 Overfull “Hboxes”

TEX is sometimes unable to typeset a line without extending it into the right margin. This can occur when TEX comes upon what it interprets as a long word that it cannot hyphenate, like a net address or a very long title. When this happens, TEX prints an error message like this:

    Overfull \hbox (20.76302pt too wide)

(In TEX, lines are in “horizontal boxes”, hence the term, “hbox”.)

TEX also provides the line number in the Texinfo source file and the text of the offending line, which is marked at all the places that TEX knows how to hyphenate words.

If the Texinfo file has an overfull hbox, you can rewrite the sentence so the overfull hbox does not occur, or you can decide to leave it. A small excursion into the right margin often does not matter and may not even be noticeable.

However, if you do leave an overfull hbox, unless told otherwise, TEX will print a large, ugly, black rectangle beside the line. This is so you will notice the location of the problem if you are correcting a draft. To prevent such a monstrosity from marring your final printout, put the following in the beginning of the Texinfo file on lines of their own, before the @titlepage command:

    @iftex @finalout @end iftex

18.8 Printing “Small” Books

By default, TEX typesets pages for printing in an 8.5 by 11 inch format. However, you can direct TEX to typeset a document in a 7 by 9.25 inch format that is suitable for bound books by inserting the following command on a line by itself at the beginning of the Texinfo file, before the @setchapternewpage command:

    @smallbook

(Since regular sized books are often about 7 by 9.25 inches, this command might better have been called the @regularbooksize command, but it came to be called the @smallbook command by comparison to the 8.5 by 11 inch format.)

The Free Software Foundation distributes printed copies of the GNU Emacs Manual in the “small” book size. See Section 10.6 [@smallexample and @smalllisp], page 87, for information about “example” environments that make it easier to produce examples for a smaller manual.

19 Creating an Info File

There are three commands for formatting an Info file: makeinfo,

texinfo-format-region, and texinfo-format-buffer.

A Texinfo file must possess an @setfilename line near its beginning, otherwise the formatting commands will fail.

For information on installing the Info file in the Info system, see Chapter 20 [Installing an Info File], page 147.

The makeinfo command creates an Info file from a Texinfo source file

more quickly than either the texinfo-format-region or the texinfo-

format-buffer command and provides better error messages. It is recommended. The texinfo-format-… commands are obsolete and are useful only if you cannot run makeinfo, which is written in C.

Since makeinfo is a C program that is independent of Emacs, you do not have to run Emacs to use makeinfo. This means that you can use makeinfo on machines that are too small to run Emacs.

Besides being faster and more helpful, makeinfo automatically fills paragraphs (as if every paragraph were followed by an @refill command).

texinfo-format-region and texinfo-format-buffer do not automati-

cally fill paragraphs. (See Appendix E[Refilling Paragraphs], page 185.)

You can run makeinfo in any one of three ways: from an operating system shell, from a shell inside of Emacs, or, in Texinfo mode in Emacs, by typing a key command.

19.1 Running makeinfo from a Shell

To create an Info file from a Texinfo file, type makeinfo followed by the name of the Texinfo file. Thus, to create the Info file for Bison, type the following at the shell prompt (where ‘%’ is the prompt):

    % makeinfo bison.texinfo

(You can create a shell inside of Emacs by typing M-x shell.)

19.2 Options for makeinfo

The makeinfo command takes several options. Most often, options are used to set the value of the fill column and specify the footnote style. Each command line option is a word preceded by a ‘+’.

For example, you could use the following command to create an Info file for ‘bison.texinfo’ that places footnotes at the end of the node (where ‘%’ is the prompt):

% makeinfo +footnote-style EN bison.texinfo

You can write two or more options in sequence, like this:

makeinfo +paragraph-indent 5 +fill-column 60 . . .

This would indent each paragraph 5 spaces and set the fill column to 60. If you wish to discover which version of makeinfo you are using, type:

% makeinfo +version

The options are:

+paragraph-indent num

Set the paragraph indentation style to num. The value of num is interpreted as follows:

• If the value is zero, do not change the existing indentation at the starts of paragraphs.

• If the value is less than zero, delete any existing indentation.

• If the value is greater than zero, indent each paragraph by

that number of spaces.

+fill-column num

Specify the maximum number of columns in a line; this is the right-hand edge of a line. Paragraphs that are filled will be filled to this width. (Filling is the process of breaking up and connecting lines so that lines are the same length as or shorter than the number specified as the fill-column. Lines are broken between words.)

+footnote-style style

Set the footnote style to style. style should either be ‘MN’ for ‘Make Node’, or ‘EN’ for ‘End Node’. When the footnote style is ‘MN’, makeinfo makes a new node containing the footnotes found in the current node. When the footnote style is ‘EN’, makeinfo places the footnote references at the end of the current node.

+no-split

Suppress the splitting stage of makeinfo. Normally, large output files (where the size is greater than 70k bytes) are split into smaller subfiles, each one approximately 50k bytes. If you specify +no-split, makeinfo will not split up the output file.

+version Print the version number of this copy of makeinfo.

+verbose Cause makeinfo to display messages saying what it is doing. Normally, makeinfo only outputs messages if there are errors or warnings.

+error-limit num

Set the maximum number of errors that makeinfo will print before exiting (on the assumption that continuing would be useless). The default number of errors that can be printed before makeinfo gives up is 100.

+no-pointer-validation

Suppress the pointer-validation phase of makeinfo. Normally, after a Texinfo file is processed, some consistency checks are made to ensure that cross references can be resolved, etc. See Section 19.3 [Pointer Validation], page 141.

+reference-limit num

Set the value of the number of references to a node that makeinfo will make without printing a warning. If a node has more than this number of references in it, makeinfo will make the references but also print a warning.

+no-warn Suppress the output of warning messages. This does not suppress the output of error messages, simply warnings. You might want this if the file you are creating has examples of Texinfo cross references within it, and the nodes that are referenced don’t actually exist.

19.3 Pointer Validation

If you do not suppress pointer-validation (by using the +no-pointer-validation option), makeinfo will check the validity of the final Info file. Mostly, this means ensuring that nodes you have referenced really exist. Here is a complete list of what is checked:

1. If a ‘Next’, ‘Previous’, or ‘Up’ node reference is a reference to a node in the current file and is not an external reference such as to ‘(dir)’, then the referenced node must exist.

2. In every node, if the ‘Previous’ node is different from the ‘Up’ node, then the ‘Previous’ node must also be pointed to by a ‘Next’ node.

3. Every node except the ‘Top’ node must have an ‘Up’ pointer.

4. The node referenced by an ‘Up’ pointer must contain a reference to the current node in some manner other than through a ‘Next’ reference. This includes menu items and cross references.

5. If the ‘Next’ reference of a node is not the same as the ‘Next’ reference of the ‘Up’ reference, then the node referenced by the ‘Next’ pointer must have a ‘Previous’ pointer that points back to the current node. This rule allows the last node in a section to point to the first node of the next chapter.

19.4 Creating Pointers Automatically

The makeinfo program has a feature for automatically creating node pointers for a hierarchically organized file that lacks pointers.

When you take advantage of this feature, you do no have to write the ‘Next’, ‘Previous’, and ‘Up’ pointers after the name of the node. However, you must write a sectioning command, such as @chapter or @section, on the line immediately following the truncated @node line. Also, the node name must appear in a menu that is one or more hierarchical levels above the node’s hierarchical level.

This feature in makeinfo is an alternative to the menu and pointer creation and update commands in Texinfo mode. See Section 2.3 [Updating Nodes and Menus], page 17.

Here is an example of using this feature:

@setfilename example.info
@node Top
@chapter Introduction
@menu
* foo::         the foo node
* bar::         the bar node
@end menu
@node foo
@section Foo
This is the foo node.
@node bar
@section Bar
This is the bar node.
@bye

produces

Info file example.info, -*- Text -*-
produced by Makeinfo,
from input file test.texinfo.

^_
File: example.info,  Node: Top
Introduction
************
* Menu:
* foo::         the foo node
* bar::         the bar node
^_
File: example.info,  Node: foo,  Next: bar,  Up: Top
Foo
===
This is the foo node.
^_
File: example.info,  Node: bar,  Prev: foo,  Up: Top
Bar
===
This is the bar node.

19.5 Running makeinfo inside Emacs

You can run makeinfo in GNU Emacs by using either the makeinforegion or the makeinfo-buffer commands. In Texinfo mode, the commands are bound to C-c C-m C-r and C-c C-m C-b by default.

C-c C-m C-r
makeinfo-region

Format the current region for Info.

C-c C-m C-b
makeinfo-buffer

Format the current buffer for Info.

When you invoke either makeinfo-region or makeinfo-buffer, Emacs

prompts for a file name, offering you name of the visited file as the default.

You can edit the default file name in the minibuffer if you wish, before typing RET to start the makeinfo process.

   The Emacs makeinfo-region and makeinfo-buffer commands run the

makeinfo C program in a temporary shell buffer. If makeinfo finds any errors, Emacs displays the error messages in the temporary buffer.

You can parse the error messages by typing C-x ` (next-error). This causes Emacs to go to and position the cursor on the line in the Texinfo source that makeinfo thinks caused the error. See section ‘Running make or Compilers Generally’ in The GNU Emacs Manual, for more information about using the next-error command.

In addition, you can kill the shell in which the makeinfo command is running or make the shell buffer display its most recent output.

C-c C-m C-k
makeinfo-kill-job
          Kill the shell created by makeinfo-region or makeinfo-buffer.
C-c C-m C-l
makeinfo-recenter-output-buffer

Redisplay the makeinfo shell buffer to display its most recent output.

(Note that the parallel commands for killing and recentering a TEX job are

C-c C-t C-k and C-c C-t C-l.)

You can specify the footnote style that makeinfo should use with the texinfo-footnote-style variable, either ‘EN’ for ‘End Node’ or ‘MN’ for ‘Make Node’.

You can specify the other options with the makeinfo-options variable. You can change the values of these variables with the M-x edit-options

command (see section ‘Editing Variable Values’ in The GNU Emacs Manual), with the M-x set-variable command (see section ‘Examining and Setting Variables’ in The GNU Emacs Manual), or with your ‘.emacs’ initialization file (see section ‘Init File’ in The GNU Emacs Manual).

19.6 The texinfo-format… Commands

In GNU Emacs in Texinfo mode, you can format part or all of a Texinfo file with the texinfo-format-region command. This formats the current region and displays the formatted text in a temporary buffer called ‘*Info

Region*’.

Similarly, you can format a buffer with the texinfo-format-buffer command. This command creates a new buffer and generates the Info file in it. Typing C-x C-s will save the Info file under the name specified by the @setfilename line which must be near the beginning of the Texinfo file.

C-c C-i C-r
texinfo-format-region

Format the current region for Info.

C-c C-i C-b
texinfo-format-buffer

Format the current buffer for Info.

   The texinfo-format-region and texinfo-format-buffer commands

provide some error checking; and other functions provide you with further help in finding formatting errors. These procedures are described in an appendix. See Appendix D[Catching Mistakes], page 175. However, the makeinfo program is faster and produces better error checking.

19.7 Tag Files and Split Files

If a Texinfo file has more than 30,000 bytes, texinfo-format-buffer automatically creates a tag table for its Info file; makeinfo always creates a tag table. With a tag table, Info can jump to new nodes more quickly than it can otherwise.

In addition, if the Texinfo file has more than 70,000 bytes in it, texinfoformat-buffer and makeinfo split the large Info file into shorter indirect subfiles of about 50,000 bytes each. Big files are split into smaller files so that Emacs does not have to make a large buffer to hold the whole of a large Info file; instead, Emacs allocates just enough memory for the small, split off file that is needed at the time. This way, Emacs avoids wasting memory when you run Info. (Before splitting was implemented, Info files were always short and include files were designed as a way to create a single, large printed manual out of the smaller Info files. See Appendix B[Include Files], page 165, for more information. Include files are still used for very large documents, such as The Emacs Lisp Reference Manual, in which each chapter is a separate file.)

When a file is split, Info itself makes use of a shortened version of the original file that contains just the tag table and references to the files that were split off. The split off files are called indirect files.

The split off files have names that are created by appending ‘-1’, ‘-2’, ‘-3’ and so on to the file names specified by the @setfilename command. The shortened version of the original file continues to have the name specified by

@setfilename.

At one stage in writing this document, for example, the Info file was saved as ‘test-texinfo’ and that file looked like this:

Info file: test-texinfo,    -*-Text-*-
produced by texinfo-format-buffer
from file: new-texinfo-manual.texinfo
^_
Indirect:
test-texinfo-1: 102
test-texinfo-2: 50422
test-texinfo-3: 101300
^_^L
Tag table:
(Indirect)
Node: overview^?104
Node: info file^?1271
Node: printed manual^?4853
Node: conventions^?6855

…

(But ‘test-texinfo’ had far more nodes than are shown here.) Each of the

split off, indirect files, ‘test-texinfo-1’, ‘test-texinfo-2’, and ‘test-

texinfo-3’, is listed in this file after the line that says ‘Indirect:’. The tag table is listed after the line that says ‘Tag table:’.

If you are using texinfo-format-buffer to create Info files, you may want to run the Info-validate command. (The makeinfo command does such a good job on its own, you do not need Info-validate.) However, you cannot run the M-x Info-validate node-checking command on indirect files. For information on how to prevent files from being split and how to validate the structure of the nodes, see Section D.5.1 [Using Info-validate], page 180.

20 Installing an Info File

Info files are usually kept in the ‘…/emacs/info’ directory. (You can find the location of this directory by typing C-h i to enter Info and then typing C-x C-f to see the full pathname to the ‘info’ directory.)

The Info tutorial describes how to use Info. (See Info file ‘info’, node ‘Info’, for more information.)

For Info to work, the ‘info’ directory must contain a file that serves as a top level directory for the Info system. By convention, this file is called ‘dir’. The ‘dir’ file is itself an Info file. It contains the top level menu for all the Info files in the system. The menu looks like this:

    * Menu:
    * Info:    (info).     Documentation browsing system.
    * Emacs:   (emacs).    The extensible, self-documenting
                           text editor.
    * Texinfo: (texinfo).  With one source file, make
                           either a printed manual using
                           TeX or an Info file.

…

Each of these menu entries points to the ‘Top’ node of the Info file that is named in parentheses. (The menu entry does not have to specify the ‘Top’ node, since Info goes to the ‘Top’ node if no node name is mentioned. See Section 6.4 [Nodes in Other Info Files], page 57.)

Thus, the ‘Info’ entry points to the ‘Top’ node of the ‘info’ file and the ‘Emacs’ entry points to the ‘Top’ node of the ‘emacs’ file.

In each of the Info files, the ‘Up’ pointer of the ‘Top’ node refers back to the dir file. For example, the node line for the ‘Top’ node of the Emacs manual looks like this:

    File: emacs  Node: Top, Up: (DIR), Next: Distrib

(Note that in this case, the file name is written in upper case letters—it can be written in either upper or lower case.)

20.1 Listing a New Info File

To add a new Info file to your system, add the name to the menu in the ‘dir’ file by editing the ‘dir’ file by hand. Also, put the new Info file in the ‘…/emacs/info’ directory. For example, if you were adding documentation for GDB, you would make the following new entry:

    * GDB: (gdb).           The source-level C debugger.

The first item is the menu item name; it is followed by a colon. The second item is the name of the Info file, in parentheses; it is followed by a period. The third part of the entry is the description of the item.

20.2 Info Files in Other Directories

If an Info file is not in the ‘info’ directory, there are two ways to specify its location:

• Write the menu’s second part as a path name, or;

• Specify an environment variable in your ‘.profile’ or ‘.cshrc’ initialization file. (Only you and others with the same environment variable will be able to find Info files whose location is specified this way.) For example, to reach a test file in the ‘~bob/manuals’ directory, you

could add an entry like this to the menu in the ‘dir’ file:

    * Test: (~bob/manuals/info-test).  Bob's own test file.

In this case, the absolute file name of the ‘info-test’ file is written as the second item of the menu entry.

Alternatively, you may set the ‘INFO_DIRECTORY’ environment variable in your ‘.cshrc’ or ‘.profile’ file. The ‘INFO_DIRECTORY’ environment variable will tell Info where to look.

If you use ‘sh’ or ‘bash’ for your shell command interpreter, you must set the ‘INFO_DIRECTORY’ environment variable in the ‘.profile’ initialization file; but if you use ‘csh’, you must set the variable in the ‘.cshrc’ initialization file. The two files require slightly different command formats.

• In a ‘.cshrc’ file, you could set the ‘INFO_DIRECTORY’ variable as fol-

lows:

         setenv INFO_DIRECTORY .:~bob/manuals:/usr/local/emacs/info

• In a ‘.profile’ file, you would achieve the same effect by writing:

         INFO_DIRECTORY=.:~bob/manuals:/usr/local/emacs/info
         export INFO_DIRECTORY

Either form would cause Info to look first in the current directory, indicated by the ‘.’, then in the ‘~bob/manuals’ directory, and finally in the ‘/usr/local/emacs/info’ directory (which is the usual location for the standard Info directory).

Appendix A @-Command List

Here is an alphabetical list of the @-commands in Texinfo. @* Force a line break. Do not end a paragraph that uses @* with an

@refill command. See Section 16.1 [Line Breaks], page 125.

@. Stands for a period that really does end a sentence. See ⟨unde-

fined⟩ [Insert Period], page ⟨undefined⟩.

@: Indicate to TEX that an immediately preceding period, question mark, exclamation mark, or colon does not end a sentence. Prevent TEX from inserting extra whitespace as it does at the end of a sentence. The command has no effect on the Info file output. See ⟨undefined⟩ [Insert Colon], page ⟨undefined⟩.

@@ Stands for ‘@’. See Section 9.1 [Inserting ‘@’], page 79. @{ Stands for a left-hand brace, {. See Section 9.1 [Inserting @

braces and periods], page 79.

@} Stands for a right-hand brace, }. See Section 9.1 [Inserting @

braces and periods], page 79.

@appendix title

Identify an appendix. The title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.4 [The @unnumbered and @appendix Commands], page 47.

@appendixsec title

Identify an appendix section within an appendix. The section title appears in the table of contents of a printed manual. In Info, the title is underlined with equal signs. See Section 5.7 [Section Commands], page 48.

@appendixsubsec title

Identify an appendix subsection within an appendix. The title appears in the table of contents of a printed manual. In Info, the title is underlined with hyphens. See Section 5.9 [Subsection Commands], page 49.

@appendixsubsubsec title

Identify an appendix subsubsection within a subappendix. The title appears in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.10 [The ‘subsub’ Commands], page 50.

@asis Used following @table and @ftable to print the table’s first column without highlighting (“as is”). See Section 12.3 [Making a Two-column Table], page 98.

@author author’s names Typeset author’s names flushleft and underline it. See Section 3.4.3 [The @title and @author Commands], page 34.

@b{text} Print text in bold font. No effect in Info. See Section 8.2.3 [Fonts], page 76.

@br Force a paragraph break. If used within a line, follow @br with braces. See Section 16.4 [@br], page 127.

@bullet{}

Generate a large round dot, or the closest possible thing to one. See Section 9.2.2 [@bullet{}], page 81.

@bye Terminate TEX processing on the file. TEX does not see any of the contents of the file following the @bye command. See Chapter 4 [Ending a File], page 41.

@c comment Indicate a comment. The rest of the line does not appear in either the Info file or the printed manual. See Section 1.4 [General Syntactic Conventions], page 7.

@center line-of-text Center the line of text following the command. See Section 3.4.2 [@center], page 33.

@chapheading title

Print a chapter-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.5 [@majorheading and @chapheading], page 48.

@chapter title

Identify a chapter. The chapter title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.3 [@chapter], page 47.

@cindex entry

Add entry to the index of concepts. See Section 15.1 [Defining the Entries of an Index], page 119.

@cite{reference} Highlight the name of a book or other reference that lacks a companion Info file. See Section 8.1.8 [@cite], page 74.

@code{sample-code} Highlight text that is an expression, a syntactically complete token of a program, or a program name. See Section 8.1.1 [@code], page 70.

@comment comment

Indicate a comment in Texinfo. The rest of the line does not appear in either the Info file or the printed manual. See Section 1.4 [General Syntactic Conventions], page 7.

@contents

Print a complete table of contents. Not effective in Info, which uses menus instead. See Section 4.2 [Generating a Table of Contents], page 42.

@copyright{}

Generate a copyright symbol. See Section 9.3.2 [@copyright{}], page 81.

@ctrl{ctrl-char} Describe an ASCII control character. Insert actual control character into Info file. See Section 8.1.9 [@ctrl], page 74.

@defcodeindex index-name

Define a new index and its indexing command. Print entries in an @code font. See Section 15.5 [Defining New Indices], page 123.

@defcv category class name Format a description for variables associated with classes in object-oriented programming. Takes three arguments: the category of thing being defined, the class to which it belongs, and its name. See Chapter 13 [Definition Commands], page 101.

@deffn category name arguments… Format a description for functions, interactive commands, and similar entities that may take arguments. Takes as arguments the category of entity being described, the name of this particular entity, and its arguments, if any. See Chapter 13 [Definition Commands], page 101.

@defindex index-name Define a new index and its indexing command. Print entries in a roman font. See Section 15.5 [Defining New Indices], page 123.

@defivar class instance-variable-name Format a description for instance variables in object-oriented programming. The command is equivalent to ‘@defcv {Instance Variable} …’. See Chapter 13 [Definition Commands], page 101.

@defmac macro-name arguments… Format a description for macros. The command is equivalent to ‘@deffn Macro …’. See Chapter 13 [Definition Commands], page 101.

@defmethod class method-name arguments… Format a description for methods in object-oriented programming. The command is equivalent to ‘@defop Method …’. Takes as arguments the name of the class of the method, the name of the method, and its arguments, if any. See Chapter 13 [Definition Commands], page 101.

@defop category class name arguments… Format a description for operations in object-oriented programming. Takes as arguments the overall name of the category of operation, the name of the class of the operation, the name of the operation, and its arguments, if any. See Chapter 13 [Definition Commands], page 101.

@defopt option-name Format a description for user options. The command is equivalent to ‘@defvr {User Option} …’. See Chapter 13 [Definition Commands], page 101.

@defspec special-form-name arguments… Format a description for special forms. The command is equivalent to ‘@deffn {Special Form} …’. See Chapter 13 [Definition Commands], page 101.

@deftp category name-of-type attributes… Format a description for data types. Takes as arguments the category, the name of the type (which is a word like ‘int’ or ‘float’), and then the names of attributes of objects of that type. See Chapter 13 [Definition Commands], page 101.

@deftypefn classification data-type name arguments… Format a description for functions and similar entities that may take arguments and that are typed. Takes as arguments the classification of entity being described, the type, the name of the entity, and its arguments, if any. See Chapter 13 [Definition Commands], page 101.

@deftypefun data-type function-name arguments… Format a description for functions in typed languages. The command is equivalent to ‘@deftypefn Function …’. See Chapter 13 [Definition Commands], page 101.

@deftypevr classification data-type name Format a description for something like a variable in a typed language—an entity that records a value. Takes as arguments the classification of entity being described, the type, and the name of the entity. See Chapter 13 [Definition Commands], page 101.

@deftypevar data-type variable-name Format a description for variables in typed languages. The command is equivalent to ‘@deftypevr Variable …’. See Chapter 13 [Definition Commands], page 101.

@defun function-name arguments… Format a description for functions. The command is equivalent to ‘@deffn Function …’. See Chapter 13 [Definition Commands], page 101.

@defvar variable-name Format a description for variables. The command is equivalent to ‘@defvr Variable …’. See Chapter 13 [Definition Commands], page 101.

@defvr category name Format a description for any kind of variable. Takes as arguments the category of the entity and the name of the entity. See Chapter 13 [Definition Commands], page 101.

@dfn{term}

Highlight the introductory or defining use of a term. See Section 8.1.7 [@dfn], page 74.

@display Indent text, do not fill, do not select a new font. Pair with @end

display. See Section 10.7 [@display], page 87.

@dots{} Stands for an ellipsis: ‘…’. See Section 9.2.1 [@dots{}], page 80.

@emph{text}

Highlight text. See Section 8.2 [Emphasizing Text], page 75.

@enumerate

Begin a numbered list, using @item for each entry. Pair with @end enumerate. See Section 12.2 [@enumerate], page 97.

@equiv{} Indicate the exact equivalence of two forms to the reader with a

special glyph: ‘ ≡ ’. See Section 11.5 [Equivalence], page 93.

@error{} Indicate to the reader with a special glyph that the following text is an error message: ‘ error ’. See Section 11.4 [Error Special Glyph], page 92.

@evenfooting [left] @| [center] @| [right]

Specify footings for even-numbered (left-hand) pages. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@evenheading [left] @| [center] @| [right]

Specify headings for even-numbered (left-hand) pages. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@everyfooting [left] @| [center] @| [right]

Specify footings for every page. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@everyheading [left] @| [center] @| [right]

Specify headings for every page. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@example Indent text, do not fill, select fixed-width font. Pair with @end

example. See Section 10.3 [@example], page 84.

@exdent line-of-text Remove any indentation a line might have. See Section 10.9 [@exdent: Undoing the Indentation of a Line], page 88.

@expansion{}

Indicate the result of a macro expansion to the reader with a special glyph: ‘↦’. See Section 11.2 [expansion], page 91.

@file{filename}

Highlight the name of a file or directory. See Section 8.1.6 [@file], page 73.

@finalout

Prevent TEX from printing large black warning rectangles beside over-wide lines. See Section 18.7 [Overfull Hboxes], page 137.

@findex entry

Add entry to the index of functions. See Section 15.1 [Defining the Entries of an Index], page 119.

@flushleft

Left justify every line but leave the right end ragged. Leave font as is. Pair with @end flushleft. See Section 10.10 [@flushleft and @flushright], page 89.

@flushright

Right justify every line but leave the left end ragged. Leave font as is. Pair with @end flushright. See Section 10.10

[@flushleft and @flushright], page 89.

@footnote{text-of-footnote} Enter a footnote. Footnote text is printed at the bottom of the page by TEX; Info may format in either ‘End Node’ or ‘Make Node’ style. See Chapter 14 [Footnotes], page 117.

@format Like @example or @display, but do not narrow the margins or select the fixed-width font. Pair with @end format. See Section 10.3 [@example], page 84.

@ftable formatting-command Begin a two-column table, using @item for each entry. Automatically enter each of the items in the first column into the index of functions. Pair with @end ftable. The same as @table, except for indexing. See Section 12.3.1 [@ftable], page 99.

@group Hold text together that must appear on one printed page. Pair with @end group. Not relevant to Info. See Section 16.6 [@group], page 127.

@heading title

Print an unnumbered section-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with equal signs. See Section 5.7 [Section Commands], page 48.

@headings on-off-single-double Turn headings on or off, or specify single-sided or double-sided headings for printing. See Section 3.4.6 [The @headings Command], page 36.

@i{text} Print text in italic font. No effect in Info. See Section 8.2.3

[Fonts], page 76.

@ifinfo Begin a stretch of text that will be ignored by TEX when it typesets the printed manual. The text appears only in the Info file. Pair with @end ifinfo. See Chapter 17 [Conditionally Visible Text], page 129.

@iftex Begin a stretch of text that will not appear in the Info file, but will be processed only by TEX. Pair with @end iftex. See Chapter 17 [Conditionally Visible Text], page 129.

@ignore Begin a stretch of text that will not appear in either the Info file or the printed output. Pair with @end ignore. See Section 1.5 [Comments and Ignored Text], page “folio .

@include filename

Incorporate the contents of the file ‘filename’ into the Info file or printed document. See Appendix B[Include Files], page 165.

@inforef{node-name, [entry-name], info-file-name} Make a cross reference to an Info file for which there is no printed manual. See Section 7.6 [Cross references using @inforef], page 67.

\input macro-definitions-file Use the specified macro definitions file. This command is used only in the first line of a Texinfo file to cause TEX to make use of the ‘texinfo’ macro definitions file. The backslash in \input is used instead of an @ because TEX does not properly recognize @ until after it has read the definitions file. See Section 3.2 [The

Texinfo File Header], page 29.

@item Indicate the beginning of a marked paragraph for @itemize and @enumerate; indicate the beginning of the text of a first column entry for @table and @ftable. See Chapter 12 [Lists and Tables], page 95.

@itemize mark-generating-character-or-command Produce a sequence of indented paragraphs, with a mark inside the left margin at the beginning of each paragraph. Pair with @end itemize. See Section 12.1 [@itemize], page 95.

@itemx Like @item but do not generate extra vertical space above the

item text. See Section 12.3.2 [@itemx], page 99.

@kbd{keyboard-characters} Indicate text that consists of characters of input to be typed by users. See Section 8.1.2 [@kbd], page 71.

@key{key-name} Use when key-name is the conventional name for a key on a keyboard. See Section 8.1.3 [@key], page 71.

@kindex entry

Add entry to the index of keys. See Section 15.1 [Defining the Entries of an Index], page 119.

@lisp Indent text, do not fill, select fixed-width font. Pair with @end

lisp. See Section 10.5 [@lisp], page 86.

@majorheading title

Print a chapter-like heading in the text, but not in the table of contents of a printed manual. Generate more vertical whitespace before the heading than the @chapheading command. In Info, the title is underlined with asterisks. See Section 5.5

[@majorheading and @chapheading], page 48.

@menu Mark the beginning of a menu of nodes in Info. No effect in a printed manual. Pair with @end menu. See Section 6.3 [@menu], page 55.

@minus{} Generate a minus sign. See Section 9.4 [@minus], page 81. @need n Start a new page in a printed manual if fewer than n mils (thou-

sandths of an inch) remain on the current page. See Section 16.7 [@need], page 128.

@node name, next, previous, up Define the beginning of a new node in Info, and serve as a locator for references for TEX. See Section 6.2 [@node], page 52.

@noindent

Prevent text from being indented as if it were a new paragraph. See Section 10.4 [@noindent], page 86.

@oddfooting [left] @| [center] @| [right]

Specify footings for odd-numbered (right-hand) pages. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@oddheading [left] @| [center] @| [right]

Specify headings for odd-numbered (right-hand) pages. Not relevant to Info. See Section C.3 [How to Make Your Own Headings], page 171.

@page Start a new page in the printed manual. No effect in Info. See

Section 16.5 [@page], page 127.

@pindex entry

Add entry to the index of programs. See Section 15.1 [Defining the Entries of an Index], page 119.

@point{} Indicate the position of point in a buffer to the reader with a special glyph: ‘⋆’. See Section 11.6 [Indicating Point in a Buffer], page 93.

@print{} Indicate printed output to the reader with a special glyph: ‘ ⊣ ’.

See Section 11.3 [Print Special Glyph], page 92.

@printindex index-name

Print an alphabetized two-column index in a printed manual or generate an alphabetized menu of index entries for Info. See Section 4.1 [Printing Indices & Menus], page 41.

@pxref{node-name, [entry], [topic], [info-file], [manual]} Make a reference that starts with a lower case ‘see’ in a printed manual. Use within parentheses only. Do not follow command with a punctuation mark. The Info formatting commands automatically insert terminating punctuation as needed, which is why you do not need to insert punctuation. Only the first argument is mandatory. See Section 7.5 [@pxref], page 66.

@quotation

Fill and indent text, print it in a roman font. This indicates text that is quoted from another real or imaginary work.. Write command on a line of its own. Pair with @end quotation. See Section 10.2 [@quotation], page 84.

@r{text} Print text in roman font. No effect in Info. See Section 8.2.3

[Fonts], page 76.

@ref{node-name, [entry], [topic], [info-file], [manual]} Make a reference. In the printed manual, the reference does not start with a ‘See’. Follow command with a punctuation mark. Only the first argument is mandatory. See Section 7.4 [@ref], page 65.

@refill In Info, refill the paragraph after all the other processing has been done. No effect on TEX, which always refills, and not needed on an Info file formatted with makeinfo, which also automatically refills. See Appendix E[Refilling Paragraphs], page 185.

@result{}

Indicate the result of an expression to the reader with a special glyph: ‘⇒’. See Section 11.1 [@result], page 91.

@samp{text}

Highlight text that is a literal example of a sequence of characters. Used for single characters, for statements and often for entire shell commands. See Section 8.1.4 [@code], page 72.

@sc{text} Set text in the printed output in the small caps font and set text in the Info file in uppercase letters. See Section 8.2.2 [Smallcaps], page 76.

@section title

Identify a section within a chapter. In a printed manual, the section title is numbered and appears in the table of contents. In Info, the title is underlined with equal signs. See Section 5.6 [@section], page 48.

@setchapternewpage on-off-odd

Specify whether chapters start on new pages, and if so, whether on odd-numbered (right-hand) new pages. See Section 3.2.5

[@setchapternewpage], page 31.

@setfilename info-file-name Provide a name for the Info file. See Section 1.4 [General Syntactic Conventions], page 7.

@settitle title

Provide a title for page headers in a printed manual. See Section 1.4 [General Syntactic Conventions], page 7.

@shortcontents

Print a short table of contents. Not relevant to Info, which does not have a table of contents, but uses menus. A synonym for @summarycontents. See Section 4.2 [Generating a Table of Contents], page 42.

@smallbook

Cause TEX to produce a printed manual in a 7 by 9.25 inch format rather than the regular 8.5 by 11 inch format. See Section 18.8 [Printing Small Books], page 137. Also, see Section 10.6

[@smallexample and @smalllisp], page 87.
@smallexample

Indent text, do not fill, select fixed-width font. In @smallbook format, print text in a smaller font than with the @example environment. Pair with @end smallexample. See Section 10.6

[@smallexample and @smalllisp], page 87.
@smalllisp

Indent text, do not fill, select fixed-width font. In @smallbook format, print text in a smaller font. Pair with @end smalllisp. See Section 10.6 [@smallexample and @smalllisp], page 87.

@sp n Skip n blank lines. See Section 16.3 [@sp], page 126.

@strong text

Emphasize text by typesetting it in a bold font for the printed manual and by surrounding it with asterisks for Info. See Section 8.2.1 [@emph{text} and @strong{text}], page 75.

@subheading title

Print an unnumbered subsection heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with hyphens. See Section 5.9 [@unnumberedsubsec

@appendixsubsec @subheading], page 49.
@subsection title

Identify a subsection within a section. In a printed manual, the subsection title is numbered and appears in the table of contents. In Info, the title is underlined with hyphens. See Section 5.8

[@subsection], page 49.
@subsubheading title

Print an unnumbered subsubsection heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.10 [The ‘subsub’ Commands], page 50.

@subsubsection title

Identify a subsubsection within a subsection. In a printed manual, the subsubsection title is numbered and appears in the table of contents. In Info, the title is underlined with periods. See Section 5.10 [The ‘subsub’ Commands], page 50.

@subtitle title

In the printed manual, set a subtitle in a normal sized font flush to the right-hand side of the page. Not relevant to Info, which does not have title pages. See Section 3.4.3 [@title @subtitle and @author Commands], page 34.

@summarycontents

Print a short table of contents. Not relevant to Info, which does not have a table of contents, but uses menus. A synonym for @shortcontents. See Section 4.2 [Generating a Table of Contents], page 42.

@syncodeindex from-index into-index Merge the index named in the first argument into the index named in the second argument, printing the entries from the first index in @code font. See Section 15.4 [Combining Indices], page 121.

@synindex from-index into-index Merge the index named in the first argument into the index named in the second argument. Sets from-index in the same (roman) font as into-index See Section 15.4 [Combining Indices], page 121.

@t{text} Print text in a fixed-width font. No effect in Info. See Sec-

tion 8.2.3 [Fonts], page 76.

@table formatting-command Begin a two column table, using @item for each entry. Write each first column entry on the same line as @item. First column entries are printed in the font resulting from formattingcommand. See Section 12.3 [Making a Two-column Table], page 98. Also see Section 12.3.1 [@ftable], page 99, and Section 12.3.2 [@itemx], page 99.

@TeX\{\} Stands for the logo TEX. See Section 9.3 [Inserting TEX and \begin{matrix}c \\ ◯ \end{matrix}],

page 81.

@tex Enter TEX completely. Pair with @end tex. See Section 17.1

[Using Ordinary TEX Commands], page 129.

@thischapter

In a heading or footing, stands for the number and name of the current chapter, in the format ‘Chapter 1: First Chapter’. See Section C.3 [How to Make Your Own Headings], page 171.

@thischaptername

In a heading or footing, stands for the name of the current chapter. See Section C.3 [How to Make Your Own Headings], page 171.

@thisfile

In a heading or footing, stands for the name of the current @include file. Does not insert anything if there is no @include file. See Section C.3 [How to Make Your Own Headings], page 171.

@thistitle

In a heading or footing, stands for the name of the document, as specified by the @settitle command. See Section C.3 [How to Make Your Own Headings], page 171.

@tindex entry

Add entry to the index of data types. See Section 15.1 [Defining the Entries of an Index], page 119.

@title title

In the printed manual, set a title flush to the left-hand side of the page in a larger than normal font and underline it with a black rule. Not relevant to Info, which does not have title pages. See section ‘The @title’ in and @author Commands.

@titlefont{text}

In the printed manual, print text in a larger than normal font. Not relevant to Info, which does not have title pages. See Section 3.4.2 [The @titlefont @center and @sp Commands], page 33.

@titlepage

Indicate to Texinfo the beginning of the title page. Write command on a line of its own. Pair with @end titlepage. Nothing within this environment appears in Info. See Section 3.4.1

[@titlepage], page 33.

@today{} Stands for the current date, in ‘1 Jan 1900’ style. See Section C.3

[How to Make Your Own Headings], page 171.

@unnumbered title

In a printed manual, identify a chapter that appears without chapter numbers of any kind. The title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.4 [unnumbered and appendix], page 47.

@unnumberedsec title

In a printed manual, identify a section that appears without section numbers of any kind. The title appears in the table of contents of a printed manual. In Info, the title is underlined with equal signs. See Section 5.7 [Section Commands], page 48.

@unnumberedsubsec title

In the printed manual, identify an unnumbered subsection within a chapter. The title appears in the table of contents of a printed manual. In Info, the title is underlined with hy-

phens. See Section 5.9 [@unnumberedsubsec @appendixsubsec
@subheading], page 49.
@unnumberedsubsubsec title

In the printed manual, identify an unnumbered subsubsection within a chapter. The title appears in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.10 [The ‘subsub’ Commands], page 50.

@var{metasyntactic-variable} Highlight a metasyntactic variable, which is something that stands for another piece of text. Thus, in this entry, the word metasyntactic-variable is highlighted with @var. See Section 8.1.5 [@var{metasyntactic-variable}], page 73.

@vindex entry

Add entry to the index of variables. See Section 15.1 [Defining the Entries of an Index], page 119.

@vskip amount

In the printed manual, insert whitespace so as to push text on the remainder of the page towards the bottom of the page. Used in formatting the copyright page with the argument ‘0pt plus 1filll’. (Note spelling of ‘filll’.) @vskip may not be used except in contexts ignored for Info. See Section 3.4.4 [The Copyright Page and Printed Permissions], page 35.

@w{text} Prevent text from being split across two lines. Do not end a paragraph that uses @w with an @refill command. In the Texinfo file, keep text on one line. See Section 16.2 [@w], page 126.

@xref{node-name, [entry], [topic], [info-file], [manual]} Make a reference that starts with ‘See’ in the printed manual. Follow command with a punctuation mark. Only the first argument is mandatory. See Section 7.3 [@xref], page 61.

Appendix B Include Files

When TEX or an Info formatting command sees an @include command in a Texinfo file, it processes the contents of the file named by the command and incorporates them into the dvi or Info file being created. Index entries from the included file are incorporated into the indices of the output file.

B.1 How to Use Include Files

To include another file within a Texinfo file, write the @include command at the beginning of a line and follow it on the same line by the name of a file to be included. For example:

    @include chap47.texinfo

An included file should simply be a segment of text that you expect to be included as-is into the overall or outer Texinfo file; it should not contain the standard beginning and end parts of a Texinfo file. In particular, you should not start an included file with a line saying ‘\input texinfo’; if you do, that phrase is inserted into the output file as is. Likewise, you should not end an included file with an @bye command; that command will stop TEX processing immediately.

In the past, you did need to write an @setfilename line at the beginning of an included file, but no longer; however, it does not matter whether you write such a line. If an @setfilename line exists in an included file, it is ignored.

Conventionally, an included file begins with an @node line that is followed by an @chapter line. Each included file is one chapter. This makes it easy to use the regular node and menu creating and updating commands to create the node pointers and menus within the included file. However, the regular node and menu creating and updating commands do not work with multiple Texinfo files. Thus you cannot use these commands to fill in the ‘Next’, ‘Previous’, and ‘Up’ pointers of the @node line that begins the included file. Also, you cannot use the regular commands to create a master menu for the whole file. Either you must insert the menus and the ‘Next’, ‘Previous’, and ‘Up’ pointers by hand, or you must use the texinfo-multiple-files-update command that is designed for @include files.

B.2 texinfo-multiple-files-update
   The texinfo-multiple-files-update command can be called without

any arguments, or with a non-nil ‘make-master-menu’ argument (a prefix argument, if called interactively), or with a non-nil ‘update-everything’ argument (a numeric prefix argument, if called interactively). To use the com-

mand, type M-x texinfo-multiple-files-update or C-u M-x texinfo-. . .

or else C-u 4 M-x texinfo-. . .. You will be prompted for the name of the outer file.

   Without an argument, the texinfo-multiple-files-update command

creates or updates the ‘Next’, ‘Previous’, and ‘Up’ pointers of the first @node line in each file included in an outer or overall Texinfo file. It also creates or updates the ‘Top’ level node pointers of the outer or overall file. In addition, it creates or updates a main menu in the outer file.

With a non-nil value for the optional ‘make-master-menu’ argument (a prefix argument, if called interactively), the texinfo-multiple-files-update command creates and inserts a master menu in the outer file. The master menu is made from all the menus in all the included files.

Finally, with a non-nil ‘update-everything’ argument (a numeric prefix argument, if called interactively), the texinfo-multiple-files-update command first updates all the menus and all the ‘Next’, ‘Previous’, and ‘Up’ pointers of all the included files before creating and inserting a master menu in the outer file. This is similar to invoking texinfo-master-menu with an argument when you are working with just one file.

B.3 Sample File with @include

   If you plan to use the texinfo-multiple-files-update command, the

outer Texinfo file that lists included files within it should contain nothing but the beginning and end parts of a Texinfo file, and a number of @include commands listing the included files. It should not even include indices, which should be listed in an included file of their own.

Moreover, each of the included files must contain exactly one highest level node (conventionally, an @chapter node or equivalent), and this node must be the first node in the included file. Furthermore, each of these highest level nodes in each included file must be at the same hierarchical level in the file structure. Usually, each is an @chapter node, an @appendix, or an @unnumbered node. Thus, normally, each included file contains one, and only one, chapter or equivalent-level node.

The outer file should not contain any nodes besides the single ‘Top’ node.

The texinfo-multiple-files-update command will not process them.

Here is an example of a outer Texinfo file with @include files within it before running texinfo-multiple-files-update, which would insert a main or master menu:

\input texinfo @c -*-texinfo-*-
@setfilename  include-example.info
@settitle Include Example
@setchapternewpage odd
@titlepage
@sp 12
@center @titlefont{Include Example}
@sp 2
@center by Whom Ever
@page
Copyright @copyright{} 1990 Free Software Foundation, Inc.
@end titlepage
@node Top, First, (dir), (dir)
@ifinfo
@chapter Master Menu
@end ifinfo
@include foo.texinfo
@include bar.texinfo
@include index.texinfo
@summarycontents
@contents
@bye

An included file, such as ‘foo.texinfo’, might look like this:

@node First, Second, , Top
@chapter First Chapter
Contents of first chapter . . .

The full contents of ‘index.texinfo’ might be as simple as this:

@unnumbered Concept Index, , Second, Top
@printindex cp

The Texinfo source for the GNU Emacs Lisp Reference Manual consists of an outer Texinfo file, ‘lispref.texinfo’ with numerous @include files within it.

B.4 Evolution of Include Files

When Info was first created, it was customary to create many small Info files on one subject. Each Info file was formatted from its own Texinfo source file. This custom meant that Emacs did not need to make a large buffer to hold the whole of a large Info file when someone wanted information; instead, Emacs allocated just enough memory for the small Info file that contained the particular information sought. This way, Emacs could avoid wasting memory.

References from one file to another were made by referring to the file name as well as the node name. (See Section 6.4 [Referring to Other Info Files], page 57. Also, see Section 7.3.4 [@xref with Four Arguments], page 63.)

Include files were designed primarily as a way to create a single, large printed manual out of several smaller Info files. In a printed manual, all the references were within the same document, so TEX could automatically determine the references’ page numbers. The Info formatting commands used include files only for creating joint indices; each of the individual Texinfo files had to be formatted for Info individually. (Each, therefore, required its

own @setfilename line.)

However, because large Info files are now split automatically, small Texinfo files are no longer strictly necessary and they are used infrequently.

Nowadays, multiple Texinfo files are used mostly for large documents, such as the GNU Emacs Lisp Reference Manual, and for projects in which several different people write different sections of a document simultaneously.

In addition, the Info formatting commands have been extended to work with the @include command so as to create a single large Info file that is split if necessary into smaller files. This means that you can write menus and cross references without naming the different Texinfo files.

Appendix C Page Headings

Most printed manuals contain headings along the top of every page except the title and copyright pages. Some manuals also contain footings. (Headings and footings have no meaning to Info, which is not paginated.)

Texinfo provides different heading formats for manuals that are printed on one side of each sheet of paper and for manuals that are printed on both sides of the paper. Usually, you will use one or other of these standard formats, but you can specify your own format, if you wish.

In addition, you can specify whether chapters should begin on a new page, or merely continue the same page as the previous chapter; and if a chapter begins on a new page, you can specify whether that page must be an odd-numbered page.

By convention, a book is printed on both sides of each sheet of paper. When you open a book flat, the right-hand page is odd-numbered, and chapters begin on right-hand pages—a preceding left-hand page is left blank if necessary. Reports, however, are often printed on just one side of paper, and chapters begin on a fresh page immediately following the end of the preceding chapter. In short or informal reports, chapters often do not begin on a new page at all, but are separated from the preceding text by a small amount of whitespace.

The @setchapternewpage command controls whether chapters begin on new pages, and whether one or other of the standard heading formats is used. In addition, Texinfo has several heading and footing commands that you can use to generate your own heading and footing formats.

In Texinfo, headings and footings are single lines at the tops and bottoms of pages; you cannot create multiline headings or footings. Each header or footer line is divided into three parts: a left part, a middle part, and a right part. Any part, or a whole line, may be left blank. Text for the left part of a header or footer line is set flushleft; text for the middle part is centered; and, text for the right part is set flushright.

C.1 Standard Heading Formats

Texinfo provides two standard heading formats, one for manuals printed on one side of each sheet of paper, and the other for manuals printed on both sides of the paper.

By default, nothing is specified for the footing of a Texinfo file, so the footings remains blank.

The standard format for single-sided printing consists of a header line in which the left-hand part contains name of the chapter, the central part is blank, and the right-hand part contains the page number.

The single-sided page looks like this:

       _______________________
      |                       |
      | chapter   page number |
      |                       |
      | Start of text ...     |
      |                       |
      |                       |

The standard format for two-sided printing depends on whether the page number is even or odd. By convention, even-numbered pages are on the leftand odd-numbered pages are on the right. (TEX will adjust the widths of the left- and right-hand margins. Usually, widths are correct, but during double-sided printing, it is wise to check that pages will bind properly— sometimes a printer will produce output in which the even-numbered pages have a larger left-hand margin thant the odd-numbered pages.)

In the standard double-sided format, the left part of the left-hand (evennumbered) page contains the page number, the central part is blank and the right part contains the title (specified by the @settitle command). The left part of the right-hand (odd-numbered) page contains the name of the chapter, the central part is blank, and the right part contains the page number.

The two pages, side by side as in an open book, look like this:

       ______________________________________________
      |                       |   |                       |
      | page number     title |   | chapter   page number |
      |                       |   |                       |
      | Start of text ...     |   | Start of text ...     |
      |                       |   |                       |
      |                       |   |                       |

The chapter name is preceded by the word ‘Chapter’, the chapter number and a colon. This makes it easier to keep track of where you are in the manual.

C.2 Specifying the Type of Heading

TEX does not begin to generate page headings for a standard Texinfo file until it reaches the @end titlepage command. Thus, the title and copyright pages are not numbered. The @end titlepage command causes TEX to begin to generate page headings according to a standard format specified by the @setchapternewpage command that precedes the @titlepage section.

There are four possibilities:

No @setchapternewpage command

Cause TEX to specify the single-sided heading format, with chapters on new pages. This is the same as @setchapternewpage on.

@setchapternewpage on

Specify the single-sided heading format, with chapters on new pages.

@setchapternewpage off

Cause TEX to start new chapters on the same page as the the last page of the preceding chapter, after skipping some vertical whitespace. It also causes TEX to typeset for double-sided printing.

@setchapternewpage odd

Specify the double-sided heading format, with chapters on new pages.

Texinfo lacks an @setchapternewpage even command.

C.3 How to Make Your Own Headings

You can use the standard headings provided with Texinfo or specify your own.

Texinfo provides six commands for specifying headings and footings. The @everyheading and @everyfooting commands generate headers and footers that are the same for both even- and odd-numbered pages. The @evenheading and @evenfooting commands generate headers and footers for even-numbered (left-hand) pages; and the @oddheading and @oddfooting commands generate headers and footers for odd-numbered (right-hand) pages.

Write custom heading specifications in the Texinfo file immediately after the @end titlepage command. Enclose your specifications between @iftex

and @end iftex commands since the texinfo-format-buffer command

may not recognize them. Also, you must cancel the predefined heading commands with the @headings off command before defining your own specifications.

Here is how to tell TEX to place the chapter name at the left, the page number in the center, and the date at the right of every header for both even- and odd-numbered pages:

    @iftex
    @headings off
    @everyheading @thischapter @| @thispage @| @today
    @end iftex

You need to divide the left part from the central part and the central part from the right had part by inserting ‘@|’ between parts. Otherwise, the specification command will not be able to tell where the text for one part ends and the next part begins.

Each part can contain text or ‘@this…’ commands. The text is printed as if the part were within an ordinary paragraph in the body of the page. The ‘@this…’ commands replace themselves with the page number, date, chapter name, or whatever.

Here are the six specification commands:

@everyheading left @| center @| right
@everyfooting left @| center @| right

The ‘every’ commands specify the format for both even- and odd-numbered pages. These commands are for documents that are printed on one side of each sheet of paper, or for documents in which you want symmetrical headers or footers.

@evenheading left @| center @| right
@oddheading left @| center @| right
@evenfooting left @| center @| right
@oddfooting left @| center @| right

The ‘even’ and ‘odd’ commands specify the format for evennumbered pages and odd-numbered pages. These commands are for books and manuals that are printed on both sides of each sheet of paper.

Use the ‘@this…’ commands to provide the names of chapters and sections and the page number. You can use the ‘@this…’ commands in the left, center, or right portions of headers and footers, or anywhere else in a Texinfo file so long as they are between @iftex and @end iftex commands.

Here are the ‘@this…’ commands:

@thispage

Expands to the current page number.

@thischapter

Expands to the name of the current chapter.

@thistitle

Expands to the name of the document, as specified by the

          @settitle command.
@thisfile

For @include files only: expands to the name of the current @include file. If the current Texinfo source file is not an @include file, this command has no effect. This command does not provide the name of the current Texinfo source file unless it is an @include file. (See Appendix B[Include Files], page 165, for more information about @include files.)

You can also use the @today{} command, which expands to the current date, in ‘1 Jan 1900’ format.

Other @-commands and text are printed in a header or footer just as if they were in the body of a page. It is useful to incorporate text, particularly when you are writing drafts:

    @iftex
    @headings off
    @everyheading @emph{Draft!} @| @thispage @| @thischapter
    @everyfooting @| @| Version: 0.27: @today{}
    @end iftex

Beware of overlong titles: they may overlap another part of the header or footer and blot it out.

172 Texinfo 2.0

Appendix D Formatting Mistakes

Besides mistakes in the content of your documentation, there are two kinds of mistake you can make with Texinfo: you can make mistakes with @-commands, and you can make mistakes with the structure of the nodes and chapters.

Emacs has two tools for catching the @-command mistakes and two for catching structuring mistakes.

For finding problems with @-commands, you can run TEX or a region formatting command on the region that has a problem; indeed, you can run these commands on each region as you write it.

For finding problems with the structure of nodes and chapters, you can

use C-c C-s (texinfo-show-structure) (and the related occur command)

and you can use the M-x Info-validate command.

The makeinfo program does an excellent job of catching errors and

reporting them—far better than texinfo-format-region or texinfo-

format-buffer. In addition, the various functions for automatically creating and updating node pointers and menus remove many opportunities for human error.

If you can, use the updating commands to create and insert pointers and menus. These prevent many errors. Then use makeinfo (or its Texinfo mode manifestations, makeinfo-region and makeinfo-buffer) to format your file and check for other errors. This is the best way to work with Texinfo. But if you cannot use makeinfo, or your problem is very puzzling, then you may want to use the tools described in this appendix.

D.1 Catching Errors with Info Formatting

After you have written part of a Texinfo file, you can use the texinfoformat-region or the makeinfo-region command to see whether the region formats properly.

Most likely, however, you are reading this section if you for some reason cannot use the makeinfo-region command; therefore, the rest of this section presumes that you are using texinfo-format-region.

If you have made a mistake with an @-command, texinfo-format-region will stop processing at or after the error and display an error message. To see where in the buffer the error occurred, switch to the ‘*Info Region*’ buffer; the cursor will be in a position that is after the location of the error. Also, the text will not be formatted after the place where the error occurred (or more precisely, where it was detected).

For example, if you accidentally end a menu with the command @end menus with an ‘s’ on the end, instead of with @end menu, you will see an error message that says:

    @end menus is not handled by texinfo

The cursor will stop at the point in the buffer where the error occurs, or not long after it. The buffer will look like this:

    ---------- Buffer: *Info Region* ----------
    * Menu:
    * Using texinfo-show-structure::  How to use
                                      `texinfo-show-structure'
                                      to catch mistakes.
    * Running Info-Validate::         How to check for
                                      unreferenced nodes.
    @end menus

⋆

    ---------- Buffer: *Info Region* ----------

The texinfo-format-region command sometimes provides slightly odd error messages. For example,

    (@xref{Catching Mistakes, for more info.)

In this case, texinfo-format-region detects the missing closing brace but displays a message that says ‘Unbalanced parentheses’ rather than ‘Unbalanced braces’. This is because the formatting command looks for mismatches between braces as if they were parentheses.

Sometimes texinfo-format-region fails to detect mistakes. For example, in the following, the closing brace is swapped with the closing parenthesis:

    (@xref{Catching Mistakes), for more info.}

Formatting produces:

    (*Note for more info.: Catching Mistakes)

The only way for you to detect this error is to realize that the reference should have looked like this:

    (*Note Catching Mistakes::, for more info.)

D.2 Catching Errors with TEX Formatting

You can also catch mistakes when you format a file with TEX.

Usually, you will want to do this after you have run texinfo-format-buffer (or, better, makeinfo-buffer) on the same file, because texinfoformat-buffer sometimes displays error messages that make more sense than TEX. (See Section D.1 [Debugging with Info], page 175, for more information.)

For example, TEX was run on a Texinfo file, part of which is shown here:

---------- Buffer: texinfo.texinfo ----------
name of the texinfo file as an extension.  The
@samp{??} are `wildcards' that cause the shell to
substitute all the raw index files.  (@xref{sorting
indices, for more information about sorting
indices.)@refill
---------- Buffer: texinfo.texinfo ----------

(The cross reference lacks a closing brace.) TEX produced the following output, after which it stopped:

---------- Buffer: *TeX-shell* ----------
Runaway argument?
{sorting indices, for more information about sorting
indices.) @refill @ETC.
! Paragraph ended before \xref was complete.
<to be read again>
                   \par
l.27
?
---------- Buffer: *TeX-shell* ----------

In this case, TEX produced an accurate and understandable error message:

Paragraph ended before \xref was complete.

(Note, however, that TEX translated the ‘@’ into a ‘\’.) Also, ‘\par’ is an internal TEX command of no relevance to Texinfo.)

Unfortunately, TEX is not always so helpful, and sometimes you have to be truly a Sherlock Holmes to discover what went wrong.

In any case, if you run into a problem like this, you can do one of two things.

1. You can tell TEX to continue running and to ignore errors as best it can by typing r RET at the ‘?’ prompt. This is often the best thing to do. However, beware: the one error may produce a cascade of additional error messages as its consequences are felt through the rest of the file. (To stop TEX when it is producing such an avalanch of error messages, type C-d (or C-c C-d, if you running a shell inside of Emacs.))

2. You can tell TEX to stop this run by typing x RET at the ‘?’ prompt. Sometimes TEX will format a file without producing error messages even

though there is a problem. This usually occurs if a command is not ended but TEX is able to continue processing anyhow. For example, if you fail to end an itemized list with the @end itemize command, TEX will write a dvi file that you can print out. The only error message that TEX will give you is the somewhat mysterious comment that

    (\end occurred inside a group at level 1)

However, if you print the dvi file, you will find that the text of the file that follows the itemized list is entirely indented as if it were part of the last item in the itemized list. The error message is the way TEX says that it expected to find an @end command somewhere in the file; but that it could not determine where it was needed.

Another source of notoriously hard-to-find errors is a missing @end group command. If you ever are stumped by incomprehensible errors, look for a missing @end group command first.

If you do not have the header lines in the file, TEX may stop in the beginning of its run and display output that looks like the following. The ‘*’ indicates that TEX is waiting for input.

    This is TeX, Version 2.0 for Berkeley UNIX
    (preloaded format=plain-cm 87.10.25)
    (#tz-bar-a02987.tex [1])
    *

In this case, simply type \end RET after the asterisk. Then put the header lines into the Texinfo file and run the TEX command again.

D.3 Using texinfo-show-structure

It is not always easy to keep track of the nodes, chapters, sections and subsections of a Texinfo file. This is especially true if you are revising or adding to a Texinfo file that someone else has written.

In GNU Emacs, in Texinfo mode, the texinfo-show-structure command lists all the lines that begin with the @-commands that specify the structure: @chapter, @section, @appendix, and so on. With an argument (prefix, if interactive), the command also shows the @node lines. The texinfo-show-structure command is bound to C-c C-s in Texinfo mode, by default.

The lines are displayed in a buffer called the ‘*Occur*’ buffer. For example, when texinfo-show-structure was run on an earlier version of this appendix, it produced the following:

     Lines matching "^@\\(chapter \\|sect\\|sub\\|unnum\\|major\\|
     heading \\|appendix\\)" in buffer texinfo.texinfo.
       4:@appendix Formatting Mistakes
      52:@appendixsec Catching Errors with Info Formatting
     222:@appendixsec Catching Errors with @TeX{} Formatting
     338:@appendixsec Using @code{texinfo-show-structure}
     407:@appendixsubsec Using @code{occur}
     444:@appendixsec Finding Badly Referenced Nodes
     513:@appendixsubsec Running @code{Info-validate}
     573:@appendixsubsec Splitting a File Manually

This means that lines 4, 52, and 222 begin with the @appendix, @appendixsec, and @appendixsec commands respectively. If you move your cursor into the ‘*Occur*’ window, you can position the cursor over one of

the lines and use the C-c C-c command (occur-mode-goto-occurrence),

to jump to the corresponding spot in the Texinfo file. See section ‘Using Occur’ in The GNU Emacs Manual, for more information about occur-

mode-goto-occurrence.

The first line in the ‘*Occur*’ window describes the regular expression specified by texinfo-heading-pattern. This regular expression is the pattern that texinfo-show-structure looks for. See section ‘Using Regular Expressions’ in The GNU Emacs Manual, for more information.

When you invoke the texinfo-show-structure command, Emacs will display the structure of the whole buffer. If you want to see the structure of just a part of the buffer, of one chapter, for example, use the C-x n (narrowto-region) command to mark the region. (See section ‘Narrowing’ in The GNU Emacs Manual.) This is how the example used above was generated. (To see the whole buffer again, use C-x w (widen).)

If you call texinfo-show-structure with a prefix argument by typing Cu C-c C-s, it will list lines beginning with @node as well as the lines beginning with the @-sign commands for @chapter, @section, and the like.

You can remind yourself of the structure of a Texinfo file by looking at the list in the ‘*Occur*’ window; and if you have mis-named a node or left out a section, you can correct the mistake.

D.4 Using occur

Sometimes the texinfo-show-structure command produces too much information. Perhaps you want to remind yourself of the overall structure of a Texinfo file, and are overwhelmed by the detailed list produced by texinfo-show-structure. In this case, you can use the occur command directly. To do this, type

    M-x occur

and then, when prompted, type a regexp, a regular expression for the pattern you want to match. (See section ‘Regular Expressions’ in The GNU Emacs Manual.) The occur command works from the current location of the cursor in the buffer to the end of the buffer. If you want to run occur on the whole buffer, place the cursor at the beginning of the buffer.

For example, to see all the lines that contain the word ‘@chapter’ in them, just type ‘@chapter’. This will produce a list of the chapters. It will also list all the sentences with ‘@chapter’ in the middle of the line.

If you want to see only those lines that start with the word ‘@chapter’, type ‘^@chapter’ when prompted by occur. If you want to see all the lines that end with a word or phrase, end the last word with a ‘$’; for example, ‘catching mistakes$’. This can be helpful when you want to see all the nodes that are part of the same chapter or section and therefore have the same ‘Up’ pointer.

See section ‘Using Occur’ in The GNU Emacs Manual, for more information.

D.5 Finding Badly Referenced Nodes

You can use the Info-validate command to check whether any of the ‘Next’, ‘Previous’, ‘Up’ or other node pointers fail to point to a node. This command checks that every node pointer points to an existing node. The Info-validate command works only on Info files, not on Texinfo files.

The makeinfo program validates pointers automatically, so you do not need to use the Info-validate command if you are using makeinfo. You only may need to use Info-validate if you are unable to run makeinfo and instead must create an Info file using texinfo-format-region or texinfoformat-buffer, or if you write an Info file from scratch.

D.5.1 Running Info-validate

To use Info-validate, visit the Info file you wish to check and type:

    M-x Info-validate

(Note that the Info-validate command requires an upper case ‘I’.)

If your file is ok, you will receive a message that says “File appears valid”. However, if you have a pointer that does not point to a node, error messages will be displayed in a buffer called ‘*problems in info file*’.

For example, Info-validate was run on a test file that contained only the first node of this manual. One of the messages said:

    In node "Overview", invalid Next: Texinfo Mode

This meant that the node called ‘Overview’ had a ‘Next’ pointer that did not point to anything (which was true in this case, since the test file had only one node in it).

Now suppose we add a node named ‘Texinfo Mode’ to our test case but we don’t specify a ‘Previous’ for this node. Then we will get the following error message:

    In node "Texinfo Mode", should have Previous: Overview

This is because every ‘Next’ pointer should be matched by a ‘Previous’ (in the node where the ‘Next’ points) which points back.

Info-validate also checks that all menu items and cross references point to actual nodes.

Note that Info-validate requires a tag table and does not work with files that have been split. (The texinfo-format-buffer command automatically splits files larger than 100,000 bytes.) In order to use Info-validate on a large file, you must run texinfo-format-buffer with an argument so that it does not split the Info file; and you must create a tag table for the unsplit file.

D.5.2 Creating an Unsplit File

You can run Info-validate only on a single Info file that has a tag table. The command will not work on the indirect subfiles that are generated when a master file is split. If you have a large file (longer than 70,000 bytes

or so), you need to run the texinfo-format-buffer or makeinfo-buffer

command in such a way that it does not create indirect subfiles. You will also need to create a tag table for the Info file. After you have done this, you can run Info-validate and look for badly referenced nodes.

The first step is to create an unsplit Info file. To prevent texinfoformat-buffer from splitting a Texinfo file into smaller Info files, give a

prefix to the M-x texinfo-format-buffer command:

    C-u  M-x texinfo-format-buffer

or else

    C-u C-c C-i C-b

When you do this, Texinfo will not split the file and will not create a tag table for it.

D.5.3 Tagifying a File

After creating an unsplit Info file, you must create a tag table for it. Visit the file you wish to tagify and type:

    M-x Info-tagify

(Note the upper case I in Info-tagify.) This creates a file with a tag table that you can validate.

The third step is to validate the file:

    M-x Info-validate

(Note the upper case I in Info-validate.) In brief, the steps are:

    C-u M-x texinfo-format-buffer
    M-x Info-tagify
    M-x Info-validate

After you have validated the node structure, you can rerun texinfoformat-buffer in the normal way so it will construct the tag table and split the file automatically or, you can make the tag table and split the file manually.

D.5.4 Splitting a File Manually

You should split a large file or else let the texinfo-format-buffer or makeinfo-buffer command do it for you automatically. (Generally you will let one of the formatting commands do this job for you. See Chapter 19 [Creating an Info File], page 139.)

The split off files are called the indirect subfiles.

Info files are split to save memory. With smaller files, Emacs does not have make such a large buffer to hold the information.

If an Info file has more than 30 nodes, you should also make a tag table for it. See Section D.5.1 [Using Info-validate], page 180, for information about creating a tag table. (Again, tag tables are usually created automatically by the formatting command; you only need to create a tag table yourself if you are doing the job manually. Most likely, you will do this for a large, unsplit file on which you have run Info-validate.)

Visit the file you wish to tagify and split and type the two commands:

    M-x Info-tagify
    M-x Info-split

(Note that the ‘I’ in ‘Info’ is upper case.)

When you use the Info-split command, the buffer is modified into a (small) Info file which lists the indirect subfiles. This file should be saved in place of the original visited file. The indirect subfiles are written in the same directory the original file is in, with names generated by appending ‘-’ and a number to the original file name.

The primary file still functions as an Info file, but it contains just the tag table and a directory of subfiles.

182 Texinfo 2.0

Appendix E Refilling Paragraphs

If a paragraph contains long @-constructs, it may look badly filled af-

ter texinfo-format-bufferregion or texinfo-format-buffer is through

with it. This is because neither the texinfo-format-region nor the texinfo-format-buffer command refills paragraphs after processing although both TEX and makeinfo do.

The @-commands are removed from the text but the lines are not refilled so some are much shorter than they were. To cause the texinfo-format-region and texinfo-format-buffer commands to refill such a paragraph, put @refill at the end of the paragraph. This command refills a paragraph in the Info file after all the other processing has been done. @refill has no effect on TEX, which always fills everything that ought to be filled.

For example,

    To use @code{foo}, pass @samp{xx%$} and
    @var{flag} and type @kbd{x} after running
    @code{make-foo}.@refill

produces (in the Info file)

    To use `foo', pass `xx%$' and FLAG and type `x' after
    running `make-foo'.

whereas without the @refill it would produce

    To use `foo', pass `xx%$' and
    FLAG and type `x' after running
    `make-foo'.

with the line broken at the same place as in the Texinfo input file.

Do not put a space before @refill; otherwise the command might be put at the beginning of the line when you refill the paragraph in the Texinfo file with M-q (fill-paragraph). If this were to happen, the @refill command might fail to work.

Do not put braces after @refill. The @refill command is the only ‘within-paragraph’ @-command that does not take braces. Because an @refill command is placed at the end of a paragraph and never at the beginning of a line, the braces are not necessary.

Also, do not end a paragraph that uses either @* or @w with an @refill

command, since texinfo-format-buffer or texinfo-format-buffer will

then refill the paragraph in spite of those commands.

184 Texinfo 2.0

Appendix F Sample Permissions

Texinfo files should contain sections that tell the readers that they have the right to copy and distribute the Info file, the printed manual and any accompanying software. This appendix contains the standard text of the Free Software Foundation copying permission notice. For an example of the text that could be used for the Distribution, General Public License and NO WARRANTY sections of a document, see the latest version of the GNU Emacs Manual.

The texts of the Free Software Foundation copying permission notice in the @ifinfo section and in the @titlepage section are slightly different.

The @ifinfo section usually begins with a line that says what the file documents. This is what a person reading the unprocessed Texinfo file or using the advanced Info command g * will see first. See Info file ‘info’, node ‘Expert’, for more information. (A reader usesusing the regular Info commands will usually start reading at the first node and skip this first section, which is not in a node.)

In the @ifinfo section, the summary sentence should be followed by a copyright notice and then by the copying permission notice. One of the copying permission paragraphs is enclosed in @ignore and @end ignore commands. This paragraph states that the Texinfo file can be processed through TEX and printed, provided the printed manual carries the proper copying permission notice. This paragraph is not made part of the Info file since it is not relevant to the Info file; but it is a mandatory part of the Texinfo file since it permits people to process the Texinfo file in TEX.

In the printed manual, the Free Software Foundation copying permission notice follows the copyright notice and publishing information and is located within the region delineated by the @titlepage and @end titlepage commands. The copying permission notice is exactly the same as the notice in the @ifinfo section except that the paragraph enclosed in @ignore and @end ignore commands is not part of the notice.

To make it simple to copy the permission notice into each section of the Texinfo file, the complete permission notices for each section are reproduced in full below.

Note that you may have to specify the correct name of a section mentioned in the permission notice. For example, in the GDB Manual, the name of the section referring to the General Public License is called the “GDB General Public License”, but in the sample shown below, that section is referred to generically as the “General Public License”.

F.1 ‘ifinfo’ Copying Permissions

In the @ifinfo section of the Texinfo file, the standard Free Software Foundation permission notices reads as follows:

This file documents . . .
Copyright 1988 Free Software Foundation, Inc.
Permission is granted to make and distribute verbatim
copies of this manual provided the copyright notice and
this permission notice are preserved on all copies.
@ignore
Permission is granted to process this file through TeX
and print the results, provided the printed document
carries a copying permission notice identical to this
one except for the removal of this paragraph (this
paragraph not being relevant to the printed manual).
@end ignore
Permission is granted to copy and distribute modified
versions of this manual under the conditions for
verbatim copying, provided also that the sections
entitled ``Distribution'' and ``General Public License''
are included exactly as in the original, and provided
that the entire resulting derived work is distributed
under the terms of a permission notice identical to this
one.
Permission is granted to copy and distribute
translations of this manual into another language, under
the above conditions for modified versions, except that
the sections entitled ``Distribution'' and ``General
Public License'' may be included in a translation
approved by the author instead of in the original
English.

F.2 Titlepage Copying Permissions

In the @titlepage section of the Texinfo file, the standard Free Software Foundation copying permission notice follows the copyright notice and publishing information. The standard phrasing is:

Permission is granted to make and distribute verbatim
copies of this manual provided the copyright notice and
this permission notice are preserved on all copies.
Permission is granted to copy and distribute modified
versions of this manual under the conditions for
verbatim copying, provided also that the sections
entitled ``Distribution'' and ``General Public License''
are included exactly as in the original, and provided
that the entire resulting derived work is distributed
under the terms of a permission notice identical to this
one.
Permission is granted to copy and distribute
translations of this manual into another language, under
the above conditions for modified versions, except that
the sections entitled ``Distribution'' and ``General
Public License'' may be included in a translation
approved by the author instead of in the original
English.

188 Texinfo 2.0

Appendix G @-Command Syntax

The character ‘@’ is used to start special Texinfo commands. (It has the same meaning that ‘\’ has in plain TEX.) Texinfo has four types of @-command:

1. Non-alphabetic commands.

These commands start with @ followed by a punctuation mark or other character that is not part of the alphabet. Non-alphabetic commands are almost always part of the text within a paragraph, and never take any argument. The two characters (@ and the other one) are complete in themselves; none is followed by braces. The non-alphabetic commands are: @., @:, @*, @@, @{, and @}.

2. Alphabetic commands that do not require arguments.

These commands start with @ followed by a word followed by leftand right-hand braces. These commands insert special symbols in the document; they do not require arguments. For exam-

ple, @dots{} ⇒ ‘. . .’, @equiv{} ⇒ ‘ ≡ ’, @TeX{} ⇒ ‘TEX’, and
@bullet{} ⇒ ‘•’.

3. Alphabetic commands that do require arguments.

These commands start with @ followed by a letter or a word, followed by an argument within braces. For example, the command @dfn indicates the introductory or defining use of a term; it

is used as follows: ‘In Texinfo, @@-commands are @dfn{mark-
up} commands.’

4. Alphabetic commands that occupy an entire line.

These commands occupy an entire line. The line starts with @, followed by the name of the command (a word) such as @center or @cindex. If no argument is needed, the word is followed by the end of the line. If there is an argument, it is separated from the command name by a space. Braces are not used.

Thus, the alphabetic commands fall into three classes that have different argument syntax. You cannot tell to which class a command belongs by the appearance of its name, but you can tell by the command’s meaning: if the command creates a special character, it is in class 2 and does not require an argument; if it makes sense to use the command together with other text as part of a paragraph, the command is in class 3 and must be followed by an argument in braces; otherwise, it is in class 4 and uses the rest of the line as its argument.

The purpose of having a different syntax for commands of classes 3 and 4 is to make Texinfo files easier to read, and also to help the GNU Emacs paragraph and filling commands work properly. There is only one exception to this rule: the command @refill, which is always used at the end of a paragraph immediately following the final period or other punctuation character. @refill takes no argument and does not require braces. @refill never confuses the Emacs paragraph commands because it cannot appear at the beginning of a line.

Command Index 191

Command Index

This is an alphabetical list of all the @-commands. To make the list easier to use, the commands are listed without their preceding ‘@’, except for the @def… commands, which are all preceded by an ‘@’ since they have been indexed automatically by the @deffn command in which they are described.

*

* (force line break) . . . . . . . . . . . . . . . . . . 125

.

. (true end of sentence) . . . . . . . . . . . . . . . 80

:

: (suppress widening). . . . . . . . . . . . . . . . .80

@

@ (single ‘@’) . . . . . . . . . . . . . . . . . . . . . . . . . 79 @. (true end of sentence) . . . . . . . . . . . . . . 80 @defcv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 @deffn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 @defivar . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 @defmac . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 @defmethod . . . . . . . . . . . . . . . . . . . . . . . . . 112 @defop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 @defopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 @defspec . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 @deftp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 @deftypefn . . . . . . . . . . . . . . . . . . . . . . . . . 107 @deftypefun. . . . . . . . . . . . . . . . . . . . . . . . 108 @deftypefvar. . . . . . . . . . . . . . . . . . . . . . .110 @deftypevr . . . . . . . . . . . . . . . . . . . . . . . . . 109 @defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 @defvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 @defvr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

{

{ (single ‘{’) . . . . . . . . . . . . . . . . . . . . . . . . . 79

}

} (single ‘}’) . . . . . . . . . . . . . . . . . . . . . . . . . 79

A

appendix . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 appendixsec . . . . . . . . . . . . . . . . . . . . . . . . . 48 appendixsection . . . . . . . . . . . . . . . . . . . . 49 appendixsubsec. . . . . . . . . . . . . . . . . . . . . .49 appendixsubsubsec . . . . . . . . . . . . . . . . . . 50 author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

B

b (bold font) . . . . . . . . . . . . . . . . . . . . . . . . . 76 br (paragraph breaks) . . . . . . . . . . . . . . . 127 buffer-end . . . . . . . . . . . . . . . . . . . . . . . . . 102 bullet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 bye . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 43

C

c (comment) . . . . . . . . . . . . . . . . . . . . . . . . 191 center . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 chapheading . . . . . . . . . . . . . . . . . . . . . . . . . 48 chapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 cindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 cite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 copyright. . . . . . . . . . . . . . . . . . . . . . . .35, 81 ctrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

D

defcodeindex. . . . . . . . . . . . . . . . . . . . . . .123 defindex . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 dfn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 dots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

E

emph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83, 95 end titlepage . . . . . . . . . . . . . . . . . . . . . . . 36 enumerate . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 exdent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

F

file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 filll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 finalout . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 findex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 flushleft . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 flushright . . . . . . . . . . . . . . . . . . . . . . . . . . 89 foo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 foobar. . . . . . . . . . . . . . . . . . . . . . . . .103, 107 foobar-int . . . . . . . . . . . . . . . . . . . . . . . . . 108 footnote . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 forward-word. . . . . . . . . . . . . . . . . . . . . . .101 ftable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

G

group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

H

heading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 headings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

I

i (italic font) . . . . . . . . . . . . . . . . . . . . . . . . 76 ifinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 iftex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 ignore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Info-validate. . . . . . . . . . . . . . . . . . . . . .180 inforef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96, 98 itemize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 itemx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

K

kbd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 kindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

L

lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

M

majorheading. . . . . . . . . . . . . . . . . . . . . . . . 48 makeinfo-buffer . . . . . . . . . . . . . . . . . . . 143 makeinfo-kill-job . . . . . . . . . . . . . . . . . 144

                                        makeinfo-recenter-output-buffer . . 144

makeinfo-region . . . . . . . . . . . . . . . . . . . 143 menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 minus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

N

need . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 noindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

O

occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180

                                        occur-mode-goto-occurrence . . . . . . . . . 17

P

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 pindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 printindex . . . . . . . . . . . . . . . . . . . . . . . . . . 41 pxref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

Q

quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

R

r (Roman font) . . . . . . . . . . . . . . . . . . . . . . 76 ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 refill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

S

samp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 sc (small caps font) . . . . . . . . . . . . . . . . . . 76 section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 setchapternewpage . . . . . . . . . . . . . . . . . . 31 setfilename . . . . . . . . . . . . . . . . . . . . . . . . . 29 settitle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 shortcontents. . . . . . . . . . . . . . . . . . . . . . .42 smallbook . . . . . . . . . . . . . . . . . . . . . . . . . . 137 smallexample. . . . . . . . . . . . . . . . . . . . . . . . 87 smalllisp . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 sp (line spacing) . . . . . . . . . . . . . . . . . . . . 126

Command Index 193

sp (titlepage line spacing) . . . . . . . . . . . . . 33 strong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 subheading . . . . . . . . . . . . . . . . . . . . . . . . . . 49 subsection . . . . . . . . . . . . . . . . . . . . . . . . . . 49 subsubheading. . . . . . . . . . . . . . . . . . . . . . .50 subsubsection. . . . . . . . . . . . . . . . . . . . . . .50 subtitle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 summarycontents . . . . . . . . . . . . . . . . . . . . 42 syncodeindex. . . . . . . . . . . . . . . . . . . . . . .122 synindex . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

T

t (typewriter font) . . . . . . . . . . . . . . . . . . . 76 table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 texindex . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

texinfo-all-menus-update . . . . . . . . . . . 19
texinfo-every-node-update . . . . . . . . . . 18
texinfo-format-buffer . . . . . 22, 144, 145
texinfo-format-region . . . . . 22, 144, 145
texinfo-indent-menu-description . . . 20
texinfo-insert-@code . . . . . . . . . . . . . . . 16
texinfo-insert-@dfn . . . . . . . . . . . . . . . . 16
texinfo-insert-@end . . . . . . . . . . . . . . . . 16
texinfo-insert-@item . . . . . . . . . . . . . . . 16
texinfo-insert-@node . . . . . . . . . . . . . . . 16
texinfo-insert-@samp . . . . . . . . . . . . . . . 16
texinfo-insert-@var . . . . . . . . . . . . . . . . 16
texinfo-insert-braces . . . . . . . . . . . . . . 16
texinfo-insert-node-lines . . . . . . . . . . 20

texinfo-make-menu . . . . . . . . . . . . . . . . . . 18

texinfo-master-menu . . . . . . . . . . . . . . . . 19

                                        texinfo-multiple-files-update . 21, 165
                                        texinfo-sequentially-update-node . . 21
                                        texinfo-show-structure. . . . . . . . . 17, 178

texinfo-tex-buffer . . . . . . . . . . . . . . . . . 22 texinfo-tex-print . . . . . . . . . . . . . . . . . . 22 texinfo-tex-region . . . . . . . . . . . . . . . . . 22

                                        texinfo-update-node . . . . . . . . . . . . . . . . 18

tindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 titlefont . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

U

unnumbered . . . . . . . . . . . . . . . . . . . . . . . . . . 47 unnumberedsec. . . . . . . . . . . . . . . . . . . . . . .48 unnumberedsubsec . . . . . . . . . . . . . . . . . . . 49

                                        unnumberedsubsubsec . . . . . . . . . . . . . . . . 50

up-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

V

var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 vindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 vskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

W

w (prevent line break). . . . . . . . . . . . . . . .126

X

xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

194 Texinfo 2.0

Concept Index 195

Concept Index

.

¯ .cshrc initialization file. . . . . . . . . . . . . . . 136 ¯ .profile initialization file. . . . . . . . . . . . . . 136

@

@-Command list . . . . . . . . . . . . . . . . . . . . . 149 @-Command Syntax . . . . . . . . . . . . . . . . . 191 @-commands . . . . . . . . . . . . . . . . . . . . . . . . . . 6

A

Abbreviations for keys . . . . . . . . . . . . . . . . 72 Adding a new info file . . . . . . . . . . . . . . . 147 Alphabetical @-command list . . . . . . . . . 149 Another Info directory . . . . . . . . . . . . . . . 148 Automatic pointer creation. . . . . . . . . . .142 Automatically insert nodes, menus . . . . . 17

B

Badly referenced nodes . . . . . . . . . . . . . . 180 Beginning a Texinfo file . . . . . . . . . . . . . . . 27 Beginning line of a Texinfo file . . . . . . . . 29 Black rectangle in hardcopy . . . . . . . . . . 137 Book, printing small . . . . . . . . . . . . . . . . . 137 Braces, inserting . . . . . . . . . . . . . . . . . . . . . 79 Breaks in a line . . . . . . . . . . . . . . . . . . . . . 125 Breaks in a paragraph . . . . . . . . . . . . . . . 127 Buffer formatting and printing. . . . . . . . .22 Bullets, inserting . . . . . . . . . . . . . . . . . . . . . 80

C

Catching errors with Info formatting . . 175 Catching errors with TEX formatting. . 176 Catching mistakes . . . . . . . . . . . . . . . . . . . 175 Chapter structuring . . . . . . . . . . . . . . . . . . 45 Characteristics of printed manual . . . . . . . 5 Checking for badly referenced nodes . . 180 Combining indices. . . . . . . . . . . . . . . . . . . 121 Command definitions . . . . . . . . . . . . . . . . 113 Command invocation . . . . . . . . . . . . . . . . . 57 Commands to insert single characters . . 79 Commands using ordinary TeX . . . . . . . 129 Commands, Indicating them in text …. 69 Commands, inserting them . . . . . . . . . . . . 15 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Compile command for formatting . . . . . 135 Conditionally visible text . . . . . . . . . . . . 129 Conditions for copying Texinfo . . . . . . . . . 1 Contents, table of . . . . . . . . . . . . . . . . . . . . 42 Conventions for writing definitions …. 113 Conventions, syntactic . . . . . . . . . . . . . . . . . 7 Copying conditions . . . . . . . . . . . . . . . . . . . . 1 Copying permissions. . . . . . . . . . . . . . . . . 187 Copyright page. . . . . . . . . . . . . . . . . . . . . . .35 Correcting mistakes . . . . . . . . . . . . . . . . . 175 Create nodes, menus automatically …. 17 Creating an Info file . . . . . . . . . . . . . . . . . 139 Creating an unsplit file . . . . . . . . . . . . . . 181 Creating index entries . . . . . . . . . . . . . . . 120 Creating indices. . . . . . . . . . . . . . . . . . . . . 119 Creating pointers automatically . . . . . . 142 Cross reference parts . . . . . . . . . . . . . . . . . 60 Cross references . . . . . . . . . . . . . . . . . . . . . . 59 Cross references using @inforef . . . . . . . 67 Cross references using @pxref . . . . . . . . . 66 Cross references using @ref . . . . . . . . . . . 65 Cross references using @xref . . . . . . . . . . 61

D

Debugging the Texinfo structure . . . . . . 175 Debugging with Info formatting . . . . . . 175 Debugging with TEX formatting . . . . . . 176 Defining indexing entries . . . . . . . . . . . . . 120 Defining new indices. . . . . . . . . . . . . . . . . 123 Definition commands . . . . . . . . . . . . . . . . 101 Definition conventions . . . . . . . . . . . . . . . 113 Definition template . . . . . . . . . . . . . . . . . . 101 Definitions, Indicating them in text …. 69 Different cross reference commands …. 59 ‘dir’ directory for Info installation …. 147 ‘dir’ file listing . . . . . . . . . . . . . . . . . . . . . 147 Display environment. . . . . . . . . . . . . . . . . . 87 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . 39 Dots, inserting . . . . . . . . . . . . . . . . . . . . . . . 80 dvi file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

E

Ellipsis, inserting . . . . . . . . . . . . . . . . . . . . . 80 Emacs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .15 Emacs shell, printing from . . . . . . . . . . . 133 Emphasizing text. . . . . . . . . . . . . . . . . . . . .75 End Node footnote style . . . . . . . . . . . . . 117 End of header line. . . . . . . . . . . . . . . . . . . . 32 End titlepage starts headings . . . . . . . . . . 36 Ending a Texinfo file . . . . . . . . . . . . . . . . . 41 Entries for an index . . . . . . . . . . . . . . . . . 120 Entries, making index . . . . . . . . . . . . . . . 119 Enumeration. . . . . . . . . . . . . . . . . . . . . . . . . 97 Equivalence, indicating it . . . . . . . . . . . . . 93 Error message, indicating it . . . . . . . . . . . 92 Evaluation special glyph . . . . . . . . . . . . . . 91 Example menu . . . . . . . . . . . . . . . . . . . . . . . 56 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Expansion, indicating it. . . . . . . . . . . . . . .91

F

File beginning . . . . . . . . . . . . . . . . . . . . . . . 27 File ending . . . . . . . . . . . . . . . . . . . . . . . . . . 41 File section-structure, showing it. . . . . . .16 Files, Indicating them within text. . . . . .69 Filling paragraphs . . . . . . . . . . . . . . . . . . . 185 Final output . . . . . . . . . . . . . . . . . . . . . . . . 137 Finding badly referenced nodes . . . . . . . 180 First line of a Texinfo file . . . . . . . . . . . . . 29 Fonts for indices . . . . . . . . . . . . . . . . . . . . 122 Fonts for printing, not Info . . . . . . . . . . . . 76 Footings. . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . .117 Format and print in Texinfo mode …. 133 Format environment . . . . . . . . . . . . . . . . . . 88 Format with the compile command … 135 Formatting a file for Info. . . . . . . . . . . . .139 Formatting commands . . . . . . . . . . . . . . . . . 6 Formatting examples . . . . . . . . . . . . . . . . . 84 Formatting for Info . . . . . . . . . . . . . . . . . . . 21 Formatting for printing . . . . . . . . . . . . . . . 22 Formatting headings and footings. . . . . 169 Formatting requirements . . . . . . . . . . . . . 135 Frequently used commands, inserting…15 Function definitions . . . . . . . . . . . . . . . . . 113

G

General syntactic conventions . . . . . . . . . . 7 Generating menus with indices . . . . . . . . 41 Glyphs for examples . . . . . . . . . . . . . . . . . . 91 GNU Emacs . . . . . . . . . . . . . . . . . . . . . . . . . 15 GNU Emacs shell, printing from . . . . . . 133 Group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .127

H

Hardcopy, printing it . . . . . . . . . . . . . . . . 131 Hboxes, overfull . . . . . . . . . . . . . . . . . . . . . 137 Header for Texinfo files . . . . . . . . . . . . . . . 29 Header of a Texinfo file . . . . . . . . . . . . . . . 29 Headings . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Headings begin. . . . . . . . . . . . . . . . . . . . . . .36 Highlighting . . . . . . . . . . . . . . . . . . . . . . . . . 69 Holding text together vertically. . . . . . . 127

I

If text conditionally visible . . . . . . . . . . . 129 ‘ifinfo’ permissions. . . . . . . . . . . . . . . . . 188 Ignored text . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Include files. . . . . . . . . . . . . . . . . . . . . . . . .165 Include files, how to use. . . . . . . . . . . . . .165 Indentation undoing . . . . . . . . . . . . . . . . . . 88 Index entries. . . . . . . . . . . . . . . . . . . . . . . .120 Index entries, making . . . . . . . . . . . . . . . . 119 Index font types. . . . . . . . . . . . . . . . . . . . .120 Indexing table entries automatically. …99 Indicating commands, definitions, etc… 69 Indicating evaluation . . . . . . . . . . . . . . . . . 91 Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 Indices, combining them . . . . . . . . . . . . . 121 Indices, defining new . . . . . . . . . . . . . . . . 123 Indices, printing and menus . . . . . . . . . . . 41 Indices, sorting. . . . . . . . . . . . . . . . . . . . . .131 Indices, two letter names. . . . . . . . . . . . . 122 Indirect subfiles . . . . . . . . . . . . . . . . . . . . . 139 Info file installation. . . . . . . . . . . . . . . . . .147 Info file requirement: @setfilename. … 29 Info file, listing new one. . . . . . . . . . . . . .147 Info file, splitting manually. . . . . . . . . . .182 Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Info formatting. . . . . . . . . . . . . . . . . . . . . . .21 Info installed in another directory. . . . .148 Info validating a large file . . . . . . . . . . . . 180 Info, creating an on-line file . . . . . . . . . . 139 Initialization file for TEX input . . . . . . . 136 Insert nodes, menus automatically . . . . . 17 Inserting @, braces, and periods . . . . . . . . 79 Inserting dots . . . . . . . . . . . . . . . . . . . . . . . . 80 Inserting ellipsis. . . . . . . . . . . . . . . . . . . . . . 80 Inserting frequently used commands. …15 Inserting special characters and symbols79 Installing an Info file . . . . . . . . . . . . . . . . 147 Installing Info in another directory. … 148 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . 39 Invoking commands, convention for …. 57 Itemization . . . . . . . . . . . . . . . . . . . . . . . . . . 95

K

Keys, recommended names . . . . . . . . . . . . 72

L

License agreement . . . . . . . . . . . . . . . . . . . . 39 Line breaks . . . . . . . . . . . . . . . . . . . . . . . . . 125 Line breaks, preventing . . . . . . . . . . . . . . 126 Line spacing . . . . . . . . . . . . . . . . . . . . . . . . 126 Lisp example . . . . . . . . . . . . . . . . . . . . . . . . 86 List of @-Commands . . . . . . . . . . . . . . . . . 149 Listing a new info file. . . . . . . . . . . . . . . .147 Lists and tables, making them . . . . . . . . . 95 Local variables . . . . . . . . . . . . . . . . . . . . . . 135 Looking for badly referenced nodes … 180

M

Macro definitions. . . . . . . . . . . . . . . . . . . .113 Make Node footnote style . . . . . . . . . . . . 117 makeinfo inside Emacs . . . . . . . . . . . . . . 143 makeinfo options. . . . . . . . . . . . . . . . . . . .139 Making a printed manual . . . . . . . . . . . . 131 Making a tag table manually . . . . . . . . . 182 Making breaks . . . . . . . . . . . . . . . . . . . . . . 125 Making cross references . . . . . . . . . . . . . . . 59 Making lists and tables . . . . . . . . . . . . . . . 95 Manual characteristics, printed . . . . . . . . . 5 Marking text within a paragraph . . . . . . 69 Marking words and phrases . . . . . . . . . . . 69 Master menu. . . . . . . . . . . . . . . . . . . . . . . . .37 Master menu parts . . . . . . . . . . . . . . . . . . . 38 Menu example . . . . . . . . . . . . . . . . . . . . . . . 56 Menu item writing . . . . . . . . . . . . . . . . . . . 55 Menus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .55 Menus generated with indices . . . . . . . . . 41 META key . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Minimal Texinfo file . . . . . . . . . . . . . . . . . . . 9 Mistakes, catching. . . . . . . . . . . . . . . . . . .175 Mode, using Texinfo . . . . . . . . . . . . . . . . . . 15 Must have in Texinfo file . . . . . . . . . . . . . . . 9

N

Names for indices . . . . . . . . . . . . . . . . . . . 122 Names recommended for keys . . . . . . . . . 72 Need space at page bottom. . . . . . . . . . .128 New index defining . . . . . . . . . . . . . . . . . . 123 New info file, listing it . . . . . . . . . . . . . . . 147 Node line writing. . . . . . . . . . . . . . . . . . . . .53 Node, Top . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Nodes in other Info files. . . . . . . . . . . . . . .57 Nodes, catching mistakes. . . . . . . . . . . . .175 Nodes, checking for badly referenced . . 180

O

Occurrences, listing with @occur . . . . . . 180 Option definitions . . . . . . . . . . . . . . . . . . . 114 Optional and repeated parameters …. 102 Options for makeinfo . . . . . . . . . . . . . . . . 139 Ordinary TEX commands, using . . . . . . 129 Other Info files’ nodes . . . . . . . . . . . . . . . . 57 Overfull “Hboxes”. . . . . . . . . . . . . . . . . . . 137 Overview of Texinfo . . . . . . . . . . . . . . . . . . . 3

P

Page breaks . . . . . . . . . . . . . . . . . . . . . . . . 127 Page headings . . . . . . . . . . . . . . . . . . . . . . 169 Page numbering. . . . . . . . . . . . . . . . . . . . . 169 Page sizes for books . . . . . . . . . . . . . . . . . 137 Pages, starting odd . . . . . . . . . . . . . . . . . . . 31 Paragraph breaks . . . . . . . . . . . . . . . . . . . 127 Paragraph, marking text within . . . . . . . 69 Parameters, optional and repeated …. 102 Part of file formatting and printing …. 22 Parts of a cross reference. . . . . . . . . . . . . .60 Parts of a master menu . . . . . . . . . . . . . . . 38 Periods, inserting. . . . . . . . . . . . . . . . . . . . .79 Permissions . . . . . . . . . . . . . . . . . . . . . . . . . 187 Permissions, printed . . . . . . . . . . . . . . . . . . 35 Point, indicating it in a buffer . . . . . . . . . 93 Pointer validation with makeinfo . . . . . 141 Preventing breaks . . . . . . . . . . . . . . . . . . . 125 Print and format in Texinfo mode …. 133 Printed manual characteristics . . . . . . . . . . 5 Printed output, indicating it . . . . . . . . . . 92 Printed permissions . . . . . . . . . . . . . . . . . . 35 Printing a region or buffer . . . . . . . . . . . . 22 Printing an index . . . . . . . . . . . . . . . . . . . . 41 Printing from an Emacs shell . . . . . . . . . 133 Printing hardcopy . . . . . . . . . . . . . . . . . . . 131 Problems, catching . . . . . . . . . . . . . . . . . . 175

Q

Quotations . . . . . . . . . . . . . . . . . . . . . . . . . . 84

R

Recommended names for keys . . . . . . . . . 72 Rectangle, ugly, black in hardcopy …. 137 References . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 References using @inforef . . . . . . . . . . . . 67 References using @pxref . . . . . . . . . . . . . . 66 References using @ref . . . . . . . . . . . . . . . . 65 References using @xref . . . . . . . . . . . . . . . 61 Referring to other Info files. . . . . . . . . . . .57 Refilling paragraphs . . . . . . . . . . . . . . . . . 185 Region formatting and printing . . . . . . . . 22 Region printing in Texinfo mode . . . . . . 133 Repeated and optional parameters …. 102 Requirements for formatting . . . . . . . . . 135 Requirements for updating commands. . 20 Result of an expression . . . . . . . . . . . . . . . 91 Running an Info formatter . . . . . . . . . . . . 21 Running Info-validate . . . . . . . . . . . . . 180 Running makeinfo in Emacs . . . . . . . . . 143

S

Sample function definition . . . . . . . . . . . 113 Sample Texinfo file . . . . . . . . . . . . . . . . . . . 10 Section-structure of a file, showing it … 16 Shell, printing from. . . . . . . . . . . . . . . . . .133 Shell, running makeinfo in . . . . . . . . . . . 143 Showing the section-structure of a file . . 16 Showing the structure of a file . . . . . . . . 178 Single characters, commands to insert . . 79 Size of printed book . . . . . . . . . . . . . . . . . 137 Small book size . . . . . . . . . . . . . . . . . . . . . 137 Small example environment . . . . . . . . . . . 87 Small lisp environment. . . . . . . . . . . . . . . .87 Sorting indices . . . . . . . . . . . . . . . . . . . . . . 131 Spaces from line to line . . . . . . . . . . . . . . 126 Special glyphs . . . . . . . . . . . . . . . . . . . . . . . 91 Special insertions. . . . . . . . . . . . . . . . . . . . .79 Special typesetting commands . . . . . . . . . 80 Specifying index entries . . . . . . . . . . . . . . 120 Splitting an Info file manually . . . . . . . . 182 Start of header line . . . . . . . . . . . . . . . . . . . 29 Starting chapters . . . . . . . . . . . . . . . . . . . . . 31 Structure of a file, showing it . . . . . . . . . . 16 Structure, catching mistakes in . . . . . . . 175 Structuring of chapters . . . . . . . . . . . . . . . 45 Subsection-like commands. . . . . . . . . . . . .49 Subsub commands. . . . . . . . . . . . . . . . . . . .50 Syntactic conventions . . . . . . . . . . . . . . . . . . 7

T

Table of contents . . . . . . . . . . . . . . . . . . . . . 42 Tables and lists, making them . . . . . . . . . 95 Tables with indexes . . . . . . . . . . . . . . . . . . 99 Tables, making two-column . . . . . . . . . . . 98 Tabs; don’t use! . . . . . . . . . . . . . . . . . . . . . 198 Tag table, making manually . . . . . . . . . . 182 Template for a definition . . . . . . . . . . . . . 101 TeX commands, using ordinary . . . . . . . 129 TeX index sorting . . . . . . . . . . . . . . . . . . . 131 TeX input initialization . . . . . . . . . . . . . . 136 Texinfo file beginning . . . . . . . . . . . . . . . . . 27 Texinfo file ending. . . . . . . . . . . . . . . . . . . .41 Texinfo file header. . . . . . . . . . . . . . . . . . . .29 Texinfo file minimum . . . . . . . . . . . . . . . . . . 9 Texinfo file section-structure, showing it 16 Texinfo mode . . . . . . . . . . . . . . . . . . . . . . . . 15 Texinfo overview . . . . . . . . . . . . . . . . . . . . . . 3 texinputs environment variable. …136 Text conditionally visible . . . . . . . . . . . . 129 Title page . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Titlepage end starts headings. . . . . . . . . .36 Titlepage permissions. . . . . . . . . . . . . . . .188 Top node . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Tree structuring. . . . . . . . . . . . . . . . . . . . . . 45 Two letter names for indices . . . . . . . . . 122 Two named items for @table . . . . . . . . . . 99 Typesetting commands for dots, etc. … 80

U

Unprocessed text . . . . . . . . . . . . . . . . . . . . . . 9 Unsplit file creation . . . . . . . . . . . . . . . . . 181 Updating nodes and menus. . . . . . . . . . . .17 Updating requirements . . . . . . . . . . . . . . . 20

V

Validating a large file . . . . . . . . . . . . . . . . 180 Validation of pointers. . . . . . . . . . . . . . . . 141 Value of an expression, indicating . . . . . . 91 Variable definitions . . . . . . . . . . . . . . . . . . 114 Vertically holding text together . . . . . . . 127 Visibility of conditional text. . . . . . . . . . 129

W

Words and phrases, marking them . . . . . 69 Writing a menu item . . . . . . . . . . . . . . . . . 55 Writing a node Line . . . . . . . . . . . . . . . . . . 53

200 Texinfo 2.0

Short Contents

Texinfo Copying Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .1 1 Overview of Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Using Texinfo Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .15 3 Beginning a Texinfo File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4 Ending a Texinfo File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .41 5 Chapter Structuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6 Nodes and Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7 Making Cross References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 8 Marking Words and Phrases . . . . . . . . . . . . . . . . . . . . . . . . . 69 9 Special Insertions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .79 10 Quotations and Examples. . . . . . . . . . . . . . . . . . . . . . . . . . .83 11 Special Glyphs for Examples . . . . . . . . . . . . . . . . . . . . . . . 91 12 Making Lists and Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 13 Definition Commands: @deffn, etc. . . . . . . . . . . . . 101 14 Footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 15 Creating Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 16 Making and Preventing Breaks . . . . . . . . . . . . . . . . . . . 125 17 Conditionally Visible Text . . . . . . . . . . . . . . . . . . . . . . . . .129 18 Printing Hardcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .131 19 Creating an Info File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 20 Installing an Info File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 Appendix A @-Command List . . . . . . . . . . . . . . . . . . . . . . . . 149 Appendix B Include Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Appendix C Page Headings . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 Appendix D Formatting Mistakes . . . . . . . . . . . . . . . . . . . . 173 Appendix E Refilling Paragraphs. . . . . . . . . . . . . . . . . . . . .183 Appendix F Sample Permissions . . . . . . . . . . . . . . . . . . . . . 185 Appendix G @-Command Syntax . . . . . . . . . . . . . . . . . . . . 189 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .195

ii Texinfo 2.0

Table of Contents

Texinfo Copying Conditions . . . . . . . . . . . . . . . . . 1

1 Overview of Texinfo . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Printed Manuals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .5

1.2.1 Obtaining TEX . . . . . . . . . . . . . . . . . . . . . . . 6 1.3 @-commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 General Syntactic Conventions . . . . . . . . . . . . . . . . 7 1.5 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . iii 1.6 What a Texinfo File Must Have . . . . . . . . . . . . . . . 9 1.7 Six Parts of a Texinfo File . . . . . . . . . . . . . . . . . . . . 10 1.8 A Short Sample Texinfo File . . . . . . . . . . . . . . . . . 10

2 Using Texinfo Mode . . . . . . . . . . . . . . . . . . . . . . 15

2.1 Inserting Frequently Used Commands. . . . . . . .15 2.2 Showing the Section-Structure of a File . . . . . . 16 2.3 Updating Nodes and Menus . . . . . . . . . . . . . . . . . . 17

2.3.1 Updating Requirements. . . . . . . . . . . . .20 2.3.2 Other Updating Commands . . . . . . . . 20

2.4 Formatting for Info . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.5 Formatting and Printing . . . . . . . . . . . . . . . . . . . . . . 22 2.6 Texinfo Mode Summary . . . . . . . . . . . . . . . . . . . . . . 23

3 Beginning a Texinfo File . . . . . . . . . . . . . . . 27

3.1 Sample Texinfo File Beginning . . . . . . . . . . . . . . . 27 3.2 The Texinfo File Header . . . . . . . . . . . . . . . . . . . . . . 29

3.2.1 The First Line of a Texinfo File …. 29 3.2.2 Start of Header. . . . . . . . . . . . . . . . . . . . . .29 3.2.3 @setfilename . . . . . . . . . . . . . . . . . . . . . . . 29 3.2.4 @settitle . . . . . . . . . . . . . . . . . . . . . . . . . . .30 3.2.5 @setchapternewpage . . . . . . . . . . . . . . . 31 3.2.6 End of Header. . . . . . . . . . . . . . . . . . . . . . .32

3.3 Summary and Copying Permissions for Info . 32 3.4 The Title and Copyright Pages . . . . . . . . . . . . . . . 32

3.4.1 @titlepage . . . . . . . . . . . . . . . . . . . . . . . . . 33

      3.4.2 @titlefont, @center, and @sp. . . . .33
      3.4.3 @title, @subtitle, and @author . . 34

3.4.4 Copyright Page and Permissions … 35 3.4.5 Heading Generation . . . . . . . . . . . . . . . . . 36 3.4.6 The @headings Command . . . . . . . . . .36

3.5 The Top Node and Master Menu . . . . . . . . . . . . .37

3.5.1 Parts of a Master Menu. . . . . . . . . . . . .38 3.6 Licensing and Distribution Information . . . . . . 39

4 Ending a Texinfo File . . . . . . . . . . . . . . . . . . . .41

4.1 Index Menus and Printing an Index . . . . . . . . . . 41 4.2 Generating a Table of Contents . . . . . . . . . . . . . . 42 4.3 @bye File Ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

5 Chapter Structuring . . . . . . . . . . . . . . . . . . . . . .45

5.1 Tree Structure of Sections . . . . . . . . . . . . . . . . . . . . 45 5.2 Types of Structuring Command . . . . . . . . . . . . . . 46 5.3 @chapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.4 @unnumbered, @appendix . . . . . . . . . . . . . . . . . . . . . 47

5.5 @majorheading, @chapheading. . . . . . . . . . . . . . .48

5.6 @section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

5.7 @unnumberedsec, @appendixsec, @heading . 48

5.8 The @subsection Command . . . . . . . . . . . . . . . . . 49 5.9 The @subsection-like Commands. . . . . . . . . . . .49 5.10 The ‘subsub’ Commands . . . . . . . . . . . . . . . . . . . . 50

6 Nodes and Menus . . . . . . . . . . . . . . . . . . . . . . . . . .51

6.1 Node and Menu Illustration . . . . . . . . . . . . . . . . . . 51 6.2 @node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.2.1 Writing a Node Line . . . . . . . . . . . . . . . . 53 6.3 @menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

6.3.1 Writing a Menu Item . . . . . . . . . . . . . . . 55 6.3.2 A Menu Example . . . . . . . . . . . . . . . . . . . 56

6.4 Referring to Other Info Files . . . . . . . . . . . . . . . . . 57 6.5 Describing Command Invocation . . . . . . . . . . . . . 57

7 Making Cross References . . . . . . . . . . . . . .59

7.1 Different Cross Reference Commands . . . . . . . . 59 7.2 Parts of a Cross Reference . . . . . . . . . . . . . . . . . . . . 60 7.3 @xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

7.3.1 @xref with One Argument . . . . . . . . . 62 7.3.2 @xref with Two Arguments . . . . . . . . 62 7.3.3 @xref with Three Arguments. . . . . . .63 7.3.4 @xref with Four Arguments . . . . . . . . 63 7.3.5 @xref with Five Arguments . . . . . . . . 64

7.4 @ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.5 @pxref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .66

7.6 @inforef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

8 Marking Words and Phrases. . . . . . . . .69

8.1 Indicating Definitions, Commands, etc. . . . . . . 69 8.1.1 @code{sample-code} . . . . . . . . . . . . . . . . 70 8.1.2 @kbd{keyboard-characters} . . . . . . . . . 71 8.1.3 @key{key-name} . . . . . . . . . . . . . . . . . . . . 71 8.1.4 @samp{text} . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.1.5 @var{metasyntactic-variable} . . . . . . 73 8.1.6 @file{file-name} . . . . . . . . . . . . . . . . . . . 73 8.1.7 @dfn{term}. . . . . . . . . . . . . . . . . . . . . . . . . .74 8.1.8 @cite{reference}. . . . . . . . . . . . . . . . . . . .74 8.1.9 @ctrl{ctrl-char} . . . . . . . . . . . . . . . . . . . . 74

8.2 Emphasizing Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 8.2.1 @emph{text} and @strong{text} … 75 8.2.2 @sc{text}: The Small Caps Font …76 8.2.3 Fonts for Printing, Not Info . . . . . . . . 76

9 Special Insertions . . . . . . . . . . . . . . . . . . . . . . . . . . 79

9.1 Inserting ‘@’, Braces, and Periods . . . . . . . . . . . . 79 9.1.1 Inserting ‘@’—@@ . . . . . . . . . . . . . . . . . . . . 79 9.1.2 Inserting ‘{’ and ‘}’—@{ and @} …. 79 9.1.3 Spacing After Colons and Periods . . 80

9.2 Inserting Ellipsis, Dots, and Bullets . . . . . . . . . . 80 9.2.1 @dots{} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 9.2.2 @bullet{} . . . . . . . . . . . . . . . . . . . . . . . . . . .81

9.3 Inserting TEX and \begin{matrix}c \\ ◯ \end{matrix}. . . . . . . . . . . . . . . . . . . . . . . . . .81

9.3.1 @TeX{}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .81 9.3.2 @copyright{} . . . . . . . . . . . . . . . . . . . . . . . 81

9.4 @minus{}: Inserting a Minus Sign . . . . . . . . . . . . 81

10 Quotations and Examples . . . . . . . . . . . 83

10.1 The Various Environments . . . . . . . . . . . . . . . . . . 83 10.2 @quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 10.3 @example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 10.4 @noindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 10.5 @lisp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .86

10.6 @smallexample and @smalllisp . . . . . . . . . . . . 87

10.7 @display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 10.8 @format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 10.9 @exdent: Undoing a Line’s Indentation. . . . .88 10.10 @flushleft and @flushright . . . . . . . . . . . . . 89

11 Special Glyphs for Examples . . . . . . 91

11.1 ⇒: Indicating Evaluation . . . . . . . . . . . . . . . . . . . 91 11.2 ↦: Indicating an Expansion . . . . . . . . . . . . . . . . 91 11.3 ⊣: Indicating Printed Output . . . . . . . . . . . . . . 92 11.4 error : Indicating an Error Message . . . . . . . 92 11.5 ≡: Indicating Equivalence . . . . . . . . . . . . . . . . . . 93 11.6 Indicating Point in a Buffer . . . . . . . . . . . . . . . . . 93

12 Making Lists and Tables . . . . . . . . . . . . . 95

12.1 Making an Itemized List. . . . . . . . . . . . . . . . . . . . .95 12.2 Making a Numbered List . . . . . . . . . . . . . . . . . . . . 97 12.3 Making a Two-column Table . . . . . . . . . . . . . . . . 98

12.3.1 @ftable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 12.3.2 @itemx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

13 Definition Commands: @deffn, etc.101

13.1 The Template for a Definition . . . . . . . . . . . . . 101 13.2 Optional and Repeated Parameters . . . . . . . .102 13.3 The Definition Commands . . . . . . . . . . . . . . . . . 103

13.3.1 Functions and Similar Entities … 103 13.3.2 Variables and Similar Entities. …106 13.3.3 Functions in Typed Languages …107 13.3.4 Variables in Typed Languages … 109 13.3.5 Object-Oriented Programming … 110 13.3.6 Data Types . . . . . . . . . . . . . . . . . . . . . . . 112

13.4 Conventions for Writing Definitions. . . . . . . .113 13.5 A Sample Function Definition. . . . . . . . . . . . . .113 13.6 A Sample Variable Definition . . . . . . . . . . . . . . 114

14 Footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .117

15 Creating Indices. . . . . . . . . . . . . . . . . . . . . . . . .119

15.1 Making Index Entries. . . . . . . . . . . . . . . . . . . . . . .119 15.2 Predefined Indices . . . . . . . . . . . . . . . . . . . . . . . . . . 119 15.3 Defining the Entries of an Index . . . . . . . . . . . 120 15.4 Combining Indices . . . . . . . . . . . . . . . . . . . . . . . . . . 121 15.5 Defining New Indices . . . . . . . . . . . . . . . . . . . . . . . 123

16 Making and Preventing Breaks.125

16.1 @*: Generate Line Breaks . . . . . . . . . . . . . . . . . . 125 16.2 @w{text}: Prevent Line Breaks. . . . . . . . . . . . .126 16.3 @sp n: Insert Blank Lines . . . . . . . . . . . . . . . . . . 126 16.4 @br: Generate Paragraph Breaks . . . . . . . . . . 127 16.5 @page: Start a New Page. . . . . . . . . . . . . . . . . . .127 16.6 @group: Prevent Page Breaks . . . . . . . . . . . . . . 127 16.7 @need mils: Prevent Page Breaks . . . . . . . . . . 128

17 Conditionally Visible Text . . . . . . . . 129

17.1 Using Ordinary TEX Commands . . . . . . . . . . . 129

18 Printing Hardcopy . . . . . . . . . . . . . . . . . . . . . 131

18.1 How to Print Using Shell Commands . . . . . . 131 18.2 Printing from an Emacs Shell . . . . . . . . . . . . . . 133 18.3 Formatting and Printing in Texinfo Mode . 133 18.4 Using the Local Variables List . . . . . . . . . . . . . 135 18.5 TEX Formatting Requirements Summary . .135 18.6 TEX Input Initialization . . . . . . . . . . . . . . . . . . . . 136 18.7 Overfull “Hboxes” . . . . . . . . . . . . . . . . . . . . . . . . . . 137 18.8 Printing “Small” Books . . . . . . . . . . . . . . . . . . . . 137

19 Creating an Info File . . . . . . . . . . . . . . . . . 139

19.1 Running makeinfo from a Shell. . . . . . . . . . . .139 19.2 Options for makeinfo . . . . . . . . . . . . . . . . . . . . . . 139 19.3 Pointer Validation . . . . . . . . . . . . . . . . . . . . . . . . . . 141 19.4 Creating Pointers Automatically . . . . . . . . . . . 142 19.5 Running makeinfo inside Emacs. . . . . . . . . . .143 19.6 The texinfo-format… Commands . . . . . . . 144 19.7 Tag Files and Split Files . . . . . . . . . . . . . . . . . . . 145

20 Installing an Info File . . . . . . . . . . . . . . . . 147

20.1 Listing a New Info File . . . . . . . . . . . . . . . . . . . . . 147 20.2 Info Files in Other Directories . . . . . . . . . . . . . 148

Appendix A @-Command List . . . . . . . . 149

Appendix B Include Files. . . . . . . . . . . . . . .163

B.1 How to Use Include Files. . . . . . . . . . . . . . . . . . . .163

B.2 texinfo-multiple-files-update . . . . . . . . . 163

B.3 Sample File with @include . . . . . . . . . . . . . . . . . 164 B.4 Evolution of Include Files . . . . . . . . . . . . . . . . . . . 166

Appendix C Page Headings . . . . . . . . . . . 167

C.1 Standard Heading Formats . . . . . . . . . . . . . . . . . 167 C.2 Specifying the Type of Heading. . . . . . . . . . . . .168 C.3 How to Make Your Own Headings. . . . . . . . . .169

Appendix D Formatting Mistakes . 173

D.1 Catching Errors with Info Formatting. . . . . .173 D.2 Catching Errors with TEX Formatting . . . . . 174

D.3 Using texinfo-show-structure . . . . . . . . . . . 176

D.4 Using occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 D.5 Finding Badly Referenced Nodes . . . . . . . . . . . 178

D.5.1 Running Info-validate. . . . . . . . . .178 D.5.2 Creating an Unsplit File . . . . . . . . . . 179 D.5.3 Tagifying a File . . . . . . . . . . . . . . . . . . . 180 D.5.4 Splitting a File Manually . . . . . . . . . 180

Appendix E Refilling Paragraphs . . 183

Appendix F Sample Permissions … 185

F.1 ‘ifinfo’ Copying Permissions . . . . . . . . . . . . . . 186 F.2 Titlepage Copying Permissions . . . . . . . . . . . . . 186

Appendix G @-Command Syntax … 189

Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .191

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195